### v1

# Why Transformers Need Adam: A Hessian Perspective

Abstract
SGD performs worse than Adam by a significant margin on Transformers, but the reason remains unclear. In this work, we provide an explanation of SGD’s failure on Transformers through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs badly on problems with block heterogeneity. To validate that heterogeneity hampers SGD, we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD works well on problems without block heterogeneity but performs badly when the heterogeneity exists. Our initial theoretical analysis indicates that SGD fails because it applies one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. The failure could be rescued if we could assign different learning rates across blocks, as designed in Adam.
1 Introduction
Transformers (Vaswani et al., 2017) have become a major workhorse behind much remarkable progress in AI development (e.g., (Achiam et al., 2023)). However, the understanding of Transformer training remains limited. For instance, Transformer training largely relies on Adam(W) (Kingma and Ba, 2014; Loshchilov and Hutter, 2017). In contrast, Stochastic Gradient Descent with momentum (SGD)  [^1] We introduce the update rules of Adam(W) and SGD in Appendix B.1., which is the de-facto optimizer for convolution neural networks (LeCun et al., 1998) (CNNs), performs poorly on Transformers (see Figure LABEL:fig:cv_figure and LABEL:fig:nlp_figure in Appendix D as evidence). Yet, the cause of failure remains unclear. It is an intriguing question why SGD fails on Transformers. First, from a theoretical perspective, this topic helps reveal more properties of the Transformer training loss landscape, which helps boost understanding of training objectives and current optimizers. Second, from a computational perspective, Adam requires heavier memory storage than SGD. Compared with SGD, Adam needs to store an additional second-order momentum, which takes at least the same memory of the model size. The memory consumption of Adam has become a major overhead in large-scale transformer training (Rasley et al., 2020). Identifying the cause of failure may help researchers design more efficient optimizers.
In this work, we explore the failure mode of SGD through the lens of Hessian. We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum as an example). By theory, the full Hessian spectrum largely determines the behavior of gradient methods (Nesterov, 2013; Goh, 2017; Sun, 2019; Goujaud et al., 2022), so we suspect it can also help uncover the reasons behind SGD failure on Transformers. Using numerical linear algebra tools (Bai et al., 1996), we empirically compare the full spectra of CNNs (the scenarios where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam). Unfortunately, in Figure LABEL:fig_full_spectrum, we find that the spectra on CNNs and Transformers are often largely similar despite the different optimizer behaviors on these tasks. As such, we have not identified critical phenomena in the full Hessian spectra associated with SGD’s failure on Transformers. To reveal the cause of failure, we need a more fine-grained investigation into Hessian.
What else could make SGD fail on Transformers, but not on CNNs? By dissecting the structures of CNNs and Transformers, we observe that CNN architectures are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the interleaved stacking of disparate parameter blocks (e.g. Query, Key, Value blocks in attention and MLP layers). We suspect that these design differences may lead to different properties in optimization. Intuitively, disparate parameter blocks would contribute differently to the total loss. So each block might prefer a distinct treatment by optimizers, which can be provided by Adam but not by SGD. This motivates us to investigate the Hessian spectrum of each parameter block, abbreviated as the blockwise Hessian spectrum.
By inspecting the blockwise Hessian spectrum, we find a possible explanation why SGD fails on Transformers. We find that Transformers are “heterogeneous": for all Transformers we checked, Hessian spectra of different parameter blocks are dramatically different. We then verify that such heterogeneity indeed hampers SGD via extensive experiments on various Transformers, CNNs, MLPs , and quadratic problems. Our contributions are summarized as follows.
  • Exploring the failure mode of SGD on transformers. We provide an explanation why SGD largely lags behind Adam on Transformers through the lens of blockwise Hessian spectrum. This explanation consists of two parts. First, we identify a phenomenon called “block heterogeneity", which states that Hessian spectra are dramatically different across parameter blocks. We find this phenomenon prevalent on all Transformers in our experiments, but not on CNNs. Second, we verify that block heterogeneity hampers SGD. More specifically, on various kinds of Transformers, CNNs, MLPs, we find that SGD always performs badly on problems with block heterogeneity but works well otherwise.
  • Theoretical results on quadratic models. We build various convex quadratic problems, some with block heterogeneity and others without block heterogeneity. We find that GD is significantly worse than Adam on problems with block heterogeneity, but performs similarly to Adam otherwise. We provide a theoretical result which shows that GD indeed works badly on quadratic problems with block heterogeneity. Our analysis suggests that one important reason for (S)GD failure is that it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. The failure could be saved if we assign different learning rates across blocks like Adam.
2 Main Results
2.1 Problem Settings
Notations.
We denote the full-batch training loss as $\mathcal{L}(w)\equiv\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(w)$, where $n$ is the number of minibatches, $\mathcal{L}_{i}(w)$ is the loss of $i$-th minibatch and $w\in\mathbb{R}^{d}$ is the neural network parameters. We denote the gradient and Hessian of the training loss w.r.t. neural network parameters as $\nabla\mathcal{L}(w)\in\mathbb{R}^{d}$ and $\nabla^{2}\mathcal{L}(w)\in\mathbb{R}^{d\times d}$, respectively. We use $\nabla\mathcal{L}_{i}(w)\in\mathbb{R}^{d}$ and $\nabla^{2}\mathcal{L}_{i}(w)\in\mathbb{R}^{d\times d}$ to denote the $i$-th minibatch counterparts. We use $[d]$ to denote the index set $\{1,2,\cdots,d\}$. Given an arbitrary partition $\{\mathcal{D}_{l}\}^{L}_{l=1}$ over $[d]$ with $d_{l}\triangleq|\mathcal{D}_{l}|$, we can split $w$ into $L$ parameter blocks $\{w_{l}\}_{l=1}^{L}$, where $w_{l}=\mathbb{R}^{d_{l}}$ consists of parameters with indexes in the $l$-th block $\mathcal{D}_{l}$. We denote $[\nabla^{2}\mathcal{L}(w)]_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ as the Hessian of $l$-th parameter-block $w_{l}$, where $[\nabla^{2}\mathcal{L}(w)]_{l,i,j}=\frac{\partial^{2}}{\partial_{w_{l,i}}\partial_{w_{l,j}}}\mathcal{L}(w_{l})$. Note that $[\nabla^{2}\mathcal{L}(w)]_{l}$ is the $l$-th principle block sub-matrix of $\nabla^{2}\mathcal{L}(w)$.
Preliminaries.
Hessian of large-scale NNs are intractable to obtain due to the numerical difficulties of computing and storing them. Fortunately, there are extensive numerical linear algebra methods to approximate the Hessian spectrum without calculating the Hessian explicitly. Among which, the Stochastic Lanczos Quadrature method (SLQ) (Bai et al., 1996) is known to be one of the oldest yet the most powerful methods. SLQ is a combination of the stochastic trace estimator (Hutchinson, 1989), the Gaussian quadrature (Golub and Meurant, 2009; Epperson, 2013), and the Lanczos algorithm (Lanczos, 1950). In short, SLQ takes a handful of Hessian-vector products as input, where the vectors are sampled from the normalized Rademacher distribution. It returns a smooth curve on $\mathbb{R}$ that approximates the histograms of eigenvalues (see Figure LABEL:fig_full_spectrum as an example). Since the derivation of SLQ is quite involved, we defer the detailed description of SLQ to Appendix B.2. The algorithmic form of SLQ can be seen in Algorithm 5 in the Appendix. We provide running time analysis and detailed setup of SLQ in Appendix C.1.
In this work, we apply SLQ to various CNNs and Transformers on different datasets. As for CNNs, we study CNNs including ResNet18 (11M) and VGG16 (138M) on ImageNet (He et al., 2016; Simonyan and Zisserman, 2014). As for Transformers, we study ViT-base (86M) on ImageNet (Dosovitskiy et al., 2020), BERT (40M) on Cornell Movie-Dialogs Corpus (Devlin et al., 2018), and GPT2 (125M) on Openwebtext (Radford et al., 2019). For a fair comparison with ResNet18, we also consider smaller models like GPT2-nano [^2] https://github.com/karpathy/nanoGPT/ (11M) on English corpus. The largest model we conduct is GPT2 with $125$ million parameters. Detailed experimental setups are shown in Appendix C.1. For all these Transformer-based tasks, SGD performs significantly worse than Adam; while for CNN-based takes, Adam and SGD are similar. The performance comparison is shown in Figure LABEL:fig:cv_figure and LABEL:fig:nlp_figure in the Appendix. We estimate the spectrum of (1) the full Hessian $\nabla^{2}\mathcal{L}(w)$ and (2) the Hessian of all parameter blocks $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$. For (2), we split $w$ according to the default partition in PyTorch implementation, e.g., Embedding layer, Query in 1st attention layer, Key in 1st attention layer, Value in 1st attention layer, etc..
Note that the term “block" is a bit different from the term “layer". For instance, Query and Key could be in the same layer, but they are different parameter blocks. In the following contexts, we refer to “full Hessian spectrum" as the estimated spectrum of $\nabla^{2}\mathcal{L}(w)$ and “blockwise Hessian spectrum" as the estimated spectrum of $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$.
2.2 Full Hessian Spectrum Is Not Informative Enough
Motivated by classical theory, we start by investigating the full Hessian spectrum of Transformers. We choose to study the full Hessian spectrum for two reasons. First, as mentioned in Section 1, the full Hessian spectrum is known to largely determine the behavior of gradient methods. For instance, the range of the spectrum will affect the convergence rate (Nesterov, 2013). Second, the full Hessian spectrum has been shown to be informative in explaining multiple phenomena in neural nets (NNs), e.g., how BatchNorm accelerates the training (Ghorbani et al., 2019). Due to these reasons, we hypothesize that the full Hessian spectrum can also help uncover the reasons behind SGD failure on Transformers.
We compare the full spectra of CNNs (the scenarios where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam). The results are shown in Figure LABEL:fig_full_spectrum. Unfortunately, our experiments show that the full Hessian spectrum might not be enough to explain the SGD failure on Transformers. We explain as follows. The primary information of the spectrum lies in its (1) dispersion, (2) shape, and (3) evolution along training. As for (1), we find that eigenvalues are dispersed very similarly and we do not observe any notably large eigenvalues in the spectra of Transformers. So (1) does not seem to be related to the failure of SGD. We further inspect (2) and (3). For all CNNs and Transformers in Figure LABEL:fig_full_spectrum, we find similar phenomena: the shape of the spectrum is approximately symmetrical around 0 at initialization. As training proceeds, the majority of large negative eigenvalues disappear and the shape evolves into a combination of a “bulk" and some “outliers". Since the spectral shape and evolution are quite similar on Transformers and CNNs, they cannot explain the failure of SGD on Transformers, either. In summary, we have not identified critical phenomena in the full Hessian spectra linked to SGD’s failure on Transformers.
2.3 Main Findings Through Blockwise Hessian Spectra
We have shown that the full Hessian spectrum seems not sufficient for uncovering the reason why SGD fails on Transformers. What else could make SGD fail on Transformers, but not on CNNs? We find two important features that are overlooked in the full Hessian spectrum analysis above.
  • (1) The Hessian structure is overlooked. There are numerical results showing that the Hessians are close to block-diagonal matrices on MLPs (Collobert, 2004; Roux et al., 2007; Martens and Grosse, 2015). We restate their findings in Appendix C.2. Further, (Collobert, 2004, Section 7) theoretically shows that the block diagonal structure comes from (i) the layer-by-layer design in NNs and (ii) the Cross-Entropy loss. Following this line of work, we also observe a near block-diagonal Hessian in small Transformers in Figure LABEL:fig_block_diagnal, where the variables in each principle block correspond to the parameters of each block in the Transformer. These results suggest that near block-diagonal Hessian might be common in NNs.
  • (2) The build-up rules of Transformers is overlooked. Transformers are built upon stacking of disparate parameter blocks, e.g. Query, Key, Value blocks in attention, and MLP layers. This is largely different from CNNs, which are constructed by the repetitive stacking of similar parameter blocks (convolution layers). Since each parameter blocks in Transformers are designed differently, they might have different properties for optimization, which might further affect the optimizer behavior.
Combining (1) and (2), we hypothesize that the blockwise Hessian spectrum might provide some insights on distinguishing CNNs and Transformers. What extra information can be contained in the blockwise spectrum but not in the full spectrum? By definition, the blockwise Hessians form the principal block sub-matrix of the full Hessian. For block diagonal matrices, blockwise Hessian encodes the location of eigenvalues, i.e., which block does an eigenvalue (of the full Hessian) lie in, while the full spectrum ignores this. In the following, we study blockwise Hessian spectra on various models. We will show that blockwise spectra indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
We here demonstrate the shape of blockwise spectra in CNNs and Transformers. We sample four blocks per model and present the spectra in Figure LABEL:fig_blockwise_spectrum. In Transformers, the Hessian spectra of embedding, attention, and MLP blocks are largely different. In contrast, in CNNs, the spectra of convolution layers are similar. We further verify this observation for the rest of the parameter blocks. We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs. We show the results in Figure LABEL:fig_blockwise_heatmap.
The results in Figure LABEL:fig_blockwise_heatmap showed an interesting phenomenon: for all the Transformers we checked, the blockwise Hessian spectra are largely different across blocks. In the following, we refer to this phenomenon as “block heterogeneity". As a control group experiment, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed. We refer to this phenomenon as “block homogeneity". These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers. Intuitively, the block-homogeneous Hessian in a CNN comes from the repetitively similar convolution layers, while the block-heterogeneous Hessian in a Transformer stems from the interleaved stacking disparate layers such as Query, Value, and MLPs. In the following, we will show that the block heterogeneity is strongly correlated with the failure of SGD on Transformers.
2.4 SGD Fails on Simple MLPs with Block Heterogeneity
We now further link block heterogeneity to the failure of SGD. We construct a simple example without attention layers but has block heterogeneity and find that SGD still fails. We consider a 4-layer MLP on MNIST. Using standard MLP, we find that the Hessian exhibits a homogeneous block structure and SGD converges as fast as Adam. We then scale the output of each layer by a constant $c$. We define $c$ as the degree of heterogeneity since larger $c$ would bring more heterogeneity to the Hessian (see Table 3 in the Appendix). As shown in Figure LABEL:fig_mlp_gap, we find that the gap between Adam and SGD increases with the degree of heterogeneity. This example shows that attention might not be the root cause of SGD failure, but the block heterogeneity (brought by attention) might be.
2.5 Implication on Choosing SGD or Adam
The above findings could bring up an interesting empirical guidance: we can compute the blockwise spectrum of initial Hessian, and then decide whether to use Adam or SGD. Such a method could be useful in scenarios in training large models that are not mainstream Transformers or CNNs, e.g., Mamba (Gu and Dao, 2023). In these cases, there is not much prior experience in choosing optimizers. It would be intriguing to decide whether SGD is suitable for the task before the training is launched. We present two reasons.
  • In large-scale model training, memory is limited but the model size grows, so Adam could become a major computational burden due to the memory of 2nd-order momentum (as mentioned in Section 1 and (Rasley et al., 2020)).
  • SGD is a lightweight alternative to Adam, but the performance of SGD may be bad. In large NN training, it might take days (or even weeks) to obtain informative feedback from training logs, so it is expensive to conduct trial and error to choose an appropriate optimizer. Specifically, if SGD performs badly, it would be difficult to judge whether the bad performance is due to improper learning rates or other more fundamental reasons. Then, it would be hard to decide whether to switch the optimizer.
Motivated by the above reasons, we here propose a quantitative metric that could predict the failure of SGD before the training is launched. With the help of this metric, we could save much expense on the trial and error for SGD. The metric is simply the averaged JS distance among blockwise Hessian spectra at initialization, i.e., the averaged value in each heatmap of Figure LABEL:fig_blockwise_heatmap. We denote it as $JS^{0}$. We present $JS^{0}$ of various models in Table 1. We note that $JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs. As argued in the previous sections, this quantity can predict the failure of SGD. Note that $JS^{0}$ is independent of optimizers and could be efficiently checked in priori before training.
Model ResNet18 VGG16 ViT-base BERT GPT2-nano GPT2
$JS^{0}$ 0.10 0.09 286.41 53.38 14.93 83.23
Table 1: $JS^{0}$ denotes the average JS distance between the initial Hessian spectra of each pair of parameter blocks. We find that the $JS^{0}$ on CNNs is at least 140×140$\times$140 × smaller than on Transformers. $JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs.
To summarize this section, we provide an explanation for SGD’s failure. We identify a phenomenon called block heterogeneity. This phenomenon naturally exists in Transformers, but not in CNNs. This phenomenon can exist in non-Transformer architectures as well. Our experiments suggest a strong connection between the failure of SGD and block heterogeneity. Our findings could provide empirical guidance to decide whether to use Adam or SGD.
3 Case Study of Quadratic Models and Preliminary Theory
Now we construct quadratic functions with block diagonal Hessian. We will compare the performance of GD and Adam on Hessian with and without block heterogeneity. Initial theoretical results on these quadratic models will be provided. Note that insights on quadratic models could be important for understanding realistic NNs, as mentioned by researchers such as LeCun et al. (2002) and OpenAI team (Kaplan et al., 2020).
3.1 Settings and Additional Notations
We consider $\min_{w}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$. We denote $\mathcal{L}^{*}$ as the minimum value of $\mathcal{L}(w)$. We set $H$ as a block diagonal matrix: $H=\operatorname{diag}(H_{1},\cdots,H_{L})$, where $H_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ and $d=\sum_{l=1}^{L}d_{l}$. We use $w_{l}\in\mathbb{R}^{d_{l}}$ to denote the variable in the $l$-th block and $w=(w_{1}^{T},\cdots,w_{L}^{T})^{T}\in\mathbb{R}^{d}$. Similarly for $h_{l}\in\mathbb{R}^{d_{l}}$. Similarly, we use $[\nabla L(w)]_{l}\in\mathbb{R}^{d_{l}}$ to denote the gradient in the $l$-th block and denote $[\mathcal{L}(w)]_{l}=\frac{1}{2}(w_{l}^{t})^{T}H_{l}w_{l}^{t}-h_{l}^{T}w_{l}$ as the objective function w.r.t. the $l$-th block. Note that $\mathcal{L}(w)=\sum_{l=1}^{L}[\mathcal{L}(w)]_{l}$. We denote $\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}$ as the eigenvalues of $H$. Similarly for $\lambda_{l,1}\cdots\lambda_{l,d_{l}}$. We denote $\kappa=\frac{\lambda_{1}}{\lambda_{d}}$ and $\kappa_{l}=\frac{\lambda_{l,1}}{\lambda_{l,d_{l}}}$ as the condition number of $H$ and $H_{l}$, respectively. We say an algorithm has complexity $\tilde{\mathcal{O}}(C)$ if it takes $\mathcal{O}(C\log(1/\epsilon))$ iterations to achieve error $\frac{\mathcal{L}(w)-\mathcal{L}^{*}}{\mathcal{L}\left(w^{0}\right)-\mathcal{L}^{*}}\leq\epsilon$, where $w^{0}$ is the initial point.
3.2 Experimental Observations
For all experiments, we choose $h=0$. We consider four types of Hessian $H$ as follows.
  • Case 1: Hessian with Transformer-type spectra. We choose $L=4$ and $d_{l}=100$. For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are independent standard Gassian random matrix and $\Lambda_{l}$ are diagonal matrices. For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the embedding layer; 3rd Query, 3rd Value, 3rd MLP (fc layer) in GPT2. Shifting and proportional scaling are performed to ensure all elements in $\Lambda_{l}$ lie in the interval $[1,2000]$. This ensures strong convexity and controls the condition number of $H$ equals $2000$. The spectra of $H_{l}$ are in Figure LABEL:fig:heter-block.
  • Case 2: Hessian with CNN-type spectra. We consider the same setup as Case 1. For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the 1st to 4th convolution layers in ResNet18. We shift and proportionally scale the elements $\Lambda_{l}$ to the interval $[1,2000]$ to ensure the strong convexity and the condition number $H$ equals 2000. The spectra of $H_{l}$ are shown in Figure LABEL:fig:homo-block.
  • Case 3: Hessian with simplified heterogeneous spectra. We choose $L=3$ and $d_{l}=3$. For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are independent standard Gassian random matrix and $\Lambda_{l}$ are diagonal matrices. We set the diagonal elements of $\Lambda_{l}$ as $\{1,2,3\},\{99,100,101\},\{1998,1999,2000\}$ for $l=1,2,3$, respectively. The spectra of $H_{l}$ are different due to their different supports. The condition number of Hessian $H$ is $2000$.
  • Case 4: Hessian with simplified homogeneous spectra. We consider the same setup as Case 3. We set the diagonal elements of $\Lambda_{l}$ as $\{1,99,1998\},\{2,100,1999\},\{3,101,2000\}$ for $l=1,2,3$, respectively. The spectra of $H_{l}$ are similar. The condition number of Hessian $H$ is 2000.
Note that in all four cases, the Hessian share the same condition number. We study two types of optimizers: one that assigns a single learning rate for all blocks, and ones that assign different learning rates across blocks. The optimizers are introduced as follows.
  • Single-learning-rate optimizer. We study gradient descent (GD).
    $w^{t+1}=w^{t}-\eta\nabla\mathcal{L}(w)=w^{t}-\eta(Hw^{t}-h)$ (1)
    We use the optimal learning rate $\eta=\frac{2}{\mu+L}$ (Nesterov, 2013). We use a standard Gaussian random vector as initialization.
  • Coordinate-wise-learning-rate optimizer. We study Adam with a constant learning rate and with no bias correction for simplicity (Algorithm 3). We set $\beta_{1}=0$ to erase the effect of momentum. This helps us to focus on the effect of coordinate-wise learning rate (or the effect of diagonal preconditioning) in Adam. We use $\epsilon=0$. We consider $\beta_{2}=1$ and $\beta_{2}=0.999$, respectively. When $\beta_{2}=1$, Adam assigns coordinate-wise learning rates according to the initial gradient, but these learning rates are fixed along iteration. The update rule is as follows.
    $w^{t+1}=w^{t}-\eta(D_{Adam}^{0})^{-1}\nabla\mathcal{L}(w)=w^{t}-\eta(D_{Adam}^{0})^{-1}(Hw^{t}-h),$ (2)
    where $D_{Adam}^{0}=\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))^{\frac{1}{2}}$ and $\nabla\mathcal{L}(w^{0})=Hw^{0}-h$. When $\beta_{2}<1$, the coordinate-wise learning rates adaptively change along iteration. The update rule is as follows.
    $w^{t+1}=w^{t}-\eta(D_{Adam}^{t})^{-1}\nabla\mathcal{L}(w)=w^{t}-\eta(D_{Adam}^{t})^{-1}(Hw^{t}-h),$ (3)
    where $D_{Adam}^{t}=\operatorname{diag}\left((1-\beta_{2})\left(\sum_{k=1}^{t}\beta_{2}^{t-k}\nabla\mathcal{L}(w^{k})\circ\nabla\mathcal{L}(w^{k})\right)+\beta^{t}\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))\right)^{\frac{1}{2}}$ and $\nabla\mathcal{L}(w^{k})=Hw^{k}-h$. We grid search $\eta$ and use a standard Gaussian random vector as initialization.
    We remark that when $\beta_{2}<1$, Adam would repeatedly bounce among non-optimal points, causing non-convergence. This will be shown in Proposition 2.
Summary of experimental observations.
We now compare the performance of Adam and GD on these four types of Hessians. The results are shown in Figure LABEL:fig_quadratic. For Hessian with heterogeneous blocks like Case 1 and 3, GD largely lags behind Adam. For Hessian with homogeneous blocks (Case 2 and 4), GD is on par or even better than Adam. We emphasize that all Hessians have the same condition number in these four cases. Further, Hessian of Case 3 and 4 share all the eigenvalues (not just the extreme ones). The performance gap between Adam and GD is due to the different blockwise spectra caused by the different locations of eigenvalues. We hypothesize that GD performs badly because it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. Such heterogeneity can be better handled using different learning rates across blocks, as designed in Adam.
3.3 Initial Theoretical Results
We now provide initial theoretical results to characterize how GD lags behind Adam in problems with heterogenous Hessian. Note that classical optimization theory depicts the rate of first-order methods by the condition number of the full Hessian $\kappa$. However, we point out that $\kappa$ is not informative enough to describe the performance gap in Figure LABEL:fig_quadratic since $\kappa$ is the same in all four cases. To distinguish Adam and GD, we need to utilize more fine-grained quantities like blockwise spectra of sub-matrices.
Unfortunately, the blockwise spectrum is rarely discussed in the optimization area. A most related notion is “block Lipschitz constant" (Beck and Tetruashvili, 2013) for studying block coordinate descent (BCD) type methods, but it was not linked to the performance of SGD or Adam before. To our knowledge, we are not aware of any theory of Adam or GD built on the block diagonal structures or the blockwise spectra of Hessian. Further, there seems to be no result distinguishing the complexity of single-learning-rate methods and the coordinate-wise counterpart, either. We now make an initial attempt in this direction. We first present the lower bound for GD.
Proposition 1.
(Lower bound for GD.) Consider $\min_{w}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$. Let $w_{GD}^{t}$ be the output of GD after $t$ steps. There exists a block diagonal matrix $H$, $h$ and an initial point $w^{0}$, s.t., for any $\eta$, we have:
$\mathcal{L}(w_{GD}^{t+1})-\mathcal{L}^{*}\geq\left(1-\frac{2}{\kappa+1}\right)\left(\mathcal{L}(w_{GD}^{t})-\mathcal{L}^{*}\right)$ (4)
where $\kappa$ is the condition number of $H$.
Proposition 1 shows that GD has complexity $\tilde{\mathcal{O}}(\kappa)$ and such complexity is tight. Now we prove that: if we are allowed to choose a customized learning rate for each block sub-matrix $H_{l}$, then the complexity will be improved. We call this method blockwise GD (BGD). Note that BGD is impractical to implement since it is hard to choose $L$ learning rates in advance. BGD serves as an artificial method for theoretical understanding.
Blockwise GD (BGD).
Choose learning rates $\{\eta_{l}\}_{l=1}^{L}$ for all blocks in advance and performs the following:
$w^{t+1}=w^{t}-(D_{BGD})^{-1}\nabla\mathcal{L}(w)=w^{t}-(D_{BGD})^{-1}(Hw^{t}-h),$
where $D_{BGD}=\operatorname{diag}\left(\left(\underbrace{\eta_{1},\cdots,\eta_{1}}_{d_{1}\text{times}},\cdots,\underbrace{\eta_{L},\cdots,\eta_{L}}_{d_{L}\text{times}}\right)^{T}\right)\in\mathbb{R}^{d\times d}.$
BGD can be also viewed as a diagonally preconditioned version of GD. We now show that this diagonal preconditioner can significantly improve the complexity of GD.
Theorem 1.
(Upper bound for GD with blockwise learning rate.) Consider $\min_{w}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$. We assume $H$ is a block diagonal matrix, i.e., $H=\operatorname{diag}(H_{1},\cdots,H_{L})$, where $H_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ and $d=\sum_{l=1}^{L}d_{l}$. Let $w_{BGD}^{t}$ be the output of BGD after $t$ steps. There exists a set of $\{\eta_{l}\}_{l=1}^{L}$ such that:
$\mathcal{L}(w_{BGD}^{t+1})-\mathcal{L}^{*}\leq\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\left(\mathcal{L}(w_{BGD}^{t})-\mathcal{L}^{*}\right)$ (5)
where $\kappa_{l}$ is the condition number of $H_{l}$.
Theorem 1 shows that: if we are allowed to choose a learning rate for each $H_{l}$, the complexity could be $\tilde{O}(\max_{l\in[L]}\kappa_{l})$, in lieu of $\tilde{O}(\kappa)$ as in GD. That is to say, the rate of BGD only depends on the "slowest block". Note that in certain heterogeneous block Hessian such Case 1 and ${\bf 3}$, $\kappa_{l}$ is about 20×20$\times$20 × and 2000×2000$\times$2000 × smaller than $\kappa$, respectively. In these cases, blockwise-learning-rate methods can be about 20×20$\times$20 × and 2000×2000$\times$2000 × faster than single-learning-rate methods like GD, respectively.
Despite the advantages of BGD, it is impractical to find the proper learning rates for all blocks, especially when the number of blocks $L$ is large. Fortunately, these learning rates could be approximated by the preconditioner in Adam, e.g., $D_{Adam}^{0}$ in (2). We now provide a complexity bound of Adam.
Assumption 1.
(Bounded initialization.) The initialization is bounded and is far from the optimal solution:
$C_{l,1}\lambda_{l,1}\leq\left|[\nabla\mathcal{L}(w^{0})]_{l,i}\right|\leq C_{l,2}\lambda_{l,1},\text{ for }i\in[d_{l}],l\in[L].$
Theorem 2.
(Upper bound for Adam with $\beta_{2}=1$.) Consider the same setting as in Theorem 1 and consider Adam with $\beta_{1}=0$ and $\beta_{2}=1$ as in (2). Assume the initialization satisfies Assumption 1. Let $w_{Adam}^{t}$ be the output of Adam after $t$ steps. Let $\eta=\min_{l\in[L]}\frac{1}{C_{l,1}}$. We have
$\mathcal{L}(w_{Adam}^{t+1})-\mathcal{L}^{*}\leq\max_{l\in[L]}\left(1-\frac{1}{\kappa_{Adam,l}}\right)\left(\mathcal{L}(w_{Adam}^{t})-\mathcal{L}^{*}\right)$ (6)
where $\kappa_{Adam,l}=r\kappa_{l}$, $r:=\max_{l\in[L]}\left(\frac{C_{l,2}}{C_{l,1}}\right)^{2}$ and $\kappa_{l}$ is the condition number of $H_{l}$.
The proofs of all the above theorems are shown in Appendix E. Theorem 2 states that Adam (with $\beta_{2}=1$) has complexity $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$. We remark that condition $\beta_{2}=1$ is necessary because any $\beta_{2}<1$ causes non-convergence issue (Da Silva and Gazeau, 2020; Bock and Weiß, 2019). We restate their results in Proposition 2. The non-convergence is also observed in Figure LABEL:fig_quadratic (c), where we find that the iterates of Adam quickly converge to near-optimal solutions, and then bounce back. As such, $\beta_{2}=1$ is necessary for asymptotic analysis. As shown in (Da Silva and Gazeau, 2020), the non-convergence is due to the constant learning rate.
Proposition 2.
(Non-convergence of constant-learning-rate Adam with $\beta_{2}<1$.) (Da Silva and Gazeau, 2020, Proposition 12) Consider $\min_{w\in\mathbb{R}}\mathcal{L}(w)=\frac{1}{2}w^{2}$. Consider Adam with $\beta_{1}=0$ and $\beta_{2}<1$ as in (3). Let $w_{Adam}^{t}$ be the output of Adam after $t$ steps. There exists a discrete limit cycle for (3), i.e., Adam produces oscillations and ${\small\liminf_{t\rightarrow\infty}\left(\mathcal{L}(w_{Adam}^{t})-\mathcal{L}^{*}\right)>0}$.
We now compare the complexity of Adam and that of GD. By Theorem 2, Adam is faster than GD when $r\cdot\max_{l\in[L]}\kappa_{l}\leq\kappa$. In the quadratic model with heterogeneous blocks (Case 3), our simulation over 1000 trials shows that $r\leq 100$ with probability $\geq\frac{2}{3}$ when using standard Gaussian random initialization. Since $\max_{l\in[L]}\kappa_{l}\approx 1$, we have $r\cdot\max_{l\in[L]}\kappa_{l}\leq 100$, w.h.p., and is 20×20$\times$20 × smaller than $\kappa=2000$. So Adam could be 20×20$\times$20 × faster than GD, w.h.p.. This is indeed observed in Figure LABEL:fig_quadratic where Adam outperforms GD by a significant margin. We summarize the complexity of GD, BGD, and Adam in Table 2.
Optimizer GD BGD Adam with Adam with
$\beta_{1}=0$ and $\beta_{2}=1$ (2) $\beta_{1}=0$ and $\beta_{2}<1$ (3)
Complexity $\tilde{O}(\kappa)$ $\tilde{O}(\max_{l\in[L]}\kappa_{l})$ $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$
Table 2: The complexity of GD, BGD and Adam for minimizing a strongly convex quadratic function with block diagonal Hessian. The symbol ✗ means non-convergence. $\kappa$ and $\kappa_{l}$ denote the condition number of the full Hessian and the block submatrix, respectively. $r:=\max_{l\in[L]}\left(\frac{C_{l,2}}{C_{l,1}}\right)^{2}$ and $C_{l,1}$, $C_{l,2}$ are constants in Assumption 1.
How to obtain a tighter complexity bound of Adam?
It is valid to ask whether the complexity upper bound $\kappa_{Adam,l}=r\kappa_{l}$ can be tightened, e.g., remove the factor of $r$. We point out it would be difficult if there is no extra structure on $H_{l}$. A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$. Intuitively, a diagonal preconditioner of $H_{l}$ is expected to be powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant. Unfortunately, it is unclear whether these structures hold in real Transformers. Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number. For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $3.9\kappa_{1}$, $14.1\kappa_{2}$, $4.2\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials). It would be interesting to explore if there are special structures of $H_{l}$ in real Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it. We leave it as a future direction.
Although Adam preconditioner might not always reduce the “local" condition number $\kappa_{l}$, the coefficient in the complexity is now independent of the “global" condition number $\kappa$. As argued above, such changes in coefficient could lead to considerable improvement over GD. Such improvement in complexity is attributed to the block diagonal structure in Hessian as well as its heterogeneous blockwise spectrum. To our knowledge, such improvement is not shown in the existing literature. In summary, our theory indicates that: for problems with block heterogeneity, the single-learning rate methods like GD can largely lag behind coordinate-wise learning rate methods like Adam.
4 Related Works
On The failure of SGD on Transformers
There is extensive literature studying the difficulties of Transformer training. We summarize these works in Appendix A. We here focus on the literature that explores why SGD fails on Transformers. One representative hypothesis is that SGD fails because it cannot handle the heavy-tailed stochastic noise in language tasks (Zhang et al., 2020). However, Chen et al. (2021); Kunstner et al. (2023) reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons. Further, SGD performs poorly on Vision Transformers on ImageNet (Figure LABEL:fig:nlp_figure), so the data modality (e.g., language or image tasks) might not be as crucial as the architecture. Zhang et al. (2019c) showed that Transformers have the “unbounded smoothness" issue and SGD with gradient clipping performs better than SGD in this case. Although clipping is an effective trick, it does not save SGD from failure as we still observe a huge gap between clipped SGD and Adam  [^3] For all NLP tasks, clipping is performed after backpropagation. So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD. . So there might be other reasons that hamper SGD. Different from these works, we find SGD fails because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
Understanding of Adam.
There was once a long-standing debate on the possible divergence of Adam (Reddi et al., 2018). The convergence for the unmodified versions is later established in (Shi et al., 2020; Zhang et al., 2022b) for RMSprop and Adam. More convergence analyses of general adaptive gradient methods are listed in Appendix A. We here focus on the literature that explores the benefit of Adam. Xie et al. (2022) show that Adam can help avoid saddle points, which is an orthogonal direction to this work. Wang et al. (2022a); Crawshaw et al. (2022); Li et al. (2023) show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time). We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes fails while Adam works well (Figure LABEL:fig_quadratic). This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate over time, and there are other reasons for Adam’s success. We show that an important benefit of Adam is its ability to handle the heterogeneity across blocks (over space).
Bernstein et al. (2018); Wu et al. (2020); Kunstner et al. (2023); Liu et al. (2023) build a relation between Adam and the sign-based methods. Wu et al. (2020) further showed that sign-based methods can be effective when the Hessian is diagonal and satisfies several other properties. However, as put by the authors, it seems “unclear to what extent these properties hold for real problems". Pan and Li (2023) numerically found that the Adam can reduce the directional sharpness along trajectories, while its relation to fast convergence remains mysterious. A recent work (Jiang et al., 2023) point out that Adam biases the trajectories towards regions where Hessian has “uniform diagonal entries" while SGD cannot. The distribution of Hessian diagonal entries is also investigated in (Liu et al., 2023). The theory in (Jiang et al., 2023) implies that Adam is faster when the Hessian is diagonal. However, as argued above, it is unclear whether the diagonal Hessian commonly holds in real problems. In fact, we find the Hessian is closer to a block-diagonal (instead of pure diagonal) structure on some small Transformers. In these cases, blockwise eigenvalues carry more information than diagonal entries, providing extra details such as the location of eigenvalues. We find that these extra details are important for distinguishing Adam and SGD.
Hessian Spectrum Analysis.
There are several important attempts to explore the Hessian spectrum of MLPs and CNNs. Sagun et al. (2016, 2017); Chaudhari et al. (2019) found that the Hessian spectra of MLPs and CNNs consist of a “bulk" together with a few “outliers". Papyan (2020); Wu et al. (2020); Liao and Mahoney (2021) further characterized the bulks and outliers in theory. Papyan (2018, 2019) numerically built the relation between these "outliers" and the Gauss-Newton matrix. Sankar et al. (2021) numerically explored the relation between Hessian of CNNs and Gauss-Newton matrix in each layer. They further showed that most CNN layers contribute similarly to the overall loss surface. We find that this result is restricted to CNNs and does not hold on Transformers due to the heterogeneity. Gur-Ari et al. (2018) showed that for MLPs and CNNs, gradient descent converges to a small subspace spanned by a few top eigenvectors of the Hessian. Yao et al. (2018); Zhang et al. (2019b) explored the relation between the Hessian spectrum of CNNs and some training phenomena such as the effect of batch sizes. Ghorbani et al. (2019); Yao et al. (2020) focused on explaining the effectiveness of techniques such as BatchNorm. Note that all these works are restricted to MLPs and CNNs, while we study the Hessian of Transformers (in addition to CNNs and MLPs) as well as its impacts on different optimizers.
5 Conclusion
In this work, we investigate the failure mode of SGD on Transformers through the lens of Hessian. By numerically explore various Transformers, CNNs, MLPs and quadratic problems, we establish a phenomenon called block heterogeneity in Hessian and link it to the failure of SGD. We point out that SGD fails because it applies one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. Initial theory are provided to support the claim.
Appendix A More Related Works
On the difficulties of Transformer training.
Transformers are known to be difficult to train. Researchers have attributed the training difficulties to various phenomena in different components of Transformers, including: the logits divergence or the rank degeneracy in the outputs of attention layers [Dong et al., 2021, Noci et al., 2022, Wortsman et al., 2023, Zhai et al., 2023, Dehghani et al., 2023, Chowdhery et al., 2023]; the growth of parameter norm in attention layers [Merrill et al., 2020]; over-reliance on residue branches [Liu et al., 2020]; and some negative impact of layer norm [Chen et al., 2018, Zhang et al., 2019a, Huang et al., 2020]. These phenomena have a strong correlation with gradient vanishing or explosion in Transformers [Zhang et al., 2019a, Liu et al., 2020, Huang et al., 2020, Xiong et al., 2020, Noci et al., 2022, Wang et al., 2022b, Wortsman et al., 2023, Molybog et al., 2023], which leads to training difficulties.
Several solutions have been proposed. Liu et al. [2020] numerically observed that adaptive gradient methods can (partly) overcome gradient vanishing by giving “consistent update magnitude", while it seems unclear how consistent update magnitude would help optimization in principle. Researchers further develop training tricks such as warmup learning rate [Liu et al., 2019, Xiong et al., 2020], temperature scaling [Noci et al., 2022], better initialization [Zhang et al., 2019a, Huang et al., 2020, Wang et al., 2022b, Bachlechner et al., 2021, Yang et al., 2022], and variants of Layer Norm [Nguyen and Salazar, 2019, Wang et al., 2019, Xiong et al., 2020, Wang et al., 2022b, Dehghani et al., 2023]. Recent researchers also suggest using z-loss regularization [Chowdhery et al., 2023, Yang et al., 2023] and tuning hyperparameters of Adam [Zhang et al., 2022b, Wortsman et al., 2023]. All these tricks can help mitigate gradient explosion or vanishing. Nevertheless, training large-scale Transformers remains challenging [Zhang et al., 2022a, Zeng et al., 2022, Wortsman et al., 2023, Molybog et al., 2023, Chowdhery et al., 2023]. Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian. We establish a strong correlation between the blockwise Hessian spectra of Transformers and the failure of SGD. We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
Convergence analysis of adaptive gradient methods
There is extensive convergence analysis for adaptive gradient methods. For instance, researchers study the convergence of AMSGrad [Reddi et al., 2018, Zhou et al., 2018], RMSprop [Zaheer et al., 2018], AdaFom [Chen et al., 2019], AdaBound [Luo et al., 2018], and Adam with iterate-dependent hyperparameters [Zou et al., 2019, Chen et al., 2022, Gadat and Gavra, 2022]. The convergence of Adam is also explored in [Défossez et al., 2022, Wang et al., 2023a]. There is also an active line of theoretical research on the convergence of AdaGrad [Duchi et al., 2011], we recommend [Wang et al., 2023b] for more detailed introduction. In this work, we do not focus on the convergence analysis. Rather, we explore the quantitative difference between the loss landscape of CNNs and Transformers and how it impact the behaviors of SGD and Adam.
Appendix B More Preliminaries
B.1 Preliminaries on Optimizers
[…]
Algorithm 1 Stochastic Gradient Descent with Momentum (SGD)
1:  Initialize $w^{0}$ and choose $10\leq\beta_{1}<$ 1 and $\eta_{0}>0$
2:  for $t=1\to\infty$ do
3:     Uniformly sample $\tau^{t}$ from the index set $\{1,2,\cdots,n\}$
4:     $m^{t}=\beta_{1}m^{t}+\nabla f_{\tau^{t}}(x^{t})$
5:     $x^{t+1}=x^{t}-\eta_{t}m^{t}$
6:  end for
Algorithm 2 AdamW
1:  Initialize $x^{0}$, $m^{0}=v^{0}=0$, $10\leq\beta_{1}<$ 1, $10\leq\beta_{2}<$ 1, $\epsilon>0$, $\eta^{0}>0$, and weight decay coefficient $\lambda$
2:  for $t=1\to\infty$ do
3:     Uniformly sample $\tau^{t}$ from the index set $\{1,2,\cdots,n\}$
4:     $w^{t+1}=w^{t}-\eta^{t}\lambda w^{t}$
5:     $m^{t}=\beta_{1}m^{t}+(1-\beta_{1})\nabla\mathcal{L}_{\tau^{t}}(w^{t})$
6:     $v^{t}=\beta_{2}v^{t}+(1-\beta_{2})\nabla\mathcal{L}_{\tau^{t}}(w^{t})\circ\nabla\mathcal{L}_{\tau^{t}}(w^{t})$
7:     $\hat{m}^{t}=\frac{m^{t}}{1-(\beta_{1})^{t}}$
8:     $\hat{v}^{t}=\frac{v^{t}}{1-(\beta_{2})^{t}}$
9:     $w^{t+1}=w^{t+1}-\eta_{t}\frac{\hat{m}^{t}}{\sqrt{\hat{v}^{t}}+\epsilon}$
10:  end for
Algorithm 3 Adam with no bias correction
1:  Initialize $x^{0}$, $m^{0}=\nabla\mathcal{L}_{\tau^{t}}(w^{0}),v^{0}=\nabla\mathcal{L}_{\tau^{t}}(w^{0})\circ\nabla\mathcal{L}_{\tau^{t}}(w^{0})$, $10\leq\beta_{1}<$ 1, $10\leq\beta_{2}<$ 1, $\epsilon>0$, $\eta^{0}>0$
2:  for $t=1\to\infty$ do
3:     Uniformly sample $\tau^{t}$ from the index set $\{1,2,\cdots,n\}$
4:     $m^{t}=\beta_{1}m^{t}+(1-\beta_{1})\nabla\mathcal{L}_{\tau^{t}}(w^{t})$
5:     $v^{t}=\beta_{2}v^{t}+(1-\beta_{2})\nabla\mathcal{L}_{\tau^{t}}(w^{t})\circ\nabla\mathcal{L}_{\tau^{t}}(w^{t})$
6:     $w^{t+1}=w^{t+1}-\eta_{t}\frac{m^{t}}{\sqrt{v^{t}}+\epsilon}$
7:  end for
B.2 Preliminaries on the Stochastic Lanczos Quadrature Method
Additional notations.
Given a real symmetric matrix $H\in\mathbb{R}^{d\times d}$, we denote $tr(H)$ as its trace and $Q^{T}\Lambda Q$ as its spectral decomposition, where $Q=\left[q_{1},\ldots,q_{d}\right],\Lambda=\operatorname{diag}\left(\lambda_{1},\ldots,\lambda_{d}\right)$ and $\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}$. We denote the condition number of $H$ as $\kappa=\lambda_{1}/\lambda_{d}$. We define matrix function as $f(H):=Q^{T}f(\Lambda)Q$, where $f(\Lambda)=\operatorname{diag}\left(f\left(\lambda_{1}\right),\ldots f\left(\lambda_{d}\right)\right)\in\mathbb{R}^{d\times d}$. We use $\mathbb{N}$ to denote the set of positive integers. We use $\|\cdot\|_{2}$ to denote the Euclidean norm.
Approximation of the Hessian spectrum can be formulated as a trace estimation problem, as introduced in [Lin et al., 2016, Ubaru et al., 2017]. First, the spectrum (eigenvalue density) of Hessian $H$ can written as: $\phi(t)=\frac{1}{d}\sum_{i=1}^{d}\delta\left(t-\lambda_{i}\right)$, where $\lambda_{i}$ are the eigenvalues of $H$ and $\delta$ is the Dirac $\delta$-function. Then, we replace the delta functions by a Gaussian blurring function: $\phi(t)\approx g(t):=\frac{1}{d}\sum_{i=1}^{d}f\left(\lambda_{i}\right)$, where $f(\lambda):=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(t-\lambda)^{2}}{2\sigma^{2}}\right)$. By definition of matrix function, it is easy to see that $g(t)=\frac{1}{d}tr(f(H))$. As such, spectrum approximation could be formulated as a trace estimation problem, i.e., estimating $\frac{1}{d}\operatorname{tr}(f(H))$, where $H\in\mathbb{R}^{d\times d}$ is a real symmetric matrix.
Trace estimation problems could be solved efficiently by the Stochastic Lanczos Quadrature Method (SLQ) [Golub and Strakoš, 1994]. For the ease of readers, we provide a detailed description of SLQ in our context. SLQ consists of the following steps.
Step 1.
We Approximate the trace of matrix function as $\frac{1}{d}tr(f(H))=\mathbb{E}(v^{T}f(H)v)\approx\frac{1}{n_{v}}\sum_{i}^{n_{v}}v_{i}^{T}f(H)v_{i}$, where $v=u/\|u\|_{2}$ and $u$ is a Rademacher random vector (each entry of $u$ independently takes ±$\pm 1$ with probability $1/2$). This step is called Huchinson’s estimation [Hutchinson, 1989].
Note that we can also replace the Rademacher random vector $u$ by a unit Gaussian vector (i.e., $u\sim N(0,I_{d\times d})$) and the unbiasedness still holds [Avron and Toledo, 2011]. In our implementation, we sample $u\sim N(0,I_{d\times d})$ because there is an efficient built-in PyTorch function for generating Gaussian vectors.
SLQ estimates $v_{i}^{T}f(H)v_{i}$ for $i\in[n_{v}]$ and then take the average. To understand SLQ, we only need to understand how it estimates each individual quadratic form. To simplify the notation regarding $i$, from now on, we will discuss how to estimate $v^{T}f(H)v$, where $v=u/\|u\|_{2}$ and $u$ is a unit Gaussian vector.
Step 2-1.
We rewrite $v^{T}f(H)v$ as a Riemann-Stieltjes integral [Golub and Meurant, 2009]:
[…]
where $\mu$ is a measure on $(\mathbb{R},\mathbb{B})$ defined as follows ($\mu(\lambda)$ denotes the measure of set $\{x;x\leq\lambda\}$):
$\mu(\lambda)=\begin{cases}0&\lambda<\lambda_{d}\\ \sum_{i=1}^{k}\left(v^{T}q_{i}\right)^{2}&\lambda_{k}\leq\lambda<\lambda_{k+1}\\ \sum_{i=1}^{d}\left(v^{T}q_{i}\right)^{2}&\lambda\geq\lambda_{1}\end{cases}.$ (8)
Step 2-2.
Unfortunately, this integral is difficult to compute. This is because the measure $\mu$ are related to the eigen-pairs of $H$, which are unknown. It seems unclear how to directly integrate over an unknown measure. As such, we further approximate this integral by a computationally friendly quantity, such as:
[…]
We hope to design $\{(c_{j},x_{j})\}_{j=1}^{m}$ with a reasonable number of $m$ such that the estimation error is small. Fortunately, the Gaussian Quadrature method provides a generic design principle of $\{(c_{j},x_{j})\}_{j=1}^{m}$ [Golub and Meurant, 2009, Epperson, 2013]. It is proved that: when $f(\lambda)$ is not "too complicated" (e.g. $f(\lambda)$ is a polynomial), then there exists $\{(c_{j},x_{j})\}_{j=1}^{m}$ which gives a high quality estimation of integral (7). The required number of $m$ is related to "how complicated the $f(\lambda)$ is". Such $\{(c_{j},x_{j})\}_{j=1}^{m}$ are called the Gaussian Quadrature rules. $c_{j}$ and $x_{j}$ are called the "weights" and the "nodes" of the Gaussian Quadrature rules. A representative theorem is as follows: when $f(\lambda)$ is a polynomial with degree $<2m$, then the Gaussian Quadrature rules give the exact approximation of integral (7).
Theorem 3.
[…]
Theorem 3 shows the existence of good $\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ and their general form. In fact, it is also shown that Gaussian Quadrature is optimal: no other$\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ can achieve zero approximation error for higher degree polynomials $f(\lambda)$ [Golub and Meurant, 2009]. However, it is often difficult to find these quadrature rules [Golub and Welsch, 1969]. There are at least three questions in sequel:
  • 1) how to find the orthogonal polynomials $\left\{p_{k}(x)\right\}_{k=1}^{m}$ w.r.t. an unknown measure $\mu$?
  • 2) how to efficiently find the roots of $p_{m}(x)$, which gives the nodes $x_{j}$?
  • 3) how to efficiently calculate the weights $c_{j}=\int_{\lambda_{d}}^{\lambda_{1}}\prod_{j\neq i}\frac{x-x_{i}}{x_{j}-x_{i}}d\mu$?
We first answer question 2) and 3) and leave question 1) for later discussion.
Now suppose that we have found the orthogonal polynomials $\left\{p_{k}(x)\right\}_{k=1}^{m}$ w.r.t. $\mu$. Recall that any orthogonal polynomial has the following "three-term" recursion [Golub and Meurant, 2009]:
$p_{k+1}(x)=\left(x-\alpha_{k+1}\right)p_{k}(x)-\beta_{k}p_{k-1}(x),k=0,1,\ldots,$
where $p_{-1}(x)\equiv 0,p_{0}(x)\equiv 1,\alpha_{k+1}=\frac{\left\langle xp_{k},p_{k}\right\rangle}{\left\langle p_{k},p_{k}\right\rangle}$ and $\beta_{k}=\frac{\left\langle p_{k},p_{k}\right\rangle}{\left\langle p_{k-1},p_{k-1}\right\rangle}$. Define $P_{m}(x)=\left(p_{0}(x),p_{1}(x),\ldots p_{m-1}(x)\right)^{T}\in\mathbb{R}^{m}$, we can rewrite the recursion formula in matrix form (given $x$): $xP_{m}=J_{m}P_{m}+\beta_{m}p_{m}(x)e^{m}$, where $e^{m}$ is the last column of identity matrix $I_{m,m}$ and $J_{m}$ is called Jacobi matrix of order $m$:
$J_{m}=\left(\begin{array}[]{ccccc}\alpha_{1}&\sqrt{\beta_{1}}&&&\\ \sqrt{\beta_{1}}&\alpha_{2}&\sqrt{\beta_{2}}&&\\ &\sqrt{\beta_{2}}&\alpha_{3}&\sqrt{\beta_{3}}&\\ &&\ddots&\ddots&\ddots\end{array}\right)\in\mathbb{R}^{m\times m}$
It turns out that $J_{m}$ can help us find the Gaussian Quadrature rules $\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ and thus provide answers for question 2) and 3). This is shown in the following theorem.
Theorem 4.
[Golub and Meurant, 2009] For the Gaussian Quadrature, $\{x_{j}\}_{j=1}^{m}$ are the eigenvalues of $J_{m}$ and $\{c_{j}\}_{j=1}^{m}$ are the squares of the first elements of the normalized eigenvectors of $J_{m}$.
The proof of Theorem 4 is based on Christoffel-Darboux relation [Brezinski, 1990]. Now, the remaining question is: how to find the Jacobian matrix $J_{m}$ of a sequence of orthogonal polynomials w.r.t. an unknown measure $\mu$? Note that we no longer need to answer question 1) if $J_{m}$ is found, since $J_{m}$ is sufficient for us to find the Gaussian quadrature rules. However, it seems impossible to find $J_{m}$ if no information of $\mu$ is provided. The good news is: when the $\mu$ is specified as in (8), there exists an efficient way to find $J_{m}$.
Step 3.
When $\mu$ is specified as in (8), $J_{m}$ can be exactly found in $m$ steps using the Lanczos algorithm [Lanczos, 1950], as shown in Algorithm 4. This method takes a real symmetric matrix as input and returns a tridiagonal matrix. It was originally proposed to solve eigenvalue problems. Later, researchers found a deep connection between the Lanczos algorithm and orthogonal polynomials, which further connects this method to the Gaussian quadrature. The method (of finding the Gaussian quadrature by the Lanczos algorithm) is called the Lanczos quadrature [Golub and Strakoš, 1994, Bai and Golub, 1996, Golub and Meurant, 2009]. An extremely elegant but highly nontrivial result is as follows:
Theorem 5.
[Golub and Meurant, 2009] Given a real symmetric matrix $H\in\mathbb{R}^{d\times d}$ and an arbitrary vector $v\in\mathbb{R}^{d}$ with unit Euclidean norm, we define the measure $\mu$ as in (8) based on this $H$ and $v$. Then $m$ steps of the Lanzcos algorithm return the Jacobian matrix $J_{m}$ of orthogonal polynomials w.r.t. to $\mu$.
After $J_{m}$ is found by the Lanczos algorithm, we perform spectral decomposition of $J_{m}\in\mathbb{R}^{m\times m}$ to get its eigen-pairs. Using Theorem 4, we successfully get the Gaussian quadrature rules and thus we can approximate the quadratic form $v^{T}f(H)v$. By averaging over different random vectors $v$ we can then approximate $\frac{1}{d}tr(f(H))$. This concludes the derivation of SLQ for the trace estimation problem.
The full procedure of SLQ is shown in Algorithm 5. We note that SLQ is efficient in theory. Ubaru et al. [2017] show that SLQ converges faster than any other polynomial expansion method for spectrum estimation (e.g., Chebyshev methods used in [Adams et al., 2018]). See [Ubaru et al., 2017, Theorem 4.1] for a formal statement.
We remark that there are at least four versions of the Lanczos algorithm in Step 3. Here, we adopt the version in Algorithm 4 since it is known to be the most numerically stable version [Cullum and Willoughby, 2002, Saad, 2011, Wikipedia, 2023]. Throughout this work, we choose $f\left(\cdot\right)$ as the Gaussian blurring function $f(\lambda):=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(t-\lambda)^{2}}{2\sigma^{2}}\right)$ for spectrum approximation. We plot the spectrum by sweeping $t$ from the minimal node to the maximal node in Gaussian Quadrature rules.
Algorithm 4 The Lanczos Algorithm
1:  Input a matrix-vector product $Hv_{1}\in\mathbb{R}^{d}$, where $H$ is a real symmetric matrix and $v_{1}$ is an arbitrary vector with Euclidean norm 1. Choose $m\in\mathbb{N}$
2:  Initialization: Let $w_{1}^{\prime}=Hv_{1}$, $\alpha_{1}=(w_{1}^{\prime})^{T}v_{1}$, $w_{1}=w_{1}^{\prime}-\alpha_{1}v_{1}$
3:  for $j=2\to m$ do
4:     Let $\beta_{j}=\left\|w_{j-1}\right\|_{2}$ (also Euclidean norm)
5:     If $\beta_{j}\neq 0$, then let $v_{j}=w_{j-1}/\beta_{j}$, else pick as $v_{j}$ an arbitrary vector with Euclidean norm 1 that is orthogonal to all of $v_{1},\ldots,v_{j-1}$
6:     Let $w_{j}^{\prime}=Av_{j}$
7:     Let $\alpha_{j}=(w_{j}^{\prime})^{T}v_{j}$
8:     Let $w_{j}=w_{j}^{\prime}-\alpha_{j}v_{j}-\beta_{j}v_{j-1}$
9:  end for
10:  Let $V$ be the matrix with columns $v_{1},\ldots,v_{m}$
11:  Let $T=\left(\begin{array}[]{cccccc}\alpha_{1}&\beta_{2}&&&&0\\ \beta_{2}&\alpha_{2}&\beta_{3}&&&\\ &\beta_{3}&\alpha_{3}&\ddots&&\\ &&\ddots&\ddots&\beta_{m-1}&\\ &&&\beta_{m-1}&\alpha_{m-1}&\beta_{m}\\ 0&&&&\beta_{m}&\alpha_{m}\end{array}\right)$
12:  Return $T$
Algorithm 5 The Stochastic Lanczos Quadrature Method
1:  Choose $\texttt{num}_{v},m\in\mathbb{N}$. Sample $\texttt{num}_{v}$ i.i.d. $v_{i}$ from normalized Rademacher distribution, $i\in[\texttt{num}_{v}]$
2:  for $i=1\to\texttt{num}_{v}$ do
3:     Run $m$ steps of the Lanczos Algorithm 4 with input $Hv_{i}$, returns $T\in\mathbb{R}^{m\times m}$
4:     Compute eigenvalue decomposition $T=Q\Lambda Q^{T}$
5:     Compute the nodes $x_{i}=\left(\Lambda_{ii}\right)_{i=1}^{m}$ and weights $c_{i}=\left(Q_{1,i}^{2}\right)_{i=1}^{m}$
6:     Return $q_{i}(t)=\sum_{i=1}^{m}c_{i}f\left(x_{i};t,\sigma^{2}\right)$
7:  end for
8:  Return $\frac{1}{\texttt{num}_{v}}\sum_{i=1}^{\texttt{num}_{v}}f\left(\ell_{i};t,\sigma^{2}\right)$
Appendix C More Eperimental Details
C.1 Implementation Details on SLQ
Implementation and Running Time Analysis.
We provide a simple PyTorch implementation of SLQ. The only query SLQ makes to the neural network is the Hessian vector product, which is attained using the auto-differentiation framework [Pearlmutter, 1994]. To assure the accuracy of the Lanczos algorithm, we remove all the randomness in the forward and backward passes, including: data shuffling order, data augmentation, and dropout, etc.. Since Flash Attention [Dao et al., 2022] does not support the calculation of Hessian-vector product, we implement all attention blocks in the naive way. For the calculation of the blockwise Hessian spectrum $\nabla^{2}\mathcal{L}(w_{l})$, we sample $u_{l}\sim N(0,I_{d_{l}\times d_{l}})$ and set $v_{l}=u_{l}/\|u_{l}\|_{2}\in\mathbb{R}^{d_{l}}$. Then we run Algorithm 5 by taking $\nabla^{2}\mathcal{L}(w_{l})$ and $v_{l}$ as inputs. We choose the hyperparameters as $m=100$ and $n_{v}=10$ in all experiments. $\sigma$ is tuned based on visual effects. These hyperparameters are reported to reach highly accurate estimation with error $<10^{-14}$ [Ghorbani et al., 2019].
We now briefly discuss the computational cost of SLQ. The major computational expense of SLQ is the repeated Hessian-vector product operations in Lanczos algorithm in Step 3. Recall $\nabla^{2}\mathcal{L}(w)d=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\mathcal{L}_{i}(w)d$, so each Hessian-vector product operation requires (i) calculating $\nabla^{2}\mathcal{L}_{i}(w)d$; (ii) repeating (i) on all data. We point out that (i) can be computed efficiently and precisely with just two backpropagation passes [Pearlmutter, 1994]. The major computational bottleneck lies in (ii) due to the large $n$. Our largest-scale experiment for Hesian spectrum is GPT2 (125M) on Openwebtext, where the number of tokens $n=9$ Billon. To calculate $\nabla^{2}\mathcal{L}(w)d$ on all these 9B tokens, it requires about 9 GPU days on eight A100-80GB GPUs. Since SLQ requires at least 1,000 times query of $\nabla^{2}\mathcal{L}(w)d$, a complete run of SLQ would take at least 9,000 days on eight A100-80GB GPUs, which is unaffordable. In this work, we use the largest possible batch size (with gradient accumulation tricks) to approximate $\nabla^{2}\mathcal{L}(w)$ under the constraints of GPU bandwidth and time limit. More detailed setup of SLQ are shown as follows.
  • ResNet18 (20M) and VGG16 (138M) on ImageNet. We use the code base of PyTorch Examples  [^4] https://github.com/pytorch/examples/blob/main/imagenet/main.py. We use batch size $=1024$. For the calculation of the blockwise Hessian spectra, we apply SLQ to all parameter blocks except for the BatchNorm layers. In total, it takes about 3 days on one V100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
  • ViT-base (86M) on ImageNet. We use the code base of PyTorch Image Models  [^5] https://github.com/huggingface/pytorch-image-models. We use batch size $=1024$. Due to the large number of parameters, we are not able to calculate the blockwise Hessian spectra for all parameter blocks. Instead, we apply SLQ to: the embedding layer; the output layer; the $1$-st, $6$-th, $12$-th attention blocks; and the $1$-st, $6$-th, $12$-th MLP blocks (note that the $12$-th attention and MLP blocks are the final ones). In total, it takes about 3 days on one V100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
  • BERT(40M) on Cornell Movie-Dialogs Corpus. We use the code base from the blog  [^6] https://medium.com/data-and-beyond/complete-guide-to-building-bert-model-from-sratch-3e6562228891. We use batch size $=327,680$ tokens. For the calculation of the blockwise Hessian spectra, we apply SLQ to all parameter blocks except for the LayerNorm layers. In total, it takes about 12 hours on one V100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
  • GPT2-nano (11M) on Shakespeare. We use the code base of NanoGPT  [^7] https://github.com/karpathy/nanoGPT/. We use batch size $=163,840$ tokens. For the calculation of the blockwise Hessian spectra, we apply SLQ to all parameter blocks with even indices, except for the LayerNorm layers. In total, it takes about 12 hours on one V100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
  • GPT2 (125M) on Openwebtext. We use the code base of NanoGPT. We use batch size $=245,760$ tokens. Due to the large number of parameters, we are not able to calculate the blockwise Hessian spectra for all parameter blocks. Instead, we apply SLQ to: the embedding layer; the output layer; the $1$-st, $4$-th, $8$-th, $12$-th attention blocks; and the $1$-st, $4$-th, $8$-th, $12$-th MLP blocks (note that the $12$-th attention and MLP blocks are the final ones). In total, it takes about 7 days on one A100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
We train all the models under the default configurations in all cases. For CNN training, we follow the rule of thumb to choose the learning rate of AdamW: we set the learning rate of AdamW to be 100×100$\times$100 × smaller than the recommended learning rate of SGD.
C.2 Implementation Details on Figure LABEL:fig_block_diagnal
We use the code base of NanoGPT to train a decoder-only Transformers on 4 consecutive tokens randomly selected from Openwebtext. We set the model configuration as: context window = 2, number of heads = 2 and the embedding dimension = 4. In MLP layers, all widths equal to 4. We change the number of layers from 1 to 4. We remove all the Layer Norms in the model. The rest of the model configurations are set to their default values in the code base. We compute the Hessian on all the parameters blocks in attention and MLP layers. In Figure LABEL:fig_block_diagnal (a), the variables in three blue boxes corresponds to the parameters in Query and Key; Value and Projection; and MLP, respectively. Similarly for the rest of the figures.
Due to the intensive overhead of computing and storing the whole Hessian, we have yet to check the block-diagonal structure on larger models. Rigorously speaking, so far we have not gotten sufficient evidence to claim this structure commonly holds in larger Transformers. It requires new numerical methods to efficiently check the block-diagonal Hessian structure in larger models without explicitly calculating them. We leave it as an interesting future direction.
For completeness, we remark that Collobert [2004], Roux et al. [2007], Martens and Grosse [2015] also observed the block-diagonal structure in (approximated) Hessian of small-scaled MLPs. We restate their findings in Figure LABEL:fig_block_diagnal_existing. [Collobert, 2004, Section 7] further theoretically proved that the block diagonal structure stems from (i) the layer-by-layer structure and (ii) the Cross-Entropy loss. These results suggest that block-diagonal structure might be common in NNs.
C.3 Implementation Details on the MLP experiments in Figure LABEL:fig_mlp_gap
We train a 4-layer MLP on MNIST. We use batch size = 128 and width = 300, 128, and 64 for the hidden layers. We use ReLU activation. We change the degree of heterogeneity by scaling the output of each layer with constant $c\in\mathbb{N}$. We scale $c$ from 1 to 15. We record the average JS distance between the initial Hessian spectra of each pair of layers and find that the average JS distance increases as $c$ increases. As such, we define $c$ as the degree of block heterogeneity. For each $c$, we train SGD and Adam with default hyperparameters by grid-searching the learning rate from 1e-4 to 1e-1 and report the best test accuracy after 1 epoch.
$c$ 1 2 3 4 5 6 7 8 9-15
Blockwise distance 1e-3 1.61 33.99 65.32 260.77 298.39 351.69 353.48 $\operatorname{NaN}$
Table 3: Consider a 4-layer MLP on MNIST. $c$ denotes the scaler that is multiplied to the output of each layer. Blockwise distance denotes the average JS distance between the initial Hessian spectra of each pair of layers. We find that the average JS distance increases as $c$ increases. For $c\geq 9$, the calculation of JS distance encounters numerical instability due to the division of near-zero values and returns $\operatorname{NaN}$.
Appendix D More Experimental Results
D.1 Performance comparison of AdamW and SGD on Different Architectures
Here, we show the performance comparison of AdamW and SGD on different models.
Appendix E Proofs
E.1 Proof of Proof of Proposition 1
Let $H=\begin{bmatrix}L&0\\ 0&\mu\end{bmatrix}$, where $L>\mu>0$. We choose the initial point as $w^{0}=(w_{1}^{0},w_{2}^{0})=(\sqrt{\mu/L},\sqrt{L/\mu})$. By the update rule of GD, we have
$\mathcal{L}(w^{t+1})$ $=$ $\mathcal{L}\left(w^{t}-\eta\nabla\mathcal{L}(w^{t})\right)$ (10)
$=$ $\frac{1}{2}(w^{t}-\eta Hw^{t})^{T}H(w^{t}-\eta Hw^{t})$
$=$ $(w^{t}_{1})^{2}|1-\eta L|L+(w^{t}_{2})^{2}|1-\eta\mu|\mu$
$=$ $|1-\eta L|^{t}L\frac{\mu}{L}+|1-\eta\mu|^{t}\mu\frac{L}{\mu}$
$=$ $\mu|1-\eta L|^{t}+L|1-\eta\mu|^{t}$
To proceed, we discuss the following cases:
When $\eta\leq 1/L$, since $|1-\eta L|^{t}$ and $|1-\eta\mu|^{t}$ are monotonically decreasing , the optimal solution is $\eta=1/L$.
When $\eta\geq 1/\mu$, since $|1-\eta L|^{t}$ and $|1-\eta\mu|^{t}$ are monotonically increasing , the optimal solution is $\eta=1/\mu$.
When $1/L\leq\eta\leq 1/\mu$, (10) can be written as $g_{t}(\eta)=\mu(\eta L-1)^{t}+L(1-\eta\mu)^{t}$. Take the first-order and the second-order derivative of the g, we can obtain $g_{t}^{\prime}(\eta)=tL\mu(\eta_{L}-1)^{t-1}-t\mu L(1-\eta\mu)^{t-1}$ and $g_{t}^{\prime\prime}(\eta)=t(t-1)L^{2}\mu(\eta L-1)^{t-2}+t(t-1)\mu^{2}(1-\eta\mu)$. Since $g^{\prime\prime}_{t}(\eta)\geq 0$ for all $\eta\in[1/L,1\mu]$, the function g is convex. By solving the equation that $g_{t}^{\prime}(\eta)=0$, we can obtain $\eta=\frac{2}{L+\mu}$ is a solution for all $t$. Plugging this result into (10) and rearranging the terms, we conclude the proof of Proposition 1.
E.2 Proof of Theorem 1
Without loss of generality, we assume $h=0$. This is because minimizing $\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ is equivalent to minimizing $\mathcal{L}(w)=\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ where $w^{*}=H^{-1}h$. By a linear transformation $z=w-w^{*}$, BGD for minimizing $\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ starting from $w^{0}$ is equivalent to BGD for minimizing $\frac{1}{2}z^{T}Hz$ starting from $z^{0}=w^{0}-w^{*}$. Thus we can assume $w^{*}=0$, or equivalently, $h=0$. When $h=0$, the update form of BGD becomes
$w^{t+1}=w^{t}-\eta D_{BGD}^{-1}Hw^{t},$
where $D_{BGD}=\operatorname{diag}\left(\left(\underbrace{\eta_{1},\cdots,\eta_{1}}_{d_{1}\text{times}},\cdots,\underbrace{\eta_{L},\cdots,\eta_{L}}_{d_{L}\text{times}}\right)^{T}\right)\in\mathbb{R}^{d\times d}.$
In the following, we will first prove the rate of BGD for each block and then combine them together to get the total rate. For the $l$-th block, we choose $\eta_{l}=\frac{1}{\lambda_{l,1}}$. The update rule of BGD for the $l$-th block becomes:
$w_{l}^{t+1}=w_{l}^{t}-\frac{1}{\lambda_{l,1}}H_{l}w_{l}^{t}$
Now we apply the standard convergence proof of GD to the $l$-th block. By descent lemma, we have
$[\mathcal{L}(w^{t+1})]_{l}$$\leq$ $[\mathcal{L}(w^{t})]_{l}+\langle w_{l}^{t+1}-w_{l}^{t},[\nabla\mathcal{L}(w^{t})]_{l}\rangle+\frac{\lambda_{l,1}}{2}\|w_{l}^{t+1}-w_{l}^{t}\|_{2}^{2}$ (11)
$\leq$ $[\mathcal{L}(w^{t})]_{l}+\langle w_{l}^{t+1}-w_{l}^{t},H_{l}w_{l}^{t}\rangle+\frac{\lambda_{l,1}}{2}\eta^{2}w_{l}^{t}H_{l}^{T}H_{l}w_{l}^{t}$ (12)
$=$ $[\mathcal{L}(w^{t})]_{l}-\frac{1}{2\lambda_{l,1}}w_{l}^{t}H_{l}^{T}H_{l}w_{l}^{t}$ (13)
$\leq$ $[\mathcal{L}(w^{t})]_{l}-\frac{\lambda_{l,d_{l}}}{2\lambda_{l,1}}w_{l}^{t}H_{l}w_{l}^{t}$ (14)
$=$ $\left(1-\frac{1}{\kappa_{l}}\right)[\mathcal{L}(w^{t})]_{l}$ (15)
Substract both sides by $\mathcal{L}^{*}$ and we get $[\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\leq\left(1-\frac{1}{\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$. Summing up both sides over $l\in[L]$,we have
$\mathcal{L}(w^{t+1})-\mathcal{L}^{*}$ $=$ $\sum_{l=1}^{L}[\left(\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$=$ $\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$
This concludes the proof of Theorem 1.
E.3 Proof of Theorem 2
Similarly as the proof of Theorem 1, we assume $h=0$. The update rule of Adam becomes
$w^{t+1}=w^{t}-\eta(D_{Adam}^{0})^{-1}Hw^{t},$
where $D_{Adam}^{0}=\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))^{\frac{1}{2}}=\operatorname{diag}(|Hw^{0}|)$. We denote $d_{t}=\eta(D_{Adam}^{0})^{-1}Hw^{t}$ and thus we have $w^{t}=\frac{1}{\eta}H^{-1}D_{Adam}^{0}d^{t}$ and $w^{t+1}=w^{t}-d_{t}$. These relations also hold for each block by changing the notation to $H_{l}$ $w_{l}^{t}$, $D_{Adam}^{0}$, and $d_{l}^{t},$ etc.. Following the framework in [Sun and Ye, 2021], we try to bound the error yet to be optimized (a.k.a., cost-to-go) and the per-step improvement, respectively. The ratio of these two terms characterizes the rate of convergence. We now express both terms using $d_{l}^{t}$. For the cost-to-go term for the $l$-th block, we have
$[\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}=\frac{1}{2}(w_{l}^{t})^{T}H_{l}w_{l}^{t}=\frac{1}{2\eta^{2}}(d_{l}^{t})^{T}D_{Adam,l}^{0}H_{l}^{-1}D_{Adam,l}^{0}d_{l}^{t}.$ (16)
For the per-step improvement, we have
$[\mathcal{L}(w^{t})]_{l}-[\mathcal{L}(w^{t+1})]_{l}$ $=$ $\frac{1}{2}(w_{l}^{t})^{T}H_{l}w_{l}^{t}-\frac{1}{2}(w_{l}^{t+1})^{T}H_{l}w_{l}^{t+1}$ (17)
$=$ $\frac{1}{2}(w_{l}^{t})^{T}H_{l}w_{l}^{t+1}-\frac{1}{2}(w_{l}^{t}-d^{t})^{T}H_{l}(w_{l}^{t}-d_{l}^{t})$
$=$ $(d_{l}^{t})^{T}H_{l}w_{l}^{t}-\frac{1}{2}(d_{l}^{t})^{T}H_{l}d_{l}^{t}$
$=$ $\frac{1}{2}(d_{l}^{t})^{T}\left(\frac{2}{\eta}D_{Adam,l}^{0}-H_{l}\right)d_{l}^{t}.$
To proceed, we denote $\hat{H}=(D_{Adam}^{0})^{-1}H$ and we denote its eigenvalues as $\hat{\lambda}_{1}\geq\hat{\lambda}_{2}\geq\cdots\hat{\lambda}_{d}$. Similarly, we denote $\hat{H}_{l}=(D_{Adam,l}^{0})^{-1}H_{l}$ and its eigenvalues $\hat{\lambda}_{l,1}\geq\hat{\lambda}_{l,2}\geq\cdots\hat{\lambda}_{l,d_{l}}$. Let $\eta=\min_{l\in[L]}C_{l,1}$, we have
$\frac{[\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}}{[\mathcal{L}(w^{t})]_{l}-[\mathcal{L}(w^{t+1})]_{l}}$ $=$ $\frac{\frac{1}{\eta^{2}}(d_{l}^{t})^{T}D_{Adam,l}^{0}H_{l}^{-1}D_{Adam,l}^{0}d_{l}^{t}}{(d_{l}^{t})^{T}\left(\frac{2}{\eta}D_{Adam,l}^{0}-H_{l}\right)d_{l}^{t}}$ (18)
$\leq$ $\left\|\frac{1}{\eta^{2}}\left(\frac{2}{\eta}D_{Adam,l}^{0}-H_{l}\right)^{-1}D_{Adam,l}^{0}H_{l}^{-1}D_{Adam,l}^{0}\right\|_{2}$
$\overset{(*)}{\leq}$ $\frac{C_{l,2}^{2}\lambda_{l,1}^{2}}{(\min_{l\in[L]}C_{l,1}^{2})\lambda_{l,1}\lambda_{l,d_{l}}}$ (19)
$\leq$ $\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l},$ (20)
[…]
$\mathcal{L}(w^{t+1})-\mathcal{L}^{*}$ $=$ $\sum_{l=1}^{L}\left([\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
$=$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$

### v2

@@
 SGD performs worse than Adam by a significant margin on Transformers, but the reason remains unclear.
-In this work, we provide an explanation of SGD’s failure on Transformers through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs badly on problems with block heterogeneity.
+In this work, we provide an explanation of SGD’s bad performance on Transformers through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs badly on problems with block heterogeneity.
 To validate that heterogeneity hampers SGD, we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD works well on problems without block heterogeneity but performs badly when the heterogeneity exists.
-Our initial theoretical analysis indicates that SGD fails because it applies one single learning rate for all blocks, which cannot handle the heterogeneity among blocks.
-The failure could be rescued if we could assign different learning rates across blocks, as designed in Adam.
+Our initial theoretical analysis indicates that SGD performs poorly because it applies one single learning rate to all blocks, which cannot handle the heterogeneity among blocks.
+This limitation could be ameliorated if we use coordinate-wise learning rates, as designed in Adam.
+[^1] Our code is available at https://github.com/zyushun/hessian-spectrum.
+{NoHyper}*: Correspondence author.
 1 Introduction
-Transformers (Vaswani et al., 2017) have become a major workhorse behind much remarkable progress in AI development (e.g., (Achiam et al., 2023)).
+Transformers [84] have become a major workhorse behind much remarkable progress in AI development (e.g., [1]).
 However, the understanding of Transformer training remains limited.
-For instance, Transformer training largely relies on Adam(W) (Kingma and Ba, 2014; Loshchilov and Hutter, 2017).
-In contrast, Stochastic Gradient Descent with momentum (SGD)  [^1] We introduce the update rules of Adam(W) and SGD in Appendix B.1., which is the de-facto optimizer for convolution neural networks (LeCun et al., 1998) (CNNs), performs poorly on Transformers (see Figure LABEL:fig:cv_figure and LABEL:fig:nlp_figure in Appendix D as evidence).
-Yet, the cause of failure remains unclear.
-It is an intriguing question why SGD fails on Transformers.
-First, from a theoretical perspective, this topic helps reveal more properties of the Transformer training loss landscape, which helps boost understanding of training objectives and current optimizers.
-Second, from a computational perspective, Adam requires heavier memory storage than SGD.
-Compared with SGD, Adam needs to store an additional second-order momentum, which takes at least the same memory of the model size.
-The memory consumption of Adam has become a major overhead in large-scale transformer training (Rasley et al., 2020).
-Identifying the cause of failure may help researchers design more efficient optimizers.
-In this work, we explore the failure mode of SGD through the lens of Hessian.
-We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum as an example).
-By theory, the full Hessian spectrum largely determines the behavior of gradient methods (Nesterov, 2013; Goh, 2017; Sun, 2019; Goujaud et al., 2022), so we suspect it can also help uncover the reasons behind SGD failure on Transformers.
-Using numerical linear algebra tools (Bai et al., 1996), we empirically compare the full spectra of CNNs (the scenarios where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam).
-Unfortunately, in Figure LABEL:fig_full_spectrum, we find that the spectra on CNNs and Transformers are often largely similar despite the different optimizer behaviors on these tasks.
-As such, we have not identified critical phenomena in the full Hessian spectra associated with SGD’s failure on Transformers.
-To reveal the cause of failure, we need a more fine-grained investigation into Hessian.
-What else could make SGD fail on Transformers, but not on CNNs?
-By dissecting the structures of CNNs and Transformers, we observe that CNN architectures are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the interleaved stacking of disparate parameter blocks (e.g.
-Query, Key, Value blocks in attention and MLP layers).
-We suspect that these design differences may lead to different properties in optimization.
-Intuitively, disparate parameter blocks would contribute differently to the total loss.
-So each block might prefer a distinct treatment by optimizers, which can be provided by Adam but not by SGD.
-This motivates us to investigate the Hessian spectrum of each parameter block, abbreviated as the blockwise Hessian spectrum.
-By inspecting the blockwise Hessian spectrum, we find a possible explanation why SGD fails on Transformers.
-We find that Transformers are “heterogeneous": for all Transformers we checked, Hessian spectra of different parameter blocks are dramatically different.
-We then verify that such heterogeneity indeed hampers SGD via extensive experiments on various Transformers, CNNs, MLPs , and quadratic problems.
-Our contributions are summarized as follows.
-  • Exploring the failure mode of SGD on transformers.
-We provide an explanation why SGD largely lags behind Adam on Transformers through the lens of blockwise Hessian spectrum.
-This explanation consists of two parts.
-First, we identify a phenomenon called “block heterogeneity", which states that Hessian spectra are dramatically different across parameter blocks.
-We find this phenomenon prevalent on all Transformers in our experiments, but not on CNNs.
-Second, we verify that block heterogeneity hampers SGD.
-More specifically, on various kinds of Transformers, CNNs, MLPs, we find that SGD always performs badly on problems with block heterogeneity but works well otherwise.
+For instance, Transformer training largely relies on the Adam optimizer [45, 56].
+In contrast, stochastic gradient descent with momentum (SGD) [^2] We introduce the update rules of Adam(W) and SGD in Appendix C.1., the de-facto optimizer for convolution neural networks (CNNs) [48], performs poorly on Transformers (e.g., Figure LABEL:fig_blockwise_spectrum).
+Yet, the reasons behind this poor performance remain unclear.
+Understanding why SGD performs badly on Transformers is an intriguing question.
+First, from a theoretical perspective, this can help us better understand the training of Transformers and more generally, neural networks.
+Second, from a computational perspective, the understanding may inspire the design of better algorithms for training neural networks.
+In this work, we explore the cause for SGD’s deficiency through the lens of Hessian.
+We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum).
+By theory, the full Hessian spectrum largely determines the behavior of gradient-based methods [63, 33, 80, 37], so we suspect it may also help explain SGD’s bad performance on Transformers.
+Using tools from numerical linear algebra [7], we empirically compare the full spectra of CNNs (where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam).
+Unfortunately, as shown in Figure LABEL:fig_full_spectrum, the spectra for CNNs and Transformers are often largely similar despite the different optimizer behaviors.
+As such, we have not identified critical features in the full Hessian spectra associated with SGD’s bad performance on Transformers.
+To reveal the cause, a more fine-grained investigation into the Hessian is needed.
+What would cause SGD to perform badly on Transformers but not on CNNs?
+By dissecting the structures of CNNs and Transformers, we notice that CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the non-sequential stacking of disparate parameter blocks (e.g.
+Query, Key, Value, Output projection blocks in attention and MLP layers).
+We hypothesize that these architectural differences might lead to different optimization properties.
+Intuitively, disparate parameter blocks contribute differently to the overall loss.
+So each block might benefit from a specialized treatment by optimizers, a flexibility offered by Adam but not by SGD.
+This observation motivates us to investigate the Hessian spectrum of each parameter block, which we refer to as the blockwise Hessian spectrum.
+By inspecting the blockwise Hessian spectrum, we discover a possible explanation for SGD’s bad performance: the “heterogeneity” inherent in Transformers.
+We provide both empirical and theoretical evidence to support this explanation.
+Our contributions can be summarized as follows:
+  • Failure mode of SGD on Transformers.
+We explain why SGD underperforms Adam on Transformers by examining the blockwise Hessian spectrum.
+First, we identify a phenomenon called ’block heterogeneity’, which refers to the large differences in the Hessian spectra across parameter blocks.
+This block heterogeneity is observed in all examined Transformers but not in CNNs.
+Second, we verify that block heterogeneity hinders SGD.
+Across various Transformers, CNNs, and MLPs, we show that SGD consistently performs poorly on problems with block heterogeneity but works well otherwise.
   • Theoretical results on quadratic models.
-We build various convex quadratic problems, some with block heterogeneity and others without block heterogeneity.
-We find that GD is significantly worse than Adam on problems with block heterogeneity, but performs similarly to Adam otherwise.
-We provide a theoretical result which shows that GD indeed works badly on quadratic problems with block heterogeneity.
-Our analysis suggests that one important reason for (S)GD failure is that it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks.
-The failure could be saved if we assign different learning rates across blocks like Adam.
+We construct convex quadratic problems with and without block heterogeneity and find that gradient descent (GD) performs significantly worse than Adam on problems with block heterogeneity but comparably otherwise.
+Our theoretical analysis shows that GD struggles on quadratic problems with block heterogeneity, suggesting that using a single learning rate for all blocks is a key reason for GD’s bad performance.
+The deficiency can be mitigated by assigning different learning rates across blocks, as Adam does.
+We emphasize that we do not claim block heterogeneity is the only cause for SGD’s bad performance, but just that it is at least one important cause.
+We verify, both empirically and theoretically, that SGD performs poorly when block heterogeneity is present.
 2 Main Results
@@
 Notations.
-We denote the full-batch training loss as $\mathcal{L}(w)\equiv\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}(w)$, where $n$ is the number of minibatches, $\mathcal{L}_{i}(w)$ is the loss of $i$-th minibatch and $w\in\mathbb{R}^{d}$ is the neural network parameters.
+We denote the training loss as $\mathcal{L}(w)$, where $w\in\mathbb{R}^{d}$ is the neural network parameters.
 We denote the gradient and Hessian of the training loss w.r.t.
 neural network parameters as $\nabla\mathcal{L}(w)\in\mathbb{R}^{d}$ and $\nabla^{2}\mathcal{L}(w)\in\mathbb{R}^{d\times d}$, respectively.
-We use $\nabla\mathcal{L}_{i}(w)\in\mathbb{R}^{d}$ and $\nabla^{2}\mathcal{L}_{i}(w)\in\mathbb{R}^{d\times d}$ to denote the $i$-th minibatch counterparts.
 We use $[d]$ to denote the index set $\{1,2,\cdots,d\}$.
@@
 We denote $[\nabla^{2}\mathcal{L}(w)]_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ as the Hessian of $l$-th parameter-block $w_{l}$, where $[\nabla^{2}\mathcal{L}(w)]_{l,i,j}=\frac{\partial^{2}}{\partial_{w_{l,i}}\partial_{w_{l,j}}}\mathcal{L}(w_{l})$.
-Note that $[\nabla^{2}\mathcal{L}(w)]_{l}$ is the $l$-th principle block sub-matrix of $\nabla^{2}\mathcal{L}(w)$.
-Preliminaries.
-Hessian of large-scale NNs are intractable to obtain due to the numerical difficulties of computing and storing them.
-Fortunately, there are extensive numerical linear algebra methods to approximate the Hessian spectrum without calculating the Hessian explicitly.
-Among which, the Stochastic Lanczos Quadrature method (SLQ) (Bai et al., 1996) is known to be one of the oldest yet the most powerful methods.
-SLQ is a combination of the stochastic trace estimator (Hutchinson, 1989), the Gaussian quadrature (Golub and Meurant, 2009; Epperson, 2013), and the Lanczos algorithm (Lanczos, 1950).
-In short, SLQ takes a handful of Hessian-vector products as input, where the vectors are sampled from the normalized Rademacher distribution.
-It returns a smooth curve on $\mathbb{R}$ that approximates the histograms of eigenvalues (see Figure LABEL:fig_full_spectrum as an example).
-Since the derivation of SLQ is quite involved, we defer the detailed description of SLQ to Appendix B.2.
-The algorithmic form of SLQ can be seen in Algorithm 5 in the Appendix.
-We provide running time analysis and detailed setup of SLQ in Appendix C.1.
-In this work, we apply SLQ to various CNNs and Transformers on different datasets.
-As for CNNs, we study CNNs including ResNet18 (11M) and VGG16 (138M) on ImageNet (He et al., 2016; Simonyan and Zisserman, 2014).
-As for Transformers, we study ViT-base (86M) on ImageNet (Dosovitskiy et al., 2020), BERT (40M) on Cornell Movie-Dialogs Corpus (Devlin et al., 2018), and GPT2 (125M) on Openwebtext (Radford et al., 2019).
-For a fair comparison with ResNet18, we also consider smaller models like GPT2-nano [^2] https://github.com/karpathy/nanoGPT/ (11M) on English corpus.
-The largest model we conduct is GPT2 with $125$ million parameters.
-Detailed experimental setups are shown in Appendix C.1.
-For all these Transformer-based tasks, SGD performs significantly worse than Adam; while for CNN-based takes, Adam and SGD are similar.
-The performance comparison is shown in Figure LABEL:fig:cv_figure and LABEL:fig:nlp_figure in the Appendix.
-We estimate the spectrum of (1) the full Hessian $\nabla^{2}\mathcal{L}(w)$ and (2) the Hessian of all parameter blocks $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$.
-For (2), we split $w$ according to the default partition in PyTorch implementation, e.g., Embedding layer, Query in 1st attention layer, Key in 1st attention layer, Value in 1st attention layer, etc..
-Note that the term “block" is a bit different from the term “layer".
-For instance, Query and Key could be in the same layer, but they are different parameter blocks.
-In the following contexts, we refer to “full Hessian spectrum" as the estimated spectrum of $\nabla^{2}\mathcal{L}(w)$ and “blockwise Hessian spectrum" as the estimated spectrum of $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$.
+Note that $[\nabla^{2}\mathcal{L}(w)]_{l}$ is the $l$-th principal block sub-matrix of $\nabla^{2}\mathcal{L}(w)$.
+Setups.
+Hessian of large-scale NNs are intractable to compute and store.
+In this work, we apply a numerical tool called Stochastic Lanczos Quadrature method (SLQ) [7] to approximate the Hessian spectrum.
+SQL uses a smooth curve on $\mathbb{R}$ to approximate the histograms of eigenvalues (see Figure LABEL:fig_full_spectrum as an example).
+A detailed introduction to SLQ is provided in Appendix C.2.
+All experimental setups in this section is shown in Appendix D.
+We focus primarily on the following models/tasks.
+  • CNNs. We study ResNet18 (11M) and VGG16 (138M) on ImageNet [40, 79]. On these tasks, SGD performs on par with Adam. See Figure LABEL:fig:cv_figure in Appendix B for the evidence.
+  • Transformers.
+We study Transformer with various scales and modalities, including GPT2 (125M) on OpenWebText [71]; ViT-base (86M) on ImageNet [27]; BERT (40M) on Cornell Movie-Dialogs Corpus [25]; GPT2-nano [^3] https://github.com/karpathy/nanoGPT/ (11M) on English corpus.
+On these tasks, SGD performs significantly worse than Adam.
+See Figure LABEL:fig:nlp_figure in Appendix B for the evidence.
+For each model, we estimated (1) the full Hessian spectrum $\nabla^{2}\mathcal{L}(w)$, and (2) the blockwise Hessian spectrum $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$.
+For the latter, we split $w$ according to the default partition in PyTorch implementation, e.g., Embedding layer, Query in each attention layer, Key in each attention layer, Value in each attention layer, etc.
+Note that the term “block" differs from the term “layer".
+For instance, Query and Key can reside in the same layer but are different parameter blocks.
 2.2 Full Hessian Spectrum Is Not Informative Enough
-Motivated by classical theory, we start by investigating the full Hessian spectrum of Transformers.
-We choose to study the full Hessian spectrum for two reasons.
-First, as mentioned in Section 1, the full Hessian spectrum is known to largely determine the behavior of gradient methods.
-For instance, the range of the spectrum will affect the convergence rate (Nesterov, 2013).
-Second, the full Hessian spectrum has been shown to be informative in explaining multiple phenomena in neural nets (NNs), e.g., how BatchNorm accelerates the training (Ghorbani et al., 2019).
-Due to these reasons, we hypothesize that the full Hessian spectrum can also help uncover the reasons behind SGD failure on Transformers.
-We compare the full spectra of CNNs (the scenarios where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam).
-The results are shown in Figure LABEL:fig_full_spectrum.
-Unfortunately, our experiments show that the full Hessian spectrum might not be enough to explain the SGD failure on Transformers.
-We explain as follows.
-The primary information of the spectrum lies in its (1) dispersion, (2) shape, and (3) evolution along training.
-As for (1), we find that eigenvalues are dispersed very similarly and we do not observe any notably large eigenvalues in the spectra of Transformers.
-So (1) does not seem to be related to the failure of SGD.
-We further inspect (2) and (3).
-For all CNNs and Transformers in Figure LABEL:fig_full_spectrum, we find similar phenomena: the shape of the spectrum is approximately symmetrical around 0 at initialization.
-As training proceeds, the majority of large negative eigenvalues disappear and the shape evolves into a combination of a “bulk" and some “outliers".
-Since the spectral shape and evolution are quite similar on Transformers and CNNs, they cannot explain the failure of SGD on Transformers, either.
-In summary, we have not identified critical phenomena in the full Hessian spectra linked to SGD’s failure on Transformers.
+We study the full Hessian spectrum of Transformers for two reasons.
+First, as stated in Section 1, the Hessian spectrum significantly influences the behavior of gradient methods [63].
+Second, previous research shows that the Hessian spectrum provides insights into neural network phenomena, like BatchNorm’s effect on training speed [32].
+Therefore, we hypothesize that the Hessian spectrum may also explain SGD’s bad performance on Transformers.
+We compare the full Hessian spectra of CNNs (where SGD works well) and those of Transformers (where SGD struggles), as shown in Figure LABEL:fig_full_spectrum.
+Unfortunately, the results suggest that the full Hessian spectrum alone may not suffice to explain why SGD struggles on Transformers.
+We elaborate as follows.
+The primary information in the spectrum lies in its (A) dispersion, (B) shape, and (C) evolution during training.
+Regarding (A), we observe that the eigenvalues are dispersed similarly across different models, with no notably large outlier for Transformers.
+Thus, dispersion does not seem to be related to the bad performance of SGD.
+We further investigate (B) and (C).
+For all CNNs and Transformers in Figure LABEL:fig_full_spectrum, we observe similar phenomena: the spectrum’s shape is approximately symmetrical around 0 at initialization.
+As training proceeds, the majority of negative eigenvalues disappear, and the shape evolves into a combination of a “bulk” and some “outliers”.
+Since the spectral shape and evolution are quite similar for both Transformers and CNNs, they cannot explain the incompetence of SGD on Transformers either.
+In summary, we have not identified any critical phenomena in the full Hessian spectra that can be linked to SGD’s bad performance on Transformers.
 2.3 Main Findings Through Blockwise Hessian Spectra
-We have shown that the full Hessian spectrum seems not sufficient for uncovering the reason why SGD fails on Transformers.
-What else could make SGD fail on Transformers, but not on CNNs?
-We find two important features that are overlooked in the full Hessian spectrum analysis above.
-  • (1) The Hessian structure is overlooked.
-There are numerical results showing that the Hessians are close to block-diagonal matrices on MLPs (Collobert, 2004; Roux et al., 2007; Martens and Grosse, 2015).
-We restate their findings in Appendix C.2.
-Further, (Collobert, 2004, Section 7) theoretically shows that the block diagonal structure comes from (i) the layer-by-layer design in NNs and (ii) the Cross-Entropy loss.
-Following this line of work, we also observe a near block-diagonal Hessian in small Transformers in Figure LABEL:fig_block_diagnal, where the variables in each principle block correspond to the parameters of each block in the Transformer.
+What other factors could cause SGD to struggle on Transformers but not on CNNs? We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
+  • 1.
+    Hessian structure.
+Existing literatures showed that the Hessians of MLPs are close to block-diagonal matrices [18, 73, 60].
+We restate their findings in Appendix B.
+Further, [18, Section 7] theoretically attributes the block diagonal structure to (i) the layer-by-layer design in NNs and (ii) the Cross-Entropy loss.
+Following this line of work, we also observe a near block-diagonal Hessian in small Transformers in Figure LABEL:fig_block_diagnal, where the variables in each principal block correspond to the parameters of each block in the Transformer.
 These results suggest that near block-diagonal Hessian might be common in NNs.
-  • (2) The build-up rules of Transformers is overlooked.
-Transformers are built upon stacking of disparate parameter blocks, e.g.
-Query, Key, Value blocks in attention, and MLP layers.
-This is largely different from CNNs, which are constructed by the repetitive stacking of similar parameter blocks (convolution layers).
-Since each parameter blocks in Transformers are designed differently, they might have different properties for optimization, which might further affect the optimizer behavior.
-Combining (1) and (2), we hypothesize that the blockwise Hessian spectrum might provide some insights on distinguishing CNNs and Transformers.
+  • 2.
+    Build-up rules of Transformers.
+As shown in Figure LABEL:fig_blockwise_spectrum, CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers).
+In contrast, Transformers consist of disparate parameter blocks, e.g.
+Query, Key, Value in attention, and MLP layers.
+Further, these blocks are stacked in a non-sequential manner.
+Due to the different designs of these blocks, they may have different properties for optimization, which can further affect the optimizer behavior.
+Combining these together, we hypothesize that the blockwise Hessian spectrum, i.e., the spectrum of principal blocks of Hessian $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$, might provide additional insights.
 What extra information can be contained in the blockwise spectrum but not in the full spectrum?
 By definition, the blockwise Hessians form the principal block sub-matrix of the full Hessian.
-For block diagonal matrices, blockwise Hessian encodes the location of eigenvalues, i.e., which block does an eigenvalue (of the full Hessian) lie in, while the full spectrum ignores this.
-In the following, we study blockwise Hessian spectra on various models.
-We will show that blockwise spectra indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
-We here demonstrate the shape of blockwise spectra in CNNs and Transformers.
-We sample four blocks per model and present the spectra in Figure LABEL:fig_blockwise_spectrum.
-In Transformers, the Hessian spectra of embedding, attention, and MLP blocks are largely different.
-In contrast, in CNNs, the spectra of convolution layers are similar.
+We note that Transformers are observed to have a near block-diagonal Hessian.
+For block diagonal matrices, blockwise spectra encode the location of eigenvalues, i.e., which block an eigenvalue (of the full matrix) resides in.
+In contrast, the full spectrum overlooks this location information.
+In the following, we study the blockwise Hessian spectra of various models and show that they indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
+We here demonstrate the shape of blockwise spectra in VGG16 [40] (CNN) and BERT [25] (Transformer).
+We sample four blocks for each model and present the spectra in Figure LABEL:fig_blockwise_spectrum.
+In BERT, the Hessian spectra of embedding, attention, and MLP blocks are largely different.
+In contrast, in ResNet, the spectra of convolution layers are similar.
 We further verify this observation for the rest of the parameter blocks.
-We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs.
-We show the results in Figure LABEL:fig_blockwise_heatmap.
-The results in Figure LABEL:fig_blockwise_heatmap showed an interesting phenomenon: for all the Transformers we checked, the blockwise Hessian spectra are largely different across blocks.
+We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs and show the results in Figure LABEL:fig_blockwise_heatmap.
+The results in Figure LABEL:fig_blockwise_heatmap showed a new phenomenon: for all Transformers we checked, the blockwise Hessian spectra are largely different across blocks.
 In the following, we refer to this phenomenon as “block heterogeneity".
-As a control group experiment, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed.
+In contrast, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed.
 We refer to this phenomenon as “block homogeneity".
 These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers.
-Intuitively, the block-homogeneous Hessian in a CNN comes from the repetitively similar convolution layers, while the block-heterogeneous Hessian in a Transformer stems from the interleaved stacking disparate layers such as Query, Value, and MLPs.
-In the following, we will show that the block heterogeneity is strongly correlated with the failure of SGD on Transformers.
-2.4 SGD Fails on Simple MLPs with Block Heterogeneity
-We now further link block heterogeneity to the failure of SGD.
-We construct a simple example without attention layers but has block heterogeneity and find that SGD still fails.
-We consider a 4-layer MLP on MNIST.
-Using standard MLP, we find that the Hessian exhibits a homogeneous block structure and SGD converges as fast as Adam.
-We then scale the output of each layer by a constant $c$.
-We define $c$ as the degree of heterogeneity since larger $c$ would bring more heterogeneity to the Hessian (see Table 3 in the Appendix).
-As shown in Figure LABEL:fig_mlp_gap, we find that the gap between Adam and SGD increases with the degree of heterogeneity.
-This example shows that attention might not be the root cause of SGD failure, but the block heterogeneity (brought by attention) might be.
+Intuitively, the block-homogeneity in CNNs comes from the repetitively similar convolution layers, while the block-heterogeneity in Transformers stems from the non-sequential stacking disparate layers such as Query, Value, and MLPs.
+In the following, we will show that the block heterogeneity is strongly correlated with SGD’s bad performance on Transformers.
+2.4 SGD Performs Badly on Various Tasks with Block Heterogeneity
+Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD performs badly on Transformers.
+(2) Transformers have block-heterogeneity.
+Now we further link block heterogeneity to SGD’s bad performance on non-Transformer models.
+This would directly establish a connection between block heterogeneity and SGD’s bad performance, without going through Transformers or attention blocks as an intermediary.
+We consider one man-made example and one real-world example.
+Example 1: A man-made MLP.
+We consider a 4-layer MLP on MNIST and change the degree of heterogeneity by scaling each layer by constant $c$.
+Figure LABEL:fig_mlp_gap (a) shows SGD fails as heterogeneity grows.
+Example 2: MLP-mixer.
+We consider MLP-mixer [82], a famous all-MLP architecture that outperforms CNNs and ViTs on some vision tasks.
+Figure LABEL:fig_mlp_gap (b) (c) shows that the initial Hessian of MLP-mixer has block heterogeneity and SGD lags behind Adam on this architecture.
+2.5 Reduced Block Heterogeneity in Pre-trained Transformers
+We remark that different Transformers exhibit different levels of block heterogeneity.
+Although all examined Transformers show strong block heterogeneity, we find that this heterogeneity can be mitigated, resulting in less performance deterioration for SGD.
+As illustrated in Figure LABEL:fig_sft_gap, pre-trained GPT2 on SFT tasks can exhibit less block heterogeneity compared to pre-training GPT2 from scratch (Figure LABEL:fig_blockwise_heatmap (f)).
+In this case, although SGD is still slower than Adam, it achieves a similar loss at convergence.
+Compared with training GPT2 from scratch (Figure LABEL:fig:nlp_figure (d) in Appendix B), the performance gap between SGD and Adam is significantly narrowed down.
+These findings suggest that the heterogeneity induced by architectural design can be alleviated by selecting “good” weights.
+This partly explains why simpler methods like SGD and even its zeroth-order version can still be effective for fine-tuning language models, albeit with slower convergence [58, 59].
 2.6 Implication on Choosing SGD or Adam
-The above findings could bring up an interesting empirical guidance: we can compute the blockwise spectrum of initial Hessian, and then decide whether to use Adam or SGD.
-Such a method could be useful in scenarios in training large models that are not mainstream Transformers or CNNs, e.g., Mamba (Gu and Dao, 2023).
+The above findings show that SGD can perform badly on various architectures. This leads to an intriguing question: Can we predict the bad performance of SGD before training begins?
+Our findings can bring up an empirical guidance: we can compute the blockwise spectrum of initial Hessian, and then decide whether to use Adam or SGD.
+Such a method could be useful in scenarios in training large models that are not mainstream Transformers or CNNs, e.g., Mamba [38].
 In these cases, there is not much prior experience in choosing optimizers.
 It would be intriguing to decide whether SGD is suitable for the task before the training is launched.
-We present two reasons.
-  • In large-scale model training, memory is limited but the model size grows, so Adam could become a major computational burden due to the memory of 2nd-order momentum (as mentioned in Section 1 and (Rasley et al., 2020)).
-  • SGD is a lightweight alternative to Adam, but the performance of SGD may be bad.
-In large NN training, it might take days (or even weeks) to obtain informative feedback from training logs, so it is expensive to conduct trial and error to choose an appropriate optimizer.
-Specifically, if SGD performs badly, it would be difficult to judge whether the bad performance is due to improper learning rates or other more fundamental reasons.
-Then, it would be hard to decide whether to switch the optimizer.
-Motivated by the above reasons, we here propose a quantitative metric that could predict the failure of SGD before the training is launched.
+One might argue that simple trial is enough: try both SGD and Adam; if Adam is remarkably better, then pick Adam; if Adam and SGD are similar, then pick SGD.
+Nevertheless, this simple approach may not be easy for large models.
+First, for large models, it may take days to know one run of an algorthm is good or not.
+Second, it requires tuning hyperparameters at least a few times to get a reasonably good judgement, making the cost of trial even higher.
+We here propose a quantitative metric that could predict the incompetence of SGD before the training.
 With the help of this metric, we could save much expense on the trial and error for SGD.
-The metric is simply the averaged JS distance among blockwise Hessian spectra at initialization, i.e., the averaged value in each heatmap of Figure LABEL:fig_blockwise_heatmap.
+The metric is simply the averaged JS distance among blockwise Hessian spectra at initialization, i.e., the averaged value in the heatmap of Figure LABEL:fig_blockwise_heatmap.
 We denote it as $JS^{0}$.
 We present $JS^{0}$ of various models in Table 1.
-We note that $JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs.
-As argued in the previous sections, this quantity can predict the failure of SGD.
-Note that $JS^{0}$ is independent of optimizers and could be efficiently checked in priori before training.
-Model ResNet18 VGG16 ViT-base BERT GPT2-nano GPT2
-$JS^{0}$ 0.10 0.09 286.41 53.38 14.93 83.23
-Table 1: $JS^{0}$ denotes the average JS distance between the initial Hessian spectra of each pair of parameter blocks.
-We find that the $JS^{0}$ on CNNs is at least 140×140$\times$140 × smaller than on Transformers.
-$JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs.
-To summarize this section, we provide an explanation for SGD’s failure.
-We identify a phenomenon called block heterogeneity.
-This phenomenon naturally exists in Transformers, but not in CNNs.
-This phenomenon can exist in non-Transformer architectures as well.
-Our experiments suggest a strong connection between the failure of SGD and block heterogeneity.
-Our findings could provide empirical guidance to decide whether to use Adam or SGD.
+Note that $JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs.
+Further, $JS^{0}$ is independent of optimizers and could be checked before training.
+Table 1: $JS^{0}$ denotes the average JS distance between the initial Hessian spectra of each pair of parameter blocks. A larger $JS^{0}$ suggests that the task is more difficult for SGD.
+Model ResNet18 VGG16 GPT2 (pretrained) MLP-mixer BERT GPT2 ViT-base
+$JS^{0}$ 0.10 0.09 18.84 34.90 53.38 83.23 286.41
+To validate the effectiveness of the quantitative metric $JS^{0}$, we summarize $JS^{0}$ of different models and the corresponding SGD performance in Figure 4.
+We find that the performance gap between SGD and Adam becomes greater as $JS^{0}$ increases.
+Thus, $JS^{0}$ can serve as a potential indicator to predict whether SGD may underperform Adam.
+Figure 4: Comparison of $JS^{0}$ and the performance of SGD on different models. We find the performance gap between SGD and Adam becomes greater as $JS^{0}$ increases.
 3 Case Study of Quadratic Models and Preliminary Theory
-Now we construct quadratic functions with block diagonal Hessian.
-We will compare the performance of GD and Adam on Hessian with and without block heterogeneity.
-Initial theoretical results on these quadratic models will be provided.
+Now we study quadratic functions with block diagonal Hessian, with or without block heterogeneity.
 Note that insights on quadratic models could be important for understanding realistic NNs, as mentioned by researchers such as LeCun et al.
-(2002) and OpenAI team (Kaplan et al., 2020).
-3.1 Settings and Additional Notations
-We consider $\min_{w}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$.
+[49] and OpenAI team [44].
+Setups and additional notations. We consider the following quadratic minimization.
+$\min_{w\in\mathbb{R}^{d}}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w,$
+where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$.
 We denote $\mathcal{L}^{*}$ as the minimum value of $\mathcal{L}(w)$.
@@
 3.1 Experimental Observations
-For all experiments, we choose $h=0$. We consider four types of Hessian $H$ as follows.
+We consider four types of Hessian $H$ as follows. For all cases, we set condition number = 5000.
   • Case 1: Hessian with Transformer-type spectra.
-We choose $L=4$ and $d_{l}=100$.
-For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are independent standard Gassian random matrix and $\Lambda_{l}$ are diagonal matrices.
+We choose $L=4$ and $d_{l}=25$.
+For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are matrices with i.i.d.
+standard Gassian entries and $\Lambda_{l}$ are diagonal matrices.
 For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the embedding layer; 3rd Query, 3rd Value, 3rd MLP (fc layer) in GPT2.
 Shifting and proportional scaling are performed to ensure all elements in $\Lambda_{l}$ lie in the interval $[1,5000]$.
-This ensures strong convexity and controls the condition number of $H$ equals $2000$.
-The spectra of $H_{l}$ are in Figure LABEL:fig:heter-block.
+This ensures strong convexity and controls the condition number of $H$ equals $5000$.
+The spectra of $H_{l}$ are in Figure LABEL:fig:heter-block in Appendix B.
+We choose $h=0$ for all cases.
   • Case 2: Hessian with CNN-type spectra.
-We consider the same setup as Case 1.
+We use the same setup as in Case 1.
 For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the 1st to 4th convolution layers in ResNet18.
-We shift and proportionally scale the elements $\Lambda_{l}$ to the interval $[1,2000]$ to ensure the strong convexity and the condition number $H$ equals 2000.
-The spectra of $H_{l}$ are shown in Figure LABEL:fig:homo-block.
+We then shift and scale $\Lambda_{l}$ to the interval $[1,5000]$ to ensure strong convexity and a condition number of 5000.
+The spectra of $H_{l}$ are shown in Figure LABEL:fig:homo-block in Appendix B.
   • Case 3: Hessian with simplified heterogeneous spectra.
@@
 For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are independent standard Gassian random matrix and $\Lambda_{l}$ are diagonal matrices.
-We set the diagonal elements of $\Lambda_{l}$ as $\{1,2,3\},\{99,100,101\},\{1998,1999,2000\}$ for $l=1,2,3$, respectively.
+We set the diagonal elements of $\Lambda_{l}$ as $\{1,2,3\},\{99,100,101\},\{4998,4999,5000\}$ for $l=1,2,3$, respectively.
 The spectra of $H_{l}$ are different due to their different supports.
-The condition number of Hessian $H$ is $2000$.
+The condition number of Hessian $H$ is $5000$.
   • Case 4: Hessian with simplified homogeneous spectra.
 We consider the same setup as Case 3.
-We set the diagonal elements of $\Lambda_{l}$ as $\{1,99,1998\},\{2,100,1999\},\{3,101,2000\}$ for $l=1,2,3$, respectively.
+We set the diagonal elements of $\Lambda_{l}$ as $\{1,99,4998\},\{2,100,4999\},\{3,101,5000\}$ for $l=1,2,3$, respectively.
 The spectra of $H_{l}$ are similar.
-The condition number of Hessian $H$ is 2000.
-Note that in all four cases, the Hessian share the same condition number.
-We study two types of optimizers: one that assigns a single learning rate for all blocks, and ones that assign different learning rates across blocks.
-The optimizers are introduced as follows.
+The condition number is 5000.
+Now we study two types of optimizers: one that assigns a single learning rate for all blocks, and one that assign different learning rates across blocks.
   • Single-learning-rate optimizer. We study gradient descent (GD).
     $w^{t+1}=w^{t}-\eta\nabla\mathcal{L}(w)=w^{t}-\eta(Hw^{t}-h)$ (1)
-    We use the optimal learning rate $\eta=\frac{2}{\mu+L}$ (Nesterov, 2013). We use a standard Gaussian random vector as initialization.
+    We use the optimal learning rate $\eta=\frac{2}{\mu+L}$ [63]. We use standard Gaussian initialization.
   • Coordinate-wise-learning-rate optimizer.
@@
 We use $\epsilon=0$.
-We consider $\beta_{2}=1$ and $\beta_{2}=0.999$, respectively.
+We consider $\beta_{2}=1$ and $\beta_{2}=0.99$, respectively.
 When $\beta_{2}=1$, Adam assigns coordinate-wise learning rates according to the initial gradient, but these learning rates are fixed along iteration.
@@
 When $\beta_{2}<1$, the coordinate-wise learning rates adaptively change along iteration.
-The update rule is as follows.
-    $w^{t+1}=w^{t}-\eta(D_{Adam}^{t})^{-1}\nabla\mathcal{L}(w)=w^{t}-\eta(D_{Adam}^{t})^{-1}(Hw^{t}-h),$ (3)
-    where $D_{Adam}^{t}=\operatorname{diag}\left((1-\beta_{2})\left(\sum_{k=1}^{t}\beta_{2}^{t-k}\nabla\mathcal{L}(w^{k})\circ\nabla\mathcal{L}(w^{k})\right)+\beta^{t}\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))\right)^{\frac{1}{2}}$ and $\nabla\mathcal{L}(w^{k})=Hw^{k}-h$.
-We grid search $\eta$ and use a standard Gaussian random vector as initialization.
-    We remark that when $\beta_{2}<1$, Adam would repeatedly bounce among non-optimal points, causing non-convergence. This will be shown in Proposition 2.
+The update rule is as follows (note that $\nabla\mathcal{L}(w^{k})=Hw^{k}-h$.).
+    $w^{t+1}=w^{t}-\eta(D_{Adam}^{t})^{-1}\nabla\mathcal{L}(w)=w^{t}-\eta(D_{Adam}^{t})^{-1}(Hw^{t}-h),\quad\text{where}$ (3)
+    $D_{Adam}^{t}=\operatorname{diag}\left((1-\beta_{2})\left(\sum_{k=1}^{t}\beta_{2}^{t-k}\nabla\mathcal{L}(w^{k})\circ\nabla\mathcal{L}(w^{k})\right)+\beta^{t}\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))\right)^{\frac{1}{2}}$
+    We grid search $\eta$ and use the standard Gaussian initialization. We remark that when $\beta_{2}<1$, Adam would bounce among non-optimal points. This will be shown in Proposition 2.
 Summary of experimental observations.
-We now compare the performance of Adam and GD on these four types of Hessians.
-The results are shown in Figure LABEL:fig_quadratic.
-For Hessian with heterogeneous blocks like Case 1 and 3, GD largely lags behind Adam.
-For Hessian with homogeneous blocks (Case 2 and 4), GD is on par or even better than Adam.
-We emphasize that all Hessians have the same condition number in these four cases.
+Figure LABEL:fig_quadratic presents two phenomena.
+For Hessian with heterogeneous blocks (Case 1 and 3), GD largely lags behind Adam.
+For Hessian with homogeneous blocks (Case 2 and 4), GD is on par with Adam.
+We emphasize that all Hessians have the same condition number.
 Further, Hessian of Case 3 and 4 share all the eigenvalues (not just the extreme ones).
-The performance gap between Adam and GD is due to the different blockwise spectra caused by the different locations of eigenvalues.
+The gap between Adam and GD is due to the different blockwise spectra caused by the different locations of eigenvalues.
 We hypothesize that GD performs badly because it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks.
@@
 To distinguish Adam and GD, we need to utilize more fine-grained quantities like blockwise spectra of sub-matrices.
-Unfortunately, the blockwise spectrum is rarely discussed in the optimization area.
-A most related notion is “block Lipschitz constant" (Beck and Tetruashvili, 2013) for studying block coordinate descent (BCD) type methods, but it was not linked to the performance of SGD or Adam before.
+Note that the blockwise spectrum is not common in the optimization area.
+The most related notion is perhaps “block Lipschitz constant" [8] for studying block coordinate descent (BCD) type methods, but it was not linked to the performance of SGD or Adam before.
 To our knowledge, we are not aware of any theory of Adam or GD built on the block diagonal structures or the blockwise spectra of Hessian.
-Further, there seems to be no result distinguishing the complexity of single-learning-rate methods and the coordinate-wise counterpart, either.
 We now make an initial attempt in this direction.
@@
 Proposition 1 shows that GD has complexity $\tilde{\mathcal{O}}(\kappa)$ and such complexity is tight.
-Now we prove that: if we are allowed to choose a customized learning rate for each block sub-matrix $H_{l}$, then the complexity will be improved.
-We call this method blockwise GD (BGD).
-Note that BGD is impractical to implement since it is hard to choose $L$ learning rates in advance.
-BGD serves as an artificial method for theoretical understanding.
-Blockwise GD (BGD).
-Choose learning rates $\{\eta_{l}\}_{l=1}^{L}$ for all blocks in advance and performs the following:
-$w^{t+1}=w^{t}-(D_{BGD})^{-1}\nabla\mathcal{L}(w)=w^{t}-(D_{BGD})^{-1}(Hw^{t}-h),$
-where $D_{BGD}=\operatorname{diag}\left(\left(\underbrace{\eta_{1},\cdots,\eta_{1}}_{d_{1}\text{times}},\cdots,\underbrace{\eta_{L},\cdots,\eta_{L}}_{d_{L}\text{times}}\right)^{T}\right)\in\mathbb{R}^{d\times d}.$
-BGD can be also viewed as a diagonally preconditioned version of GD. We now show that this diagonal preconditioner can significantly improve the complexity of GD.
+Now we prove that Adam can achieves better complexity.
+This is because it chooses different learning rates for different block sub-matrix $H_{l}$ via its diagonal preconditinoner $D_{Adam}^{0}$.
+We consider generic random initialization that covers commonly used distributions such as Gaussian, Uniform, etc.
+Assumption 1.
+(Random initialization.) Assume the initialization $w^{0}$ is sampled from a continuous distribution, i.e., the probability measure (induced by $w^{0}$) of any zero-Lebesgue-measure set is 0.
 Theorem 1.
-(Upper bound for GD with blockwise learning rate.) Consider $\min_{w}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$.
-We assume $H$ is a block diagonal matrix, i.e., $H=\operatorname{diag}(H_{1},\cdots,H_{L})$, where $H_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ and $d=\sum_{l=1}^{L}d_{l}$.
-Let $w_{BGD}^{t}$ be the output of BGD after $t$ steps.
-There exists a set of $\{\eta_{l}\}_{l=1}^{L}$ such that:
-$\mathcal{L}(w_{BGD}^{t+1})-\mathcal{L}^{*}\leq\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\left(\mathcal{L}(w_{BGD}^{t})-\mathcal{L}^{*}\right)$ (5)
-where $\kappa_{l}$ is the condition number of $H_{l}$.
-Theorem 1 shows that: if we are allowed to choose a learning rate for each $H_{l}$, the complexity could be $\tilde{O}(\max_{l\in[L]}\kappa_{l})$, in lieu of $\tilde{O}(\kappa)$ as in GD.
-That is to say, the rate of BGD only depends on the "slowest block".
-Note that in certain heterogeneous block Hessian such Case 1 and ${\bf 3}$, $\kappa_{l}$ is about 20×20$\times$20 × and 2000×2000$\times$2000 × smaller than $\kappa$, respectively.
-In these cases, blockwise-learning-rate methods can be about 20×20$\times$20 × and 2000×2000$\times$2000 × faster than single-learning-rate methods like GD, respectively.
-Despite the advantages of BGD, it is impractical to find the proper learning rates for all blocks, especially when the number of blocks $L$ is large.
-Fortunately, these learning rates could be approximated by the preconditioner in Adam, e.g., $D_{Adam}^{0}$ in (2).
-We now provide a complexity bound of Adam.
-Assumption 1.
-(Bounded initialization.) The initialization is bounded and is far from the optimal solution:
-$C_{l,1}\lambda_{l,1}\leq\left|[\nabla\mathcal{L}(w^{0})]_{l,i}\right|\leq C_{l,2}\lambda_{l,1},\text{ for }i\in[d_{l}],l\in[L].$
-Theorem 2.
-(Upper bound for Adam with $\beta_{2}=1$.) Consider the same setting as in Theorem 1 and consider Adam with $\beta_{1}=0$ and $\beta_{2}=1$ as in (2).
+(Upper bound for Adam with $\beta_{2}=1$.) Consider the same setting as Proposition 1 and consider Adam with $\beta_{1}=0$ and $\beta_{2}=1$ as in (2).
 Assume the initialization satisfies Assumption 1.
@@
 Let $\eta=\min_{l\in[L]}\frac{1}{C_{l,1}}$.
-We have
+Then w.p.1., we have
 $\mathcal{L}(w_{Adam}^{t+1})-\mathcal{L}^{*}\leq\max_{l\in[L]}\left(1-\frac{1}{\kappa_{Adam,l}}\right)\left(\mathcal{L}(w_{Adam}^{t})-\mathcal{L}^{*}\right)$ (5)
-where $\kappa_{Adam,l}=r\kappa_{l}$, $r:=\max_{l\in[L]}\left(\frac{C_{l,2}}{C_{l,1}}\right)^{2}$ and $\kappa_{l}$ is the condition number of $H_{l}$.
-The proofs of all the above theorems are shown in Appendix E.
+where $\kappa_{Adam,l}=r\kappa_{l}$, $\kappa_{l}$ is the condition number of $H_{l}$, constant $r$ relates to $w^{0}$ defined as:
+$r=\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}},\text{ where }C_{l,1}=\min_{i\in[d_{l}]}\frac{|[\nabla\mathcal{L}(w^{0})]_{l,i}|}{\lambda_{l,1}},C_{l,2}=\max_{i\in[d_{l}]}\frac{|[\nabla\mathcal{L}(w^{0})]_{l,i}|}{\lambda_{l,1}}.$ (6)
+The proofs of the above theorems are shown in Appendix E.
 Theorem 1 states that Adam (with $\beta_{2}=1$) has complexity $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$.
-We remark that condition $\beta_{2}=1$ is necessary because any $\beta_{2}<1$ causes non-convergence issue (Da Silva and Gazeau, 2020; Bock and Weiß, 2019).
+We note that coefficient $r$ depends on the ratio between initial gradient and the principal eigenvalue for each block, and smaller ratio would give faster convergence.
+We further remark that condition $\beta_{2}=1$ is necessary because any $\beta_{2}<1$ causes non-convergence issue [10, 21].
 We restate their results in Proposition 2.
@@
 As such, $\beta_{2}=1$ is necessary for asymptotic analysis.
-As shown in (Da Silva and Gazeau, 2020), the non-convergence is due to the constant learning rate.
+The analysis for $\beta_{2}=1$ is still meaningful since it still shows the effect of Adam’s preconditioner.
+As shown in [21], the non-convergence is due to the constant learning rate.
+Reducing the learning rate reduces the gap between $\mathcal{L}(w_{Adam}^{t})$ and $\mathcal{L}^{*}$, but does not remove it.
 Proposition 2.
-(Non-convergence of constant-learning-rate Adam with $\beta_{2}<1$.) (Da Silva and Gazeau, 2020, Proposition 12) Consider $\min_{w\in\mathbb{R}}\mathcal{L}(w)=\frac{1}{2}w^{2}$.
+(Non-convergence of constant-learning-rate Adam with $\beta_{2}<1$.) [21, Proposition 12, Figure 1] Consider $\min_{w\in\mathbb{R}}\mathcal{L}(w)=\frac{1}{2}w^{2}$.
 Consider Adam with $\beta_{1}=0$ and $\beta_{2}<1$ as in (3).
 Let $w_{Adam}^{t}$ be the output of Adam after $t$ steps.
-There exists a discrete limit cycle for (3), i.e., Adam produces oscillations and ${\small\liminf_{t\rightarrow\infty}\left(\mathcal{L}(w_{Adam}^{t})-\mathcal{L}^{*}\right)>0}$.
+There exists a discrete limit cycle for (3) and $\liminf_{t\rightarrow\infty}\left(\mathcal{L}(w_{Adam}^{t})-\mathcal{L}^{*}\right)>0$.
 We now compare the complexity of Adam and that of GD.
@@
 In the quadratic model with heterogeneous blocks (Case 3), our simulation over 1000 trials shows that $r\leq 1000$ with probability $\geq\frac{2}{3}$ when using standard Gaussian random initialization.
-Since $\max_{l\in[L]}\kappa_{l}\approx 1$, we have $r\cdot\max_{l\in[L]}\kappa_{l}\leq 100$, w.h.p., and is 20×20$\times$20 × smaller than $\kappa=2000$.
-So Adam could be 20×20$\times$20 × faster than GD, w.h.p..
+Since $\max_{l\in[L]}\kappa_{l}\approx 1$, we have $r\cdot\max_{l\in[L]}\kappa_{l}\leq 1000$, w.h.p., and is about 5×5$\times$5 × smaller than $\kappa=5000$.
+So Adam could be 5×5$\times$5 × faster than GD, w.h.p..
 This is indeed observed in Figure LABEL:fig_quadratic where Adam outperforms GD by a significant margin.
-We summarize the complexity of GD, BGD, and Adam in Table 2.
-Optimizer GD BGD Adam with Adam with
-$\beta_{1}=0$ and $\beta_{2}=1$ (2) $\beta_{1}=0$ and $\beta_{2}<1$ (3)
-Complexity $\tilde{O}(\kappa)$ $\tilde{O}(\max_{l\in[L]}\kappa_{l})$ $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$
-Table 2: The complexity of GD, BGD and Adam for minimizing a strongly convex quadratic function with block diagonal Hessian.
+We summarize the complexity of GD and Adam in Table 2.
+We remark that there is room to improve the complexity bound of Adam, but it could be challenging.
+We provide more technical discussions in Appendix B.
+Table 2: The complexity of GD and Adam for minimizing a strongly convex quadratic function with block diagonal Hessian.
 The symbol ✗ means non-convergence.
 $\kappa$ and $\kappa_{l}$ denote the condition number of the full Hessian and the block submatrix, respectively.
-$r:=\max_{l\in[L]}\left(\frac{C_{l,2}}{C_{l,1}}\right)^{2}$ and $C_{l,1}$, $C_{l,2}$ are constants in Assumption 1.
+$r$ is defined in (6).
+Optimizer GD Adam with Adam with
+$\beta_{1}=0$ and $\beta_{2}=1$ (2) $\beta_{1}=0$ and $\beta_{2}<1$ (3)
+Complexity $\tilde{O}(\kappa)$ $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$
+4 Conclusion
+In this work, we explore why SGD performs badly on Transformers.
+By numerically exploring various neural nets and quadratic problems, we establish a phenomenon called block heterogeneity in Hessian and link it to the deficiency of SGD.
+Initial theory is provided to support the claim.
+Appendix A Related Works
+On the deficiency of SGD on Transformers
+There is an active line of works that explores why SGD performs badly on Transformers.
+One representative hypothesis is that SGD cannot handle the heavy-tailed stochastic noise in language tasks [106].
+However, Chen et al.
+[14], Kunstner et al.
+[46] reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons.
+Further, SGD performs poorly on Vision Transformers on ImageNet (See Figure LABEL:fig:nlp_figure.
+Also see [93] for more evidence), so the data modality (e.g., language or vision tasks) might not be as crucial as the architecture.
+[105] showed that NLP tasks have “unbounded smoothness" issue and SGD with gradient clipping performs better than SGD in this case.
+Although clipping is an effective trick, we still observe a huge gap between clipped SGD and Adam  [^4] For all NLP tasks, clipping is performed immediately after backpropagation.
+So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD.
+.
+So there might be other reasons that hamper SGD.
+Different from these works, we find SGD performs badly because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
+Understanding of Adam.
+There was once a long-standing debate on the possible divergence of Adam [72].
+The convergence for the unmodified versions is later established in [78, 108] for RMSprop and Adam.
+More convergence analyses of general adaptive gradient methods are listed later in this section.
+We here focus on the literature that explores the benefit of Adam.
+Xie et al.
+[94] show that Adam can help avoid saddle points, which is an orthogonal direction to this work.
+Wang et al.
+[85], Crawshaw et al.
+[19], Li et al.
+[50] show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time).
+We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes performs badly while Adam works well (Figure LABEL:fig_quadratic).
+This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate (over time), and there are other reasons for Adam’s success.
+We show that an important benefit of Adam is its ability to handle the heterogeneity across blocks (over space).
+Recent works [9, 92, 46, 53, 3] build a relation between Adam and the sign-based methods.
+Wu et al.
+[92] further showed that sign-based methods can be effective when the Hessian is diagonal and satisfies several other properties.
+However, as put by the authors, it seems “unclear to what extent these properties hold for real problems".
+Pan and Li [66] numerically found that the Adam can reduce the directional sharpness along trajectories, while its relation to fast convergence remains mysterious.
+A recent work [43] point out that Adam biases the trajectories towards regions where Hessian has “uniform diagonal entries" while SGD cannot.
+The distribution of Hessian diagonal entries is also investigated in [53].
+The theory in [43] implies that Adam is faster when the Hessian is diagonal.
+However, as argued above, it is unclear whether the diagonal Hessian structure commonly holds in real problems.
+In fact, we find the Hessian is closer to a block-diagonal (instead of pure diagonal) structure on some small Transformers.
+In these cases, blockwise eigenvalues carry more information than diagonal entries, providing extra details such as the location of eigenvalues.
+We find that these extra details are important for distinguishing Adam and SGD.
+Hessian Spectrum Analysis.
+There are several important attempts to explore the Hessian spectrum of MLPs and CNNs.
+Early works [75, 76, 12] found that the Hessian spectra of MLPs and CNNs consist of a “bulk" together with a few “outliers".
+Papyan [69], Wu et al.
+[92], Liao and Mahoney [51] further characterized the bulks and outliers in theory.
+Papyan [67, 68] numerically built the relation between these "outliers" and the Gauss-Newton matrix.
+Sankar et al.
+[77] numerically explored the relation between Hessian of CNNs and Gauss-Newton matrix in each layer.
+They further found that most CNN layers contribute similarly to the overall loss surface.
+We find that this result is restricted to CNNs and does not hold on Transformers due to the heterogeneity.
+Gur-Ari et al.
+[39] showed that for MLPs and CNNs, gradient descent converges to a small subspace spanned by a few top eigenvectors of the Hessian.
+Yao et al.
+[98], Zhang et al.
+[104] explored the relation between the Hessian spectrum of CNNs and some training phenomena such as the effect of batch sizes.
+Ghorbani et al.
+[32], Yao et al.
+[99] focused on explaining the effectiveness of techniques such as BatchNorm.
+Note that all these works are restricted to MLPs and CNNs, while we study the Hessian of Transformers (in addition to CNNs and MLPs) as well as its impacts on different optimizers.
+On the difficulties of Transformer training.
+Transformers are known to be difficult to train.
+Researchers have attributed the training difficulties to various phenomena in different components of Transformers, including: the logits divergence or the rank degeneracy in the outputs of attention layers [26, 65, 91, 102, 24, 17]; the growth of parameter norm in attention layers [61]; over-reliance on residue branches [55]; and some negative impact of layer norm [15, 103, 41].
+These phenomena have a strong correlation with gradient vanishing or explosion in Transformers [103, 55, 41, 95, 65, 88, 91, 62], which leads to training difficulties.
+Several solutions have been proposed.
+Liu et al.
+[55] numerically observed that adaptive gradient methods can (partly) overcome gradient vanishing by giving “consistent update magnitude", while it seems unclear how consistent update magnitude would help optimization in principle.
+Researchers further develop training tricks such as warmup learning rate [54, 95], temperature scaling [65], better initialization [103, 41, 88, 5, 97], and variants of Layer Norm [64, 89, 95, 88, 24].
+Recent researchers also suggest using z-loss regularization [17, 96] and tuning hyperparameters of Adam [108, 91].
+All these tricks can help mitigate gradient explosion or vanishing.
+Nevertheless, training large-scale Transformers remains challenging [107, 101, 91, 62, 17].
+Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian.
+We establish a strong correlation between the blockwise Hessian spectra of Transformers and the poor performance of SGD.
+We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
+Convergence analysis of general adaptive gradient methods
+There is extensive convergence analysis for adaptive gradient methods.
+For instance, researchers study the convergence of AMSGrad [72, 109], RMSprop [100], AdaFom [16], AdaBound [57], and Adam with iterate-dependent hyperparameters [110, 13, 31].
+The convergence of Adam is also explored in [23, 86].
+There is also an active line of theoretical research on the convergence of AdaGrad [28], we recommend [87] for more detailed introduction.
+In this work, we do not focus on the convergence analysis.
+Rather, we explore the quantitative difference between the loss landscape of CNNs and Transformers and how it impact the behaviors of SGD and Adam.
+Appendix B More Results and Discussions
+Performance comparison of AdamW and SGD on different Architectures.
+Here, we show the performance comparison of AdamW and SGD on different models.
+All the vision models are trained on ImageNet.
+Language models are trained on different English corpus.
+See Appendix D.1 for more implementation details.
+Block-diagonal structure in the existing literature.
+We remark that Collobert [18], Roux et al.
+[73], Martens and Grosse [60] also observed the block-diagonal structure in (approximated) Hessian of small-scaled MLPs.
+We restate their findings in Figure LABEL:fig_block_diagnal_existing.
+[18, Section 7] further theoretically proved that the block diagonal structure stems from (i) the layer-by-layer structure and (ii) the Cross-Entropy loss.
+These results suggest that block-diagonal structure might be common in NNs.
+Blockwise spectra for quadratic models in Section 3.1.
+We here visualize the blockwise spectrum for the quadratic models in Case 1 and Case 2. These spectra are collected from GPT2 and ResNet18, respectively.
 How to obtain a tighter complexity bound of Adam?
-It is valid to ask whether the complexity upper bound $\kappa_{Adam,l}=r\kappa_{l}$ can be tightened, e.g., remove the factor of $r$.
+It is valid to ask whether the complexity upper bound in Theorem 1 $\kappa_{Adam,l}=r\kappa_{l}$ can be tightened, e.g., improve the factor of $r$.
 We point out it would be difficult if there is no extra structure on $H_{l}$.
 A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$.
-Intuitively, a diagonal preconditioner of $H_{l}$ is expected to be powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant.
-Unfortunately, it is unclear whether these structures hold in real Transformers.
+Intuitively, a diagonal preconditioner of $H_{l}$ is powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant [30].
+Unfortunately, it is unclear whether these structures hold in Transformers.
 Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number.
-For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $3.9\kappa_{1}$, $14.1\kappa_{2}$, $4.2\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials).
-It would be interesting to explore if there are special structures of $H_{l}$ in real Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it.
+For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $7.09\kappa_{1}$, $18.98\kappa_{2}$, $18.76\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials).
+It would be interesting to explore if there are special structures of $H_{l}$ in Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it.
 We leave it as a future direction.
@@
 In summary, our theory indicates that: for problems with block heterogeneity, the single-learning rate methods like GD can largely lag behind coordinate-wise learning rate methods like Adam.
-4 Related Works
-On The failure of SGD on Transformers
-There is extensive literature studying the difficulties of Transformer training.
-We summarize these works in Appendix A.
-We here focus on the literature that explores why SGD fails on Transformers.
-One representative hypothesis is that SGD fails because it cannot handle the heavy-tailed stochastic noise in language tasks (Zhang et al., 2020).
-However, Chen et al.
-(2021); Kunstner et al.
-(2023) reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons.
-Further, SGD performs poorly on Vision Transformers on ImageNet (Figure LABEL:fig:nlp_figure), so the data modality (e.g., language or image tasks) might not be as crucial as the architecture.
-Zhang et al.
-(2019c) showed that Transformers have the “unbounded smoothness" issue and SGD with gradient clipping performs better than SGD in this case.
-Although clipping is an effective trick, it does not save SGD from failure as we still observe a huge gap between clipped SGD and Adam  [^3] For all NLP tasks, clipping is performed after backpropagation.
-So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD.
-.
-So there might be other reasons that hamper SGD.
-Different from these works, we find SGD fails because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
-Understanding of Adam.
-There was once a long-standing debate on the possible divergence of Adam (Reddi et al., 2018).
-The convergence for the unmodified versions is later established in (Shi et al., 2020; Zhang et al., 2022b) for RMSprop and Adam.
-More convergence analyses of general adaptive gradient methods are listed in Appendix A.
-We here focus on the literature that explores the benefit of Adam.
-Xie et al.
-(2022) show that Adam can help avoid saddle points, which is an orthogonal direction to this work.
-Wang et al.
-(2022a); Crawshaw et al.
-(2022); Li et al.
-(2023) show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time).
-We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes fails while Adam works well (Figure LABEL:fig_quadratic).
-This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate over time, and there are other reasons for Adam’s success.
-We show that an important benefit of Adam is its ability to handle the heterogeneity across blocks (over space).
-Bernstein et al.
-(2018); Wu et al.
-(2020); Kunstner et al.
-(2023); Liu et al.
-(2023) build a relation between Adam and the sign-based methods.
-Wu et al.
-(2020) further showed that sign-based methods can be effective when the Hessian is diagonal and satisfies several other properties.
-However, as put by the authors, it seems “unclear to what extent these properties hold for real problems".
-Pan and Li (2023) numerically found that the Adam can reduce the directional sharpness along trajectories, while its relation to fast convergence remains mysterious.
-A recent work (Jiang et al., 2023) point out that Adam biases the trajectories towards regions where Hessian has “uniform diagonal entries" while SGD cannot.
-The distribution of Hessian diagonal entries is also investigated in (Liu et al., 2023).
-The theory in (Jiang et al., 2023) implies that Adam is faster when the Hessian is diagonal.
-However, as argued above, it is unclear whether the diagonal Hessian commonly holds in real problems.
-In fact, we find the Hessian is closer to a block-diagonal (instead of pure diagonal) structure on some small Transformers.
-In these cases, blockwise eigenvalues carry more information than diagonal entries, providing extra details such as the location of eigenvalues.
-We find that these extra details are important for distinguishing Adam and SGD.
-Hessian Spectrum Analysis.
-There are several important attempts to explore the Hessian spectrum of MLPs and CNNs.
-Sagun et al.
-(2016, 2017); Chaudhari et al.
-(2019) found that the Hessian spectra of MLPs and CNNs consist of a “bulk" together with a few “outliers".
-Papyan (2020); Wu et al.
-(2020); Liao and Mahoney (2021) further characterized the bulks and outliers in theory.
-Papyan (2018, 2019) numerically built the relation between these "outliers" and the Gauss-Newton matrix.
-Sankar et al.
-(2021) numerically explored the relation between Hessian of CNNs and Gauss-Newton matrix in each layer.
-They further showed that most CNN layers contribute similarly to the overall loss surface.
-We find that this result is restricted to CNNs and does not hold on Transformers due to the heterogeneity.
-Gur-Ari et al.
-(2018) showed that for MLPs and CNNs, gradient descent converges to a small subspace spanned by a few top eigenvectors of the Hessian.
-Yao et al.
-(2018); Zhang et al.
-(2019b) explored the relation between the Hessian spectrum of CNNs and some training phenomena such as the effect of batch sizes.
-Ghorbani et al.
-(2019); Yao et al.
-(2020) focused on explaining the effectiveness of techniques such as BatchNorm.
-Note that all these works are restricted to MLPs and CNNs, while we study the Hessian of Transformers (in addition to CNNs and MLPs) as well as its impacts on different optimizers.
-5 Conclusion
-In this work, we investigate the failure mode of SGD on Transformers through the lens of Hessian.
-By numerically explore various Transformers, CNNs, MLPs and quadratic problems, we establish a phenomenon called block heterogeneity in Hessian and link it to the failure of SGD.
-We point out that SGD fails because it applies one single learning rate for all blocks, which cannot handle the heterogeneity among blocks.
-Initial theory are provided to support the claim.
-Appendix A More Related Works
-On the difficulties of Transformer training.
-Transformers are known to be difficult to train.
-Researchers have attributed the training difficulties to various phenomena in different components of Transformers, including: the logits divergence or the rank degeneracy in the outputs of attention layers [Dong et al., 2021, Noci et al., 2022, Wortsman et al., 2023, Zhai et al., 2023, Dehghani et al., 2023, Chowdhery et al., 2023]; the growth of parameter norm in attention layers [Merrill et al., 2020]; over-reliance on residue branches [Liu et al., 2020]; and some negative impact of layer norm [Chen et al., 2018, Zhang et al., 2019a, Huang et al., 2020].
-These phenomena have a strong correlation with gradient vanishing or explosion in Transformers [Zhang et al., 2019a, Liu et al., 2020, Huang et al., 2020, Xiong et al., 2020, Noci et al., 2022, Wang et al., 2022b, Wortsman et al., 2023, Molybog et al., 2023], which leads to training difficulties.
-Several solutions have been proposed.
-Liu et al.
-[2020] numerically observed that adaptive gradient methods can (partly) overcome gradient vanishing by giving “consistent update magnitude", while it seems unclear how consistent update magnitude would help optimization in principle.
-Researchers further develop training tricks such as warmup learning rate [Liu et al., 2019, Xiong et al., 2020], temperature scaling [Noci et al., 2022], better initialization [Zhang et al., 2019a, Huang et al., 2020, Wang et al., 2022b, Bachlechner et al., 2021, Yang et al., 2022], and variants of Layer Norm [Nguyen and Salazar, 2019, Wang et al., 2019, Xiong et al., 2020, Wang et al., 2022b, Dehghani et al., 2023].
-Recent researchers also suggest using z-loss regularization [Chowdhery et al., 2023, Yang et al., 2023] and tuning hyperparameters of Adam [Zhang et al., 2022b, Wortsman et al., 2023].
-All these tricks can help mitigate gradient explosion or vanishing.
-Nevertheless, training large-scale Transformers remains challenging [Zhang et al., 2022a, Zeng et al., 2022, Wortsman et al., 2023, Molybog et al., 2023, Chowdhery et al., 2023].
-Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian.
-We establish a strong correlation between the blockwise Hessian spectra of Transformers and the failure of SGD.
-We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
-Convergence analysis of adaptive gradient methods
-There is extensive convergence analysis for adaptive gradient methods.
-For instance, researchers study the convergence of AMSGrad [Reddi et al., 2018, Zhou et al., 2018], RMSprop [Zaheer et al., 2018], AdaFom [Chen et al., 2019], AdaBound [Luo et al., 2018], and Adam with iterate-dependent hyperparameters [Zou et al., 2019, Chen et al., 2022, Gadat and Gavra, 2022].
-The convergence of Adam is also explored in [Défossez et al., 2022, Wang et al., 2023a].
-There is also an active line of theoretical research on the convergence of AdaGrad [Duchi et al., 2011], we recommend [Wang et al., 2023b] for more detailed introduction.
-In this work, we do not focus on the convergence analysis.
-Rather, we explore the quantitative difference between the loss landscape of CNNs and Transformers and how it impact the behaviors of SGD and Adam.
-Appendix B More Preliminaries
-B.1 Preliminaries on Optimizers
+Appendix C More Preliminaries
+C.1 Preliminaries on Optimizers
 […]
@@
 3:     Uniformly sample $\tau^{t}$ from the index set $\{1,2,\cdots,n\}$
-4:     $m^{t}=\beta_{1}m^{t}+\nabla f_{\tau^{t}}(x^{t})$
+4:     $m^{t}=\beta_{1}m^{t}+\nabla\mathcal{L}_{\tau^{t}}(x^{t})$
 5:     $x^{t+1}=x^{t}-\eta_{t}m^{t}$
@@
 7:  end for
-B.2 Preliminaries on the Stochastic Lanczos Quadrature Method
+C.2 Preliminaries on the Stochastic Lanczos Quadrature Method
 Additional notations.
@@
 We use $\|\cdot\|_{2}$ to denote the Euclidean norm.
-Approximation of the Hessian spectrum can be formulated as a trace estimation problem, as introduced in [Lin et al., 2016, Ubaru et al., 2017].
+Approximation of the Hessian spectrum can be formulated as a trace estimation problem, as introduced in [52, 83].
 First, the spectrum (eigenvalue density) of Hessian $H$ can written as: $\phi(t)=\frac{1}{d}\sum_{i=1}^{d}\delta\left(t-\lambda_{i}\right)$, where $\lambda_{i}$ are the eigenvalues of $H$ and $\delta$ is the Dirac $\delta$-function.
@@
 As such, spectrum approximation could be formulated as a trace estimation problem, i.e., estimating $\frac{1}{d}\operatorname{tr}(f(H))$, where $H\in\mathbb{R}^{d\times d}$ is a real symmetric matrix.
-Trace estimation problems could be solved efficiently by the Stochastic Lanczos Quadrature Method (SLQ) [Golub and Strakoš, 1994].
-For the ease of readers, we provide a detailed description of SLQ in our context.
+Trace estimation problems could be solved efficiently by the Stochastic Lanczos Quadrature Method (SLQ) [35].
+For the ease of readers, we re-organize and summarize the existing literature ([35, 83, 32]) and provide a detailed description of SLQ in our context.
 SLQ consists of the following steps.
@@
 We Approximate the trace of matrix function as $\frac{1}{d}tr(f(H))=\mathbb{E}(v^{T}f(H)v)\approx\frac{1}{n_{v}}\sum_{i}^{n_{v}}v_{i}^{T}f(H)v_{i}$, where $v=u/\|u\|_{2}$ and $u$ is a Rademacher random vector (each entry of $u$ independently takes ±$\pm 1$ with probability $1/2$).
-This step is called Huchinson’s estimation [Hutchinson, 1989].
-Note that we can also replace the Rademacher random vector $u$ by a unit Gaussian vector (i.e., $u\sim N(0,I_{d\times d})$) and the unbiasedness still holds [Avron and Toledo, 2011].
+This step is called Huchinson’s estimation [42].
+Note that we can also replace the Rademacher random vector $u$ by a unit Gaussian vector (i.e., $u\sim N(0,I_{d\times d})$) and the unbiasedness still holds [4].
 In our implementation, we sample $u\sim N(0,I_{d\times d})$ because there is an efficient built-in PyTorch function for generating Gaussian vectors.
@@
 Step 2-1.
-We rewrite $v^{T}f(H)v$ as a Riemann-Stieltjes integral [Golub and Meurant, 2009]:
+We rewrite $v^{T}f(H)v$ as a Riemann-Stieltjes integral [34]:
 […]
@@
 We hope to design $\{(c_{j},x_{j})\}_{j=1}^{m}$ with a reasonable number of $m$ such that the estimation error is small.
-Fortunately, the Gaussian Quadrature method provides a generic design principle of $\{(c_{j},x_{j})\}_{j=1}^{m}$ [Golub and Meurant, 2009, Epperson, 2013].
+Fortunately, the Gaussian Quadrature method provides a generic design principle of $\{(c_{j},x_{j})\}_{j=1}^{m}$ [34, 29].
 It is proved that: when $f(\lambda)$ is not "too complicated" (e.g.
@@
 Theorem 2 shows the existence of good $\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ and their general form.
-In fact, it is also shown that Gaussian Quadrature is optimal: no other$\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ can achieve zero approximation error for higher degree polynomials $f(\lambda)$ [Golub and Meurant, 2009].
-However, it is often difficult to find these quadrature rules [Golub and Welsch, 1969].
+In fact, it is also shown that Gaussian Quadrature is optimal: no other$\left\{\left(c_{j},x_{j}\right)\right\}_{j=1}^{m}$ can achieve zero approximation error for higher degree polynomials $f(\lambda)$ [34].
+However, it is often difficult to find these quadrature rules [36].
 There are at least three questions in sequel:
@@
 We first answer question 2) and 3) and leave question 1) for later discussion.
-Now suppose that we have found the orthogonal polynomials $\left\{p_{k}(x)\right\}_{k=1}^{m}$ w.r.t.
-$\mu$.
-Recall that any orthogonal polynomial has the following "three-term" recursion [Golub and Meurant, 2009]:
+Now suppose that we have found the orthogonal polynomials $\left\{p_{k}(x)\right\}_{k=1}^{m}$ w.r.t. $\mu$. Recall that any orthogonal polynomial has the following "three-term" recursion [34]:
 $p_{k+1}(x)=\left(x-\alpha_{k+1}\right)p_{k}(x)-\beta_{k}p_{k-1}(x),k=0,1,\ldots,$
@@
 Theorem 3.
-[Golub and Meurant, 2009] For the Gaussian Quadrature, $\{x_{j}\}_{j=1}^{m}$ are the eigenvalues of $J_{m}$ and $\{c_{j}\}_{j=1}^{m}$ are the squares of the first elements of the normalized eigenvectors of $J_{m}$.
-The proof of Theorem 4 is based on Christoffel-Darboux relation [Brezinski, 1990].
+[34] For the Gaussian Quadrature, $\{x_{j}\}_{j=1}^{m}$ are the eigenvalues of $J_{m}$ and $\{c_{j}\}_{j=1}^{m}$ are the squares of the first elements of the normalized eigenvectors of $J_{m}$.
+The proof of Theorem 3 is based on Christoffel-Darboux relation [11].
 Now, the remaining question is: how to find the Jacobian matrix $J_{m}$ of a sequence of orthogonal polynomials w.r.t.
@@
 Step 3.
-When $\mu$ is specified as in (8), $J_{m}$ can be exactly found in $m$ steps using the Lanczos algorithm [Lanczos, 1950], as shown in Algorithm 4.
+When $\mu$ is specified as in (8), $J_{m}$ can be exactly found in $m$ steps using the Lanczos algorithm [47], as shown in Algorithm 4.
 This method takes a real symmetric matrix as input and returns a tridiagonal matrix.
@@
 Later, researchers found a deep connection between the Lanczos algorithm and orthogonal polynomials, which further connects this method to the Gaussian quadrature.
-The method (of finding the Gaussian quadrature by the Lanczos algorithm) is called the Lanczos quadrature [Golub and Strakoš, 1994, Bai and Golub, 1996, Golub and Meurant, 2009].
+The method (of finding the Gaussian quadrature by the Lanczos algorithm) is called the Lanczos quadrature [35, 6, 34].
 An extremely elegant but highly nontrivial result is as follows:
 Theorem 4.
-[Golub and Meurant, 2009] Given a real symmetric matrix $H\in\mathbb{R}^{d\times d}$ and an arbitrary vector $v\in\mathbb{R}^{d}$ with unit Euclidean norm, we define the measure $\mu$ as in (8) based on this $H$ and $v$.
+[34] Given a real symmetric matrix $H\in\mathbb{R}^{d\times d}$ and an arbitrary vector $v\in\mathbb{R}^{d}$ with unit Euclidean norm, we define the measure $\mu$ as in (8) based on this $H$ and $v$.
 Then $m$ steps of the Lanzcos algorithm return the Jacobian matrix $J_{m}$ of orthogonal polynomials w.r.t.
@@
 Ubaru et al.
-[2017] show that SLQ converges faster than any other polynomial expansion method for spectrum estimation (e.g., Chebyshev methods used in [Adams et al., 2018]).
-See [Ubaru et al., 2017, Theorem 4.1] for a formal statement.
+[83] show that SLQ converges faster than any other polynomial expansion method for spectrum estimation (e.g., Chebyshev methods used in [2]).
+See [83, Theorem 4.1] for a formal statement.
 We remark that there are at least four versions of the Lanczos algorithm in Step 3.
-Here, we adopt the version in Algorithm 4 since it is known to be the most numerically stable version [Cullum and Willoughby, 2002, Saad, 2011, Wikipedia, 2023].
+Here, we adopt the version in Algorithm 4 since it is known to be the most numerically stable version [20, 74, 90].
 Throughout this work, we choose $f\left(\cdot\right)$ as the Gaussian blurring function $f(\lambda):=\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(t-\lambda)^{2}}{2\sigma^{2}}\right)$ for spectrum approximation.
@@
 8:  Return $\frac{1}{\texttt{num}_{v}}\sum_{i=1}^{\texttt{num}_{v}}f\left(\ell_{i};t,\sigma^{2}\right)$
-Appendix C More Eperimental Details
-C.1 Implementation Details on SLQ
+Appendix D More Eperimental Details
+D.1 Implementation Details on SLQ and Training Configurations
 Implementation and Running Time Analysis.
 We provide a simple PyTorch implementation of SLQ.
-The only query SLQ makes to the neural network is the Hessian vector product, which is attained using the auto-differentiation framework [Pearlmutter, 1994].
+The only query SLQ makes to the neural network is the Hessian vector product, which is attained using the auto-differentiation framework [70].
 To assure the accuracy of the Lanczos algorithm, we remove all the randomness in the forward and backward passes, including: data shuffling order, data augmentation, and dropout, etc..
-Since Flash Attention [Dao et al., 2022] does not support the calculation of Hessian-vector product, we implement all attention blocks in the naive way.
+Since Flash Attention [22] does not support the calculation of Hessian-vector product, we implement all attention blocks in the naive way.
 For the calculation of the blockwise Hessian spectrum $\nabla^{2}\mathcal{L}(w_{l})$, we sample $u_{l}\sim N(0,I_{d_{l}\times d_{l}})$ and set $v_{l}=u_{l}/\|u_{l}\|_{2}\in\mathbb{R}^{d_{l}}$.
@@
 $\sigma$ is tuned based on visual effects.
-These hyperparameters are reported to reach highly accurate estimation with error $<10^{-14}$ [Ghorbani et al., 2019].
+These hyperparameters are reported to reach highly accurate estimation with error $<10^{-14}$ [32].
 We now briefly discuss the computational cost of SLQ.
@@
 Recall $\nabla^{2}\mathcal{L}(w)d=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}\mathcal{L}_{i}(w)d$, so each Hessian-vector product operation requires (i) calculating $\nabla^{2}\mathcal{L}_{i}(w)d$; (ii) repeating (i) on all data.
-We point out that (i) can be computed efficiently and precisely with just two backpropagation passes [Pearlmutter, 1994].
+We point out that (i) can be computed efficiently and precisely with just two backpropagation passes [70].
 The major computational bottleneck lies in (ii) due to the large $n$.
@@
 More detailed setup of SLQ are shown as follows.
-  • ResNet18 (20M) and VGG16 (138M) on ImageNet.
+  • ResNet18 (18M) and VGG16 (138M) on ImageNet.
 We use the code base of PyTorch Examples  [^5] https://github.com/pytorch/examples/blob/main/imagenet/main.py.
@@
 In total, it takes about 12 hours on one V100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
-  • GPT2 (125M) on Openwebtext.
+  • GPT2 (125M) on Openwebtext  [^9] https://huggingface.co/datasets/Skylion007/openwebtext.
 We use the code base of NanoGPT.
@@
 In total, it takes about 7 days on one A100 GPU to estimate all the blockwise Hessian spectra and the full Hessian spectrum.
-We train all the models under the default configurations in all cases.
-For CNN training, we follow the rule of thumb to choose the learning rate of AdamW: we set the learning rate of AdamW to be 100×100$\times$100 × smaller than the recommended learning rate of SGD.
-C.2 Implementation Details on Figure LABEL:fig_block_diagnal
-We use the code base of NanoGPT to train a decoder-only Transformers on 4 consecutive tokens randomly selected from Openwebtext.
+Training configuration.
+In all cases, we train all the models under the default configurations in the above codebase.
+We grid-search the learning rates for SGD and Adam under the same budget and report the best result for each optimizer.
+We use the cosine-decay learning rate schedule for vision tasks.
+For SFT task, we use nanoGPT codebase.
+We first pre-train GPT2 on OpenwebText for 25B tokens and then fine-tune it on a subset of Alpaca Eval  [^10] https://huggingface.co/datasets/tatsu-lab/alpaca_eval.
+D.2 Implementation Details on Figure LABEL:fig_block_diagnal
+We use the code base of NanoGPT to train decoder-only Transformers on 4 consecutive tokens randomly selected from Openwebtext.
 We set the model configuration as: context window = 2, number of heads = 2 and the embedding dimension = 4.
 In MLP layers, all widths equal to 4.
-We change the number of layers from 1 to 4.
+We choose the number of layers to be 2, 4, and 8.
 We remove all the LayerNorms in the model.
@@
 We compute the Hessian on all the parameters blocks in attention and MLP layers.
-In Figure LABEL:fig_block_diagnal (a), the variables in three blue boxes corresponds to the parameters in Query and Key; Value and Projection; and MLP, respectively.
+In Figure LABEL:fig_block_diagnal (a), the variables each block corresponds to the parameters in Query and Key; Value and Projection; and MLP, respectively.
+For better visualization, We take the absolute value of each entry to distinguish non-zero values (including negatives) from those near 0 and then report the average value in each block.
 Similarly for the rest of the figures.
@@
 Rigorously speaking, so far we have not gotten sufficient evidence to claim this structure commonly holds in larger Transformers.
-It requires new numerical methods to efficiently check the block-diagonal Hessian structure in larger models without explicitly calculating them.
+It requires new numerical methods to efficiently check the block-diagonal Hessian structure without explicitly calculating them.
 We leave it as an interesting future direction.
-For completeness, we remark that Collobert [2004], Roux et al.
-[2007], Martens and Grosse [2015] also observed the block-diagonal structure in (approximated) Hessian of small-scaled MLPs.
-We restate their findings in Figure LABEL:fig_block_diagnal_existing.
-[Collobert, 2004, Section 7] further theoretically proved that the block diagonal structure stems from (i) the layer-by-layer structure and (ii) the Cross-Entropy loss.
-These results suggest that block-diagonal structure might be common in NNs.
-C.3 Implementation Details on the MLP experiments in Figure LABEL:fig_mlp_gap
+D.3 More Details on the MLP experiments in Figure LABEL:fig_mlp_gap
 We train a 4-layer MLP on MNIST.
@@
 We scale $c$ from 1 to 15.
-We record the average JS distance between the initial Hessian spectra of each pair of layers and find that the average JS distance increases as $c$ increases.
-As such, we define $c$ as the degree of block heterogeneity.
 For each $c$, we train SGD and Adam with default hyperparameters by grid-searching the learning rate from 1e-4 to 1e-1 and report the best test accuracy after 1 epoch.
-$c$ 1 2 3 4 5 6 7 8 9-15
-Blockwise distance 1e-3 1.61 33.99 65.32 260.77 298.39 351.69 353.48 $\operatorname{NaN}$
-Table 3: Consider a 4-layer MLP on MNIST.
-$c$ denotes the scaler that is multiplied to the output of each layer.
-Blockwise distance denotes the average JS distance between the initial Hessian spectra of each pair of layers.
-We find that the average JS distance increases as $c$ increases.
-For $c\geq 9$, the calculation of JS distance encounters numerical instability due to the division of near-zero values and returns $\operatorname{NaN}$.
-Appendix D More Experimental Results
-D.1 Performance comparison of AdamW and SGD on Different Architectures
-Here, we show the performance comparison of AdamW and SGD on different models.
 Appendix E Proofs
@@
 E.2 Proof of Theorem 1
+We first show that $C_{l,2}$ and $C_{l,1}$ are non-zero w.p.1.
+under the random initialization in Assumption 1.
+We define set $S_{i}=\{w;h_{i}^{T}w=0\}$ where $h_{i}\in\mathbb{R}^{d}$ is the $i$-th row of $H$.
+Since $H$ is positive definite, there is at least one non-zero entry in $h_{i}$, $i\in[d]$.
+As such, $S_{i}$ is a $(d-1)$-dimensional subspace of $\mathbb{R}^{d}$ and thus has zero Lebesgue measure in $\mathbb{R}^{d}$.
+Since $w^{0}$ follows continuous distribution, we have $\Pr\left(\{w^{0};h_{i}^{T}w^{0}=0\}\right)=0$, for $i=[d]$.
+Then we have
+$\Pr\left(\nabla\mathcal{L}(w^{0})\text{ has at least one zero entry}\right)$ $=$ $\Pr\left(Hw^{0}\text{ has at least one zero entry}\right)$ (11)
+$=$ $\Pr\left(\cup_{i=1}^{d}\{w^{0};h_{i}^{T}w^{0}=0\}\right)$ (12)
+$\leq$ $\sum_{i=1}^{d}\Pr\left(\{w^{0};h_{i}^{T}w^{0}=0\}\right)$ (13)
+$=$ $0$ (14)
+Therefore, $\nabla\mathcal{L}(w^{0})$ is elementwise non-zero w.p.1.., so $C_{l,1}$ and $C_{l,2}$ are non-zero for all $l\in[L]$, w.p.1..
+In the following analysis, We will assume $C_{l,1}$ and $C_{l,2}$ are non-zero.
 Without loss of generality, we assume $h=0$.
 This is because minimizing $\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ is equivalent to minimizing $\mathcal{L}(w)=\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ where $w^{*}=H^{-1}h$.
-By a linear transformation $z=w-w^{*}$, BGD for minimizing $\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ starting from $w^{0}$ is equivalent to BGD for minimizing $\frac{1}{2}z^{T}Hz$ starting from $z^{0}=w^{0}-w^{*}$.
+By a linear transformation $z=w-w^{*}$, Adam for minimizing $\frac{1}{2}\left(w-w^{*}\right)^{T}H\left(w-w^{*}\right)$ starting from $w^{0}$ is equivalent to Adam for minimizing $\frac{1}{2}z^{T}Hz$ starting from $z^{0}=w^{0}-w^{*}$.
 Thus we can assume $w^{*}=0$, or equivalently, $h=0$.
-When $h=0$, the update form of BGD becomes
-$w^{t+1}=w^{t}-\eta D_{BGD}^{-1}Hw^{t},$
-where $D_{BGD}=\operatorname{diag}\left(\left(\underbrace{\eta_{1},\cdots,\eta_{1}}_{d_{1}\text{times}},\cdots,\underbrace{\eta_{L},\cdots,\eta_{L}}_{d_{L}\text{times}}\right)^{T}\right)\in\mathbb{R}^{d\times d}.$
-In the following, we will first prove the rate of BGD for each block and then combine them together to get the total rate.
-For the $l$-th block, we choose $\eta_{l}=\frac{1}{\lambda_{l,1}}$.
-The update rule of BGD for the $l$-th block becomes:
-$w_{l}^{t+1}=w_{l}^{t}-\frac{1}{\lambda_{l,1}}H_{l}w_{l}^{t}$
-Now we apply the standard convergence proof of GD to the $l$-th block. By descent lemma, we have
-$[\mathcal{L}(w^{t+1})]_{l}$$\leq$ $[\mathcal{L}(w^{t})]_{l}+\langle w_{l}^{t+1}-w_{l}^{t},[\nabla\mathcal{L}(w^{t})]_{l}\rangle+\frac{\lambda_{l,1}}{2}\|w_{l}^{t+1}-w_{l}^{t}\|_{2}^{2}$ (11)
-$\leq$ $[\mathcal{L}(w^{t})]_{l}+\langle w_{l}^{t+1}-w_{l}^{t},H_{l}w_{l}^{t}\rangle+\frac{\lambda_{l,1}}{2}\eta^{2}w_{l}^{t}H_{l}^{T}H_{l}w_{l}^{t}$ (12)
-$=$ $[\mathcal{L}(w^{t})]_{l}-\frac{1}{2\lambda_{l,1}}w_{l}^{t}H_{l}^{T}H_{l}w_{l}^{t}$ (13)
-$\leq$ $[\mathcal{L}(w^{t})]_{l}-\frac{\lambda_{l,d_{l}}}{2\lambda_{l,1}}w_{l}^{t}H_{l}w_{l}^{t}$ (14)
-$=$ $\left(1-\frac{1}{\kappa_{l}}\right)[\mathcal{L}(w^{t})]_{l}$ (15)
-Substract both sides by $\mathcal{L}^{*}$ and we get $[\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\leq\left(1-\frac{1}{\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$.
-Summing up both sides over $l\in[L]$,we have
-$\mathcal{L}(w^{t+1})-\mathcal{L}^{*}$ $=$ $\sum_{l=1}^{L}[\left(\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
-$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
-$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
-$=$ $\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$
-This concludes the proof of Theorem 1.
-E.3 Proof of Theorem 2
-Similarly as the proof of Theorem 1, we assume $h=0$. The update rule of Adam becomes
+The update rule of Adam becomes
 $w^{t+1}=w^{t}-\eta(D_{Adam}^{0})^{-1}Hw^{t},$
@@
 These relations also hold for each block by changing the notation to $H_{l}$ $w_{l}^{t}$, $D_{Adam}^{0}$, and $d_{l}^{t},$ etc..
-Following the framework in [Sun and Ye, 2021], we try to bound the error yet to be optimized (a.k.a., cost-to-go) and the per-step improvement, respectively.
+Following the framework in [81], we try to bound the error yet to be optimized (a.k.a., cost-to-go) and the per-step improvement, respectively.
 The ratio of these two terms characterizes the rate of convergence.
@@
 $\overset{(*)}{\leq}$ $\frac{C_{l,2}^{2}\lambda_{l,1}^{2}}{(\min_{l\in[L]}C_{l,1}^{2})\lambda_{l,1}\lambda_{l,d_{l}}}$ (18)
-$\leq$ $\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l},$ (20)
+$\leq$ $\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\kappa_{l},$ (19)
 […]
 $\mathcal{L}(w^{t+1})-\mathcal{L}^{*}$ $=$ $\sum_{l=1}^{L}\left([\mathcal{L}(w^{t+1})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
-$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
-$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
-$=$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\max_{l\in[L]}\frac{C_{l,2}^{2}}{C_{l,1}^{2}}\right)\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$
+$\leq$ $\sum_{l=1}^{L}\left(1-\frac{1}{\left(\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\right)\kappa_{l}}\right)\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
+$\leq$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\right)\kappa_{l}}\right)\sum_{l=1}^{L}\left([\mathcal{L}(w^{t})]_{l}-[\mathcal{L}^{*}]_{l}\right)$
+$=$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\right)\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$

### v3

@@
 SGD performs worse than Adam by a significant margin on Transformers, but the reason remains unclear.
-In this work, we provide an explanation of SGD’s bad performance on Transformers through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs badly on problems with block heterogeneity.
-To validate that heterogeneity hampers SGD, we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD works well on problems without block heterogeneity but performs badly when the heterogeneity exists.
-Our initial theoretical analysis indicates that SGD performs poorly because it applies one single learning rate to all blocks, which cannot handle the heterogeneity among blocks.
+In this work, we provide an explanation through the lens of Hessian: (i) Transformers are “heterogeneous”: the Hessian spectrum across parameter blocks vary dramatically, a phenomenon we call “block heterogeneity"; (ii) Heterogeneity hampers SGD: SGD performs worse than Adam on problems with block heterogeneity.
+To validate (i) and (ii), we check various Transformers, CNNs, MLPs, and quadratic problems, and find that SGD can perform on par with Adam on problems without block heterogeneity, but performs worse than Adam when the heterogeneity exists.
+Our initial theoretical analysis indicates that SGD performs worse because it applies one single learning rate to all blocks, which cannot handle the heterogeneity among blocks.
 This limitation could be ameliorated if we use coordinate-wise learning rates, as designed in Adam.
@@
 1 Introduction
-Transformers [84] have become a major workhorse behind much remarkable progress in AI development (e.g., [1]).
+Transformers [84] have become a major workhorse behind AI development (e.g., [1]).
 However, the understanding of Transformer training remains limited.
 For instance, Transformer training largely relies on the Adam optimizer [45, 56].
-In contrast, stochastic gradient descent with momentum (SGD) [^2] We introduce the update rules of Adam(W) and SGD in Appendix C.1., the de-facto optimizer for convolution neural networks (CNNs) [48], performs poorly on Transformers (e.g., Figure LABEL:fig_blockwise_spectrum).
-Yet, the reasons behind this poor performance remain unclear.
-Understanding why SGD performs badly on Transformers is an intriguing question.
+In contrast, stochastic gradient descent with momentum (SGD) [^2] We introduce the update rules of Adam(W) and SGD in Appendix C.1., the de-facto optimizer for convolution neural networks (CNNs) [48], performs significantly worse than Adam on Transformers (e.g., Figure LABEL:fig_blockwise_spectrum).
+Yet, the reasons behind this performance gap remain unclear.
+Understanding why SGD performs worse than Adam on Transformers is an intriguing question.
 First, from a theoretical perspective, this can help us better understand the training of Transformers and more generally, neural networks.
 Second, from a computational perspective, the understanding may inspire the design of better algorithms for training neural networks.
-In this work, we explore the cause for SGD’s deficiency through the lens of Hessian.
+In this work, we explore why SGD largely underperforms Adam on Transformers through the lens of Hessian.
 We start by investigating the full Hessian spectrum of Transformers, i.e., the full eigenvalue density of Hessian (see Figure LABEL:fig_full_spectrum).
-By theory, the full Hessian spectrum largely determines the behavior of gradient-based methods [63, 33, 80, 37], so we suspect it may also help explain SGD’s bad performance on Transformers.
+By theory, the full Hessian spectrum largely determines the behavior of gradient-based methods [63, 33, 80, 37], so we suspect it may also help explain SGD’s unsatisfactory performance.
 Using tools from numerical linear algebra [7], we empirically compare the full spectra of CNNs (where SGD is on par with Adam) and those of Transformers (where SGD largely lags behind Adam).
 Unfortunately, as shown in Figure LABEL:fig_full_spectrum, the spectra for CNNs and Transformers are often largely similar despite the different optimizer behaviors.
-As such, we have not identified critical features in the full Hessian spectra associated with SGD’s bad performance on Transformers.
+As such, we have not identified critical features in the full Hessian spectra associated with the gap between Adam and SGD on Transformers.
 To reveal the cause, a more fine-grained investigation into the Hessian is needed.
-What would cause SGD to perform badly on Transformers but not on CNNs?
+What would cause SGD to perform significantly worse than Adam on Transformers, but not on CNNs?
 By dissecting the structures of CNNs and Transformers, we notice that CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers), while Transformers involve the non-sequential stacking of disparate parameter blocks (e.g.
@@
 This observation motivates us to investigate the Hessian spectrum of each parameter block, which we refer to as the blockwise Hessian spectrum.
-By inspecting the blockwise Hessian spectrum, we discover a possible explanation for SGD’s bad performance: the “heterogeneity” inherent in Transformers.
+By inspecting the blockwise Hessian spectrum, we discover a possible explanation for why SGD is worse: the “heterogeneity” inherent in Transformers.
 We provide both empirical and theoretical evidence to support this explanation.
 Our contributions can be summarized as follows:
-  • Failure mode of SGD on Transformers.
-We explain why SGD underperforms Adam on Transformers by examining the blockwise Hessian spectrum.
-First, we identify a phenomenon called ’block heterogeneity’, which refers to the large differences in the Hessian spectra across parameter blocks.
+  • Why SGD underperforms Adam on Transformers.
+We explain why SGD is worse than Adam on Transformers by examining the blockwise Hessian spectrum.
+First, we identify a phenomenon called “block heterogeneity", which refers to the large differences in the Hessian spectra across parameter blocks.
 This block heterogeneity is observed in all examined Transformers but not in CNNs.
 Second, we verify that block heterogeneity hinders SGD.
-Across various Transformers, CNNs, and MLPs, we show that SGD consistently performs poorly on problems with block heterogeneity but works well otherwise.
+Across various Transformers, CNNs, and MLPs, we show that SGD consistently performs worse than Adam on problems with block heterogeneity but can perform similarly to Adam, otherwise.
   • Theoretical results on quadratic models.
-We construct convex quadratic problems with and without block heterogeneity and find that gradient descent (GD) performs significantly worse than Adam on problems with block heterogeneity but comparably otherwise.
-Our theoretical analysis shows that GD struggles on quadratic problems with block heterogeneity, suggesting that using a single learning rate for all blocks is a key reason for GD’s bad performance.
+We construct convex quadratic problems with and without block heterogeneity and find that gradient descent (GD) largely underperforms Adam on problems with block heterogeneity, but can perform comparably otherwise.
+Our theoretical analysis shows that GD can be slower than Adam on quadratic problems with block heterogeneity.
+We point out GD is slower than Adam because it uses a single learning rate for all blocks.
 The deficiency can be mitigated by assigning different learning rates across blocks, as Adam does.
-We emphasize that we do not claim block heterogeneity is the only cause for SGD’s bad performance, but just that it is at least one important cause.
-We verify, both empirically and theoretically, that SGD performs poorly when block heterogeneity is present.
+We emphasize that we do not claim block heterogeneity is the only cause for the performance gap between Adam and SGD, but just that it is at least one important cause.
+We verify, both empirically and theoretically, that SGD underperforms Adam when block heterogeneity is present.
 2 Main Results
@@
 Second, previous research shows that the Hessian spectrum provides insights into neural network phenomena, like BatchNorm’s effect on training speed [32].
-Therefore, we hypothesize that the Hessian spectrum may also explain SGD’s bad performance on Transformers.
-We compare the full Hessian spectra of CNNs (where SGD works well) and those of Transformers (where SGD struggles), as shown in Figure LABEL:fig_full_spectrum.
-Unfortunately, the results suggest that the full Hessian spectrum alone may not suffice to explain why SGD struggles on Transformers.
+Therefore, we hypothesize that the Hessian spectrum may also explain why SGD largely lags behind Adam on Transformers.
+We compare the full Hessian spectra of CNNs (where SGD performs similarly to Adam) and those of Transformers (where SGD underperforms Adam), as shown in Figure LABEL:fig_full_spectrum.
+Unfortunately, the results suggest that the full Hessian spectrum alone may not suffice to explain the gap between Adam and SGD on Transformers.
 We elaborate as follows.
@@
 Regarding (A), we observe that the eigenvalues are dispersed similarly across different models, with no notably large outlier for Transformers.
-Thus, dispersion does not seem to be related to the bad performance of SGD.
+Thus, dispersion does not seem to be related to why SGD is worse than Adam.
 We further investigate (B) and (C).
@@
 As training proceeds, the majority of negative eigenvalues disappear, and the shape evolves into a combination of a “bulk” and some “outliers”.
-Since the spectral shape and evolution are quite similar for both Transformers and CNNs, they cannot explain the incompetence of SGD on Transformers either.
-In summary, we have not identified any critical phenomena in the full Hessian spectra that can be linked to SGD’s bad performance on Transformers.
+Since the spectral shape and evolution are quite similar for both Transformers and CNNs, they cannot explain why SGD is worse than Adam on Transformers, either.
+In summary, we have not identified any critical phenomena in the full Hessian spectra that can be linked to the performance gap between Adam and SGD on Transformers.
 2.3 Main Findings Through Blockwise Hessian Spectra
-What other factors could cause SGD to struggle on Transformers but not on CNNs? We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
+What other factors could cause SGD to perform significantly worse than Adam on Transformers but not on CNNs?
+We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
   • 1.
@@
 Intuitively, the block-homogeneity in CNNs comes from the repetitively similar convolution layers, while the block-heterogeneity in Transformers stems from the non-sequential stacking disparate layers such as Query, Value, and MLPs.
-In the following, we will show that the block heterogeneity is strongly correlated with SGD’s bad performance on Transformers.
-2.4 SGD Performs Badly on Various Tasks with Block Heterogeneity
-Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD performs badly on Transformers.
+In the following, we will show that the block heterogeneity is strongly correlated with the performance gap between SGD and Adam on Transformers.
+2.4 SGD Performs Worse than Adam on Various Tasks with Block Heterogeneity
+Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD is worse than Adam on Transformers.
 (2) Transformers have block-heterogeneity.
-Now we further link block heterogeneity to SGD’s bad performance on non-Transformer models.
-This would directly establish a connection between block heterogeneity and SGD’s bad performance, without going through Transformers or attention blocks as an intermediary.
+Now we further link block heterogeneity to SGD’s unsatisfactory performance on non-Transformer models.
+This would directly establish a connection between “block heterogeneity" and “why SGD is worse than Adam", without going through Transformers or attention blocks as an intermediary.
 We consider one man-made example and one real-world example.
@@
 We consider a 4-layer MLP on MNIST and change the degree of heterogeneity by scaling each layer by constant $c$.
-Figure LABEL:fig_mlp_gap (a) shows SGD fails as heterogeneity grows.
+Figure LABEL:fig_mlp_gap (a) shows SGD gradually performs worse than Adam as heterogeneity grows.
 Example 2: MLP-mixer.
@@
 2.6 Implication on Choosing SGD or Adam
-The above findings show that SGD can perform badly on various architectures. This leads to an intriguing question: Can we predict the bad performance of SGD before training begins?
+We have shown that SGD can largely underperform Adam on various architectures. This leads to an intriguing question: Can we predict the incompetence of SGD before the training begins?
 Our findings can bring up an empirical guidance: we can compute the blockwise spectrum of initial Hessian, and then decide whether to use Adam or SGD.
@@
 4 Conclusion
-In this work, we explore why SGD performs badly on Transformers.
-By numerically exploring various neural nets and quadratic problems, we establish a phenomenon called block heterogeneity in Hessian and link it to the deficiency of SGD.
+In this work, we explore why SGD largely underperforms Adam on Transformers.
+we establish a phenomenon called block heterogeneity in Hessian and link it to the performance gap between Adam and SGD.
 Initial theory is provided to support the claim.
 Appendix A Related Works
-On the deficiency of SGD on Transformers
-There is an active line of works that explores why SGD performs badly on Transformers.
+On the unsatisfactory performance of SGD on Transformers
+There is an active line of works that explores why SGD performs significantly worse than Adam on Transformers.
 One representative hypothesis is that SGD cannot handle the heavy-tailed stochastic noise in language tasks [106].
@@
 [46] reported that the gap between Adam and SGD maintains even in the full-batch case with no stochasticity, so there might be other reasons.
-Further, SGD performs poorly on Vision Transformers on ImageNet (See Figure LABEL:fig:nlp_figure.
+Further, SGD performs worse than Adam on Vision Transformers on ImageNet (See Figure LABEL:fig:nlp_figure.
 Also see [93] for more evidence), so the data modality (e.g., language or vision tasks) might not be as crucial as the architecture.
@@
 So in Figure LABEL:fig:nlp_figure, SGD in NLP tasks essentially refers to clipped SGD.
-.
-So there might be other reasons that hamper SGD.
-Different from these works, we find SGD performs badly because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
+, so there might be other reasons.
+Different from these works, we find SGD underperforms Adam because it uses one single learning rate for all blocks, which cannot handle the Hessian heterogeneity among blocks.
 Understanding of Adam.
@@
 [50] show that Adam and its variant outperform SGD under relaxed smoothness conditions, based on the intuition that Adam can adaptively change its learning rate along iteration (over time).
-We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes performs badly while Adam works well (Figure LABEL:fig_quadratic).
+We pointed out that the theory is not complete: even for quadratic functions where the smoothness is fixed, SGD sometimes performs largely worse than Adam (Figure LABEL:fig_quadratic).
 This indicates that the benefit of Adam is not merely due to its ability to adaptively change the learning rate (over time), and there are other reasons for Adam’s success.
@@
 Different from all aforementioned works, we investigate the training difficulties of Transformers through the eigenvalues of Hessian.
-We establish a strong correlation between the blockwise Hessian spectra of Transformers and the poor performance of SGD.
+We establish a strong correlation between “the blockwise Hessian spectra of Transformers" and “why SGD largely underperforms Adam on Transformers".
 We realize that our attempt is just a first step towards understanding Transformer training, and we believe there is rich information hidden in Hessian and we leave more fine-grained analysis as future works.
@@
 Language models are trained on different English corpus.
+We grid-search the learning rates for SGD and Adam under the same budget and report the best result for each optimizer.
 See Appendix D.1 for more implementation details.

### v4

@@
 We verify, both empirically and theoretically, that SGD underperforms Adam when block heterogeneity is present.
-2 Main Results
+2 Problem Settings and Initial Attempts
 2.1 Problem Settings
@@
 A detailed introduction to SLQ is provided in Appendix C.2.
-All experimental setups in this section is shown in Appendix D.
+All experimental setups are shown in Appendix D.
 We focus primarily on the following models/tasks.
@@
 In summary, we have not identified any critical phenomena in the full Hessian spectra that can be linked to the performance gap between Adam and SGD on Transformers.
-2.3 Main Findings Through Blockwise Hessian Spectra
+2.3 Motivations of Investigating Blockwise Hessian Spectra
 What other factors could cause SGD to perform significantly worse than Adam on Transformers but not on CNNs?
-We identify some critical features that have been overlooked in the full Hessian spectrum analysis above.
-  • 1.
-    Hessian structure.
-Existing literatures showed that the Hessians of MLPs are close to block-diagonal matrices [18, 73, 60].
-We restate their findings in Appendix B.
-Further, [18, Section 7] theoretically attributes the block diagonal structure to (i) the layer-by-layer design in NNs and (ii) the Cross-Entropy loss.
-Following this line of work, we also observe a near block-diagonal Hessian in small Transformers in Figure LABEL:fig_block_diagnal, where the variables in each principal block correspond to the parameters of each block in the Transformer.
-These results suggest that near block-diagonal Hessian might be common in NNs.
-  • 2.
-    Build-up rules of Transformers.
+We identify one critical feature that has been overlooked in the full Hessian spectrum analysis above: the building-up rules of Transformers.
 As shown in Figure LABEL:fig_blockwise_spectrum, CNNs are constructed by the repetitive stacking of similar parameter blocks (convolution layers).
@@
 Further, these blocks are stacked in a non-sequential manner.
-Due to the different designs of these blocks, they may have different properties for optimization, which can further affect the optimizer behavior.
-Combining these together, we hypothesize that the blockwise Hessian spectrum, i.e., the spectrum of principal blocks of Hessian $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$, might provide additional insights.
-What extra information can be contained in the blockwise spectrum but not in the full spectrum?
-By definition, the blockwise Hessians form the principal block sub-matrix of the full Hessian.
-We note that Transformers are observed to have a near block-diagonal Hessian.
-For block diagonal matrices, blockwise spectra encode the location of eigenvalues, i.e., which block an eigenvalue (of the full matrix) resides in.
-In contrast, the full spectrum overlooks this location information.
-In the following, we study the blockwise Hessian spectra of various models and show that they indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
-We here demonstrate the shape of blockwise spectra in VGG16 [40] (CNN) and BERT [25] (Transformer).
+We hypothesize that the “different designs among parameter blocks" can be reflected in the Hessian of these parameter blocks, which might affect algorithmic behavior.
+This inspires us to investigate the blockwise Hessian spectra, i.e., the spectrum of principal blocks of Hessian $[\nabla^{2}\mathcal{L}(w)]_{l},l\in[L]$.
+In parallel to the motivation above, we further provide another evidence that blockwise spectra might be helpful.
+Classical literature showed that the Hessians of neural nets are near-block-diagonal matrices [18], i.e., the magnitudes in the Hessian principle blocks are much larger than those in the off-diagonal blocks.
+We restate their findings in Figure LABEL:fig_block_diagnal (a).
+This implies that the majority of Hessian information indeed lies in its principle blocks, and the blockwise Hessian of neural nets might contain valuable information.
+To summarize, the “heterogeneous" building-up rules of Transformers inspire us to check the blockwise Hessian, i.e., the principle blocks of the Hessian.
+The classical results of neural nets [18] further support us to explore this direction since they find that the majority of Hessian information indeed lies in its principle blocks.
+In the following, we study the blockwise Hessian spectra of various neural networks.
+For ease of implementation, we define parameter blocks under the PyTorch partition.
+We show that the blockwise spectra indeed carry more information than the full spectrum for distinguishing CNNs and Transformers.
+Remark: why near-block-diagonal?
+We briefly restate the analysis in [18, Section 7] to explain the near-block-diagonal Hessian structure of neural nets.
+Consider minimizing $\ell(f(\theta,x),y)$ where $\ell(\cdot,\cdot)$ is the Cross-Entropy (CE) loss, $f(\theta,x)=\sum_{i=1}^{n}v_{i}\phi(w_{i}^{\top}x)$ is an 1-hidden-layer neural network with input $x\in\mathbb{R}^{d}$, weight $w_{i}\in\mathbb{R}^{d}$, $v_{i}\in\mathbb{R}$, and label $y\in\{0,1\}$, then the off-diagonal-block Hessian elements will contain
+$\frac{\partial^{2}\ell(f(\theta,x),y)}{\partial w_{i}\partial w_{j}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{\theta}(y|x)\left(1-p_{\theta}(y|x)\right)}v_{i}v_{j}\phi^{\prime}\left(w_{i}^{\top}x\right)\phi^{\prime}\left(w_{j}^{\top}x\right)xx^{\top}\quad\text{for }i\neq j,$ (1)
+where $p_{\theta}(y|x)=1/(1+\exp(-yf(\theta,x)))$ and $\phi^{\prime}(\cdot)$ is the derivative of $\phi(\cdot)$.
+Note that the term $p_{\theta}(y|x)\left(1-p_{\theta}(y|x)\right)$ will vanish rapidly since the training objective is to maximize $p_{\theta}(y|x)$.
+Consequently, this drives the Hessian towards a near-block-diagonal configuration, with each block representing an output neuron.
+This result is validated in Figure LABEL:fig_block_diagnal: we find that the near-block-diagonal structure appears at 1% step and it maintains along the training.
+3 Main Results
+3.1 Transformers Exhibit Block Heterogeneity in Hessian, while CNNs Do Not
+We now compare the shape of blockwise spectra in VGG16 [40] (CNN) and BERT [25] (Transformer).
 We sample four blocks for each model and present the spectra in Figure LABEL:fig_blockwise_spectrum.
-In BERT, the Hessian spectra of embedding, attention, and MLP blocks are largely different.
+In BERT, the spectra of embedding, attention, and MLP blocks are largely different.
 In contrast, in ResNet, the spectra of convolution layers are similar.
@@
 We calculate the Jensen-Shannon (JS) distance between two eigenvalue densities of all possible block pairs and show the results in Figure LABEL:fig_blockwise_heatmap.
-The results in Figure LABEL:fig_blockwise_heatmap showed a new phenomenon: for all Transformers we checked, the blockwise Hessian spectra are largely different across blocks.
-In the following, we refer to this phenomenon as “block heterogeneity".
-In contrast, the blockwise Hessian spectra of CNNs are similar and the block heterogeneity is not observed.
-We refer to this phenomenon as “block homogeneity".
-These results indicate that block heterogeneity is informative in distinguishing CNNs and Transformers.
-Intuitively, the block-homogeneity in CNNs comes from the repetitively similar convolution layers, while the block-heterogeneity in Transformers stems from the non-sequential stacking disparate layers such as Query, Value, and MLPs.
+We summarize our findings in Observation 1.
+Observation 1: For all Transformers we checked, the blockwise Hessian spectra are largely different from each other. In contrast, the blockwise Hessian spectra of CNNs are similar.
+In the following, we refer to the phenomenon of Transformers as “block heterogeneity", and refer to that of CNN as “block homogeneity".
+The observations in Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap indicate that block heterogeneity is informative in distinguishing CNNs and Transformers.
 In the following, we will show that the block heterogeneity is strongly correlated with the performance gap between SGD and Adam on Transformers.
@@
 Figure LABEL:fig_blockwise_spectrum and LABEL:fig_blockwise_heatmap have shown that (1) SGD is worse than Adam on Transformers.
-(2) Transformers have block-heterogeneity.
+(2) Transformers have block heterogeneity.
 Now we further link block heterogeneity to SGD’s unsatisfactory performance on non-Transformer models.
@@
 We consider MLP-mixer [81], a famous all-MLP architecture that outperforms CNNs and ViTs on some vision tasks.
-Figure LABEL:fig_mlp_gap (b) (c) shows that the initial Hessian of MLP-mixer has block heterogeneity and SGD lags behind Adam on this architecture.
+Figure LABEL:fig_mlp_gap (b) (c) show that the initial Hessian of MLP-mixer has block heterogeneity and SGD lags behind Adam on this architecture.
+We summarize the findings so far in Observation 2.
+Observation 2: For all tasks that we checked, SGD is worse than Adam when block heterogeneity exists, regardless of whether Transformers or attention mechanisms are utilized.
 3.3 Reduced Block Heterogeneity in Pre-trained Transformers
@@
 This partly explains why simpler methods like SGD and even its zeroth-order version can still be effective for fine-tuning language models, albeit with slower convergence [59, 60].
+In Figure LABEL:fig:hessian_along_training_vit in Appendix B, we further report the evolution of the block heterogeneity of ViT-base along the training.
+Similarly to GPT2 in Figure LABEL:fig_sft_gap, we find that the block heterogeneity of ViT-base tends to reduce after the training.
+In addition, we find that SGD can perform better when initializing at the weight with less heterogeneity, e.g., initializing at 50% total training steps.
+We hypothesize that “the attenuation of Hessian heterogeneity" is a common phenomenon after training, and we leave detailed investigation as a future direction.
+Observation 3: Block heterogeneity in Hessian tends to reduce after (pre)-training.
 3.4 Implication on Choosing SGD or Adam
@@
 Further, $JS^{0}$ is independent of optimizers and could be checked before training.
+To validate the effectiveness of the quantitative metric $JS^{0}$, we summarize $JS^{0}$ of different models and the corresponding SGD performance in Figure 4.
+We find that the performance gap between SGD and Adam becomes greater as $JS^{0}$ increases.
+Thus, $JS^{0}$ can serve as a potential indicator to predict whether SGD may underperform Adam.
 Table 1: $JS^{0}$ denotes the average JS distance between the initial Hessian spectra of each pair of parameter blocks. A larger $JS^{0}$ suggests that the task is more difficult for SGD.
@@
 $JS^{0}$ 0.10 0.09 18.84 34.90 53.38 83.23 286.41
-To validate the effectiveness of the quantitative metric $JS^{0}$, we summarize $JS^{0}$ of different models and the corresponding SGD performance in Figure 4.
-We find that the performance gap between SGD and Adam becomes greater as $JS^{0}$ increases.
-Thus, $JS^{0}$ can serve as a potential indicator to predict whether SGD may underperform Adam.
 Figure 4: Comparison of $JS^{0}$ and the performance of SGD on different models. We find the performance gap between SGD and Adam becomes greater as $JS^{0}$ increases.
+Finally, we remark $JS^{0}$ is rather expensive to compute due to the overhead of SQL: it requires comparable time to one training run.
+Fortunately, we find the original SQL is rather redundant for measuring hessian heterogeneity.
+We propose some simple tricks to significantly reduce the computation time, while still effectively detecting the Hessian heterogeneity.
+We call it simplified SQL and we present it in Table 3 in Appendix B.
+As a result, the simplified SQL can obtain the same message as in Table 1 while only taking negligible time (e.g., $<0.001$s for ResNet18).
 4 Case Study of Quadratic Models and Preliminary Theory
@@
 We emphasize that all Hessians have the same condition number.
-Further, Hessian of Case 3 and 4 share all the eigenvalues (not just the extreme ones).
-The gap between Adam and GD is due to the different blockwise spectra caused by the different locations of eigenvalues.
+Further, Hessian in Case 3 and 4 share all the eigenvalues (not just the extreme ones).
+The gap between Adam and GD is purely due to the different blockwise spectra caused by the different locations of eigenvalues.
+Case 3 and 4 help reveal the causal relation between “block heterogeneity in Hessian" and “GD is worse than Adam".
 We hypothesize that GD performs badly because it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks.
@@
 We summarize the complexity of GD and Adam in Table 2.
-We remark that there is room to improve the complexity bound of Adam, but it could be challenging.
-We provide more technical discussions in Appendix B.
+Remark: some common misconceptions.
+During the review process, we find that readers might conclude that “Theorem 1 implies Adam under homogeneity has worse complexity than Adam under heterogeneity".
+We now clarify that this claim is not correct, and there is no conclusion on “whether Adam under homogeneity is faster or slower than Adam under heterogeneity".
+Similarly, Theorem 1 does not imply “Adam always converges similarly as GD under homogeneity".
+Though it is observed on CNNs, there is no general conclusion of this kind.
+For interested readers, we provide a detailed explanation in Appendix B.
 Table 2: The complexity of GD and Adam for minimizing a strongly convex quadratic function with block diagonal Hessian.
@@
 Complexity $\tilde{O}(\kappa)$ $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$
+How to obtain a tighter complexity bound of Adam?
+It is valid to ask whether the complexity upper bound in Theorem 1 can be tightened, e.g., improve the factor of $r$.
+We point out it would be difficult if there is no extra structure on $H_{l}$.
+A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$.
+Intuitively, a diagonal preconditioner of $H_{l}$ is powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant [30].
+Unfortunately, it is unclear whether these structures hold in Transformers.
+Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number.
+For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $7.09\kappa_{1}$, $18.98\kappa_{2}$, $18.76\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials).
+It would be interesting to explore if there are special structures of $H_{l}$ in Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it.
+We leave it as a future direction.
+More discussions on the theoretical advantage of Adam.
+Although Adam preconditioner might not always reduce the “local" condition number $\kappa_{l}$, the coefficient in the complexity is now independent of the “global" condition number $\kappa$.
+As argued above, such changes in coefficient could lead to considerable improvement over GD.
+Such improvement in complexity is attributed to the block diagonal structure in Hessian as well as its heterogeneous blockwise spectrum.
+To our knowledge, such improvement is not shown in the existing literature.
+One possible reason is that: for the optimization community, it is very rare to analyze (near-) block-diagonal Hessian structure since typical problems do not have such structure.
+For instance, in the classical non-linear programming dataset [48], all problems have non-block-diagonal Hessian.
+We suggest a different perspective to characterize modern optimization problems.
+We believe our perspective is new because it is built upon multiple non-trivial findings.
+In summary, our theory indicates that: for problems with block heterogeneity, the single-learning rate methods like GD can largely lag behind coordinate-wise learning rate methods like Adam.
 5 Conclusion
@@
 we establish a phenomenon called block heterogeneity in Hessian and link it to the performance gap between Adam and SGD.
-Initial theory is provided to support the claim.
+We numerically verify our claim on various Transformers, CNNs, MLPs, and quadratic problems.
+Initial theory is also provided to support the claim.
+Acknowledgements
+Yushun Zhang would like to thank Yinyu Ye, Wentao Ding, Guiyu Hong, Yingru Li, and Bohan Wang for the valuable discussions.
+The work of Ruoyu Sun is supported by NSFC (No.
+12326608); Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone Project (No.HZQSWS-KCCYB-2024016); University Development Fund UDF01001491, the Chinese University of Hong Kong, Shenzhen; Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001).
+The work of Z.-Q.
+Luo was supported by the Guangdong Major Project of Basic and Applied Basic Research (No.2023B0303000001), the Guangdong Provincial Key Laboratory of Big Data Computing, and the National Key Research and Development Project under grant 2022YFA1003900.
+Broader Impacts
+We explore why SGD performs worse than Adam for training Transformers.
+Our work can help the community better understand large AI model training.
+However, it would be a potential threat if the AI models are used for illegal usage.
 Appendix A Related Works
@@
 See Appendix D.1 for more implementation details.
-Block-diagonal structure in the existing literature.
-We remark that Collobert [18], Roux et al.
-[73], Martens and Grosse [60] also observed the block-diagonal structure in (approximated) Hessian of small-scaled MLPs.
-We restate their findings in Figure LABEL:fig_block_diagnal_existing.
-[18, Section 7] further theoretically proved that the block diagonal structure stems from (i) the layer-by-layer structure and (ii) the Cross-Entropy loss.
-These results suggest that block-diagonal structure might be common in NNs.
+More results for SGD under careful tuning.
+For ViT-base, GPT2-nano, and BERT, we further present the performance of SGD under learning rate grid search.
+For ViT-base training on ImageNet, we report the results after 30 epochs (or equivalently, about 30k iterations).
+We cannot afford further training ViT-base due to the limited hardware resources (a complete run of 90 epochs would take $>2$ weeks for each curve).
+As shown in Figure LABEL:fig:sgd_grid_search, SGD consistently performs worse than Adam on ViT-base, GPT2, and BERT.
+The best results for SGD are picked out and presented in Figure LABEL:fig:nlp_figure.
+Note that we are not the first ones to report that “SGD performs worse than Adam on ViT".
+An influential work [92] also reports that SGD is worse than Adam on vanilla ViT.
+The authors report that “SGD yields significantly worse results than AdamW (on ViT)", and “ViT often fails to converge with SGD " (their Figure 3).
+These results align with our findings in Figure LABEL:fig:sgd_grid_search.
+Detailed training curves of Figure LABEL:fig_mlp_gap.
+In Figure LABEL:fig_mlp_gap_appendix, we present the detailed training curves for the man-made MLP in Figure LABEL:fig_mlp_gap.
+We find that SGD performs worse as heterogeneity grows, while Adam still performs well.
+The evolution of Hessian heterogeneity along training.
+For ViT-base, we further investigate the evolution of block heterogeneity of Hessian along the training.
+As shown in Figure LABEL:fig:hessian_along_training_vit, we find that heterogeneity attenuates along the training.
+We further take the checkpoint of ViT-base at 50% training step and switch AdamW to SGD, we observe that now SGD performs better than training from scratch as in Figure LABEL:fig:nlp_figure (a).
+SGD performs better here because there is less heterogeneity when initializing in the middle of training.
+Simplifed SQL for calculating $JS^{0}$ in Section 3.4.
+We note that $JS^{0}$ in Table 1 is rather expensive to compute due to the computational overhead of SQL: it requires comparable time to one training run.
+Fortunately, we find the original SQL is redundant for measuring hessian heterogeneity.
+We propose the following simple tweaks to significantly reduce the computation time, while still effectively detecting the Hessian heterogeneity.
+We call it simplified SQL.
+  • Change the hyperparemters of SQL, including:
+    • We change $\texttt{num}_{v}=10$ to $\texttt{num}_{v}=1$.
+In SQL, $\texttt{num}_{v}$ decides the number of random Gaussian vectors to approximate the expected quadrature.
+It is reasonable to reduce $\texttt{num}_{v}$ because in high dimensional space, random vectors tend to concentrate around their mean, so one random sample can already be informative enough.
+    • We change the Lanzcos step $m=100$ to $m=10$.
+The reduction on Lanzcos step will have a coarse estimation on the middle eigenvalue, but won’t affect much the heterogeneity, which is more dependent on the extreme eigenvalues.
+  • Randomly sample a subset of blocks and reduce batch size for estimating the spectrum. We uniformly sample 50% blocks and choose batch size = 32 (previously batch size = 1024).
+We report the result and runtime in Table 3.
+As a result, the simplified SQL can obtain the same message as the original SQL: $JS^{0}$ of ResNet is about 100x smaller than that of BERT.
+Further, the simplified SQL is highly efficient to compute.
+With this simplified SQL, we believe our method can efficiently scale to larger models.
+The result is tested on a single V100.
+Table 3: $JS^{0}$ computed by simplified SQL.
+We find that the simplified SQL can obtain the same message as the original SQL: $JS^{0}$ of ResNet is about 100x smaller than that of BERT.
+Further, the simplified SQL is efficient to compute.
+Model JSO Time for JSO Time for Training Time for JSO / Time for training
+BERT 98.8344 20 s 4 h 0.0014
+ResNet18 0.3569 65 s 87.5 h 0.0002
 Blockwise spectra for quadratic models in Section 4.1.
 We here visualize the blockwise spectrum for the quadratic models in Case 1 and Case 2. These spectra are collected from GPT2 and ResNet18, respectively.
-How to obtain a tighter complexity bound of Adam?
-It is valid to ask whether the complexity upper bound in Theorem 1 $\kappa_{Adam,l}=r\kappa_{l}$ can be tightened, e.g., improve the factor of $r$.
-We point out it would be difficult if there is no extra structure on $H_{l}$.
-A key technical step is to bound the condition number of the preconditioned matrix $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$.
-Intuitively, a diagonal preconditioner of $H_{l}$ is powerful when $H_{l}$ itself has a near-diagonal structure, e.g., pure diagonal, tridiagonal or diagonal dominant [30].
-Unfortunately, it is unclear whether these structures hold in Transformers.
-Without any assumption on $H_{l}$, we find that the diagonal preconditioner of $D_{Adam}^{0}$ could increase the condition number.
-For instance, when using standard Gaussian initialization, in case 3, we find $\kappa\left((D_{Adam,l}^{0})^{-1}H_{l}\right)$ equals $7.09\kappa_{1}$, $18.98\kappa_{2}$, $18.76\kappa_{3}$ for the 3 blocks, respectively (all averaged over 1000 trials).
-It would be interesting to explore if there are special structures of $H_{l}$ in Transformers such that Adam preconditioner can reduce $\kappa_{l}$, rather than increase it.
-We leave it as a future direction.
-Although Adam preconditioner might not always reduce the “local" condition number $\kappa_{l}$, the coefficient in the complexity is now independent of the “global" condition number $\kappa$.
-As argued above, such changes in coefficient could lead to considerable improvement over GD.
-Such improvement in complexity is attributed to the block diagonal structure in Hessian as well as its heterogeneous blockwise spectrum.
-To our knowledge, such improvement is not shown in the existing literature.
-In summary, our theory indicates that: for problems with block heterogeneity, the single-learning rate methods like GD can largely lag behind coordinate-wise learning rate methods like Adam.
+Does our theory imply that Adam under homogeneity is slower than Adam under heterogeneity?
+ [^5] We would like to thank the anonymous Reviewer pTmy for raising this insightful question.
+During the review process, we find that readers might conclude that “Theorem 1 implies Adam under homogeneity has worse complexity than Adam under heterogeneity".
+We would like to clarify that this conclusion is not correct, and there is no conclusion on “whether Adam under homogeneity is faster or slower than Adam under heterogeneity ".
+We explain as follows.
+Our theoretical result states that Adam has complexity $\mathcal{O}\left(\max_{l}\kappa_{l}\right)$.
+If our result implies the above conclusion, one needs the following argument: when changing heterogeneity to homogeneity, $\max_{l}\kappa_{l}$ increases, and thus Adam is slower.
+However, “changing heterogeneity to homogeneity" does not necessarily mean “$\max_{l}\kappa_{l}$ increases".
+Actually, $\max_{l}\kappa_{l}$ can change in an arbitrary way (can increase, decrease, or keep the same) when changing the heterogeneity.
+We provide three examples below.
+We will use Adam (homo) to denote the convergence rate of Adam on homogeneous Hessian, similarly for Adam (hetero).
+Example 1: Adam (homo) is same as Adam (hetero).
+  • Case 1-1 (homogeneous): eigenvalues are {1,2}, {1,2}
+  • Case 1-2 (heterogeneous): eigenvalues are {1,2}, {11,12}
+Since $\max_{l}\kappa_{l}$ are the same for both Case 1-1 and 1-2, Adam (homo) is the same as Adam (hetero).
+Example 2: Adam (homo) is faster than Adam (hetero).
+  • Case 2-1 (homogeneous): eigenvalues are {1,1.5}, {1,1.5}
+  • Case 2-2 (heterogeneous): eigenvalues are {1,2}, {11,12}
+Since Case 2-1 has smaller $\max_{l}\kappa_{l}$ than Case 2-2, Adam (homo) is faster than Adam (hetero).
+Example 3: Adam (homo) is slower than Adam (hetero).
+  • Case 3-1 (homogeneous): eigenvalues are {1,12}, {1,12}
+  • Case 3-2 (heterogeneous): eigenvalues are {1,2}, {11,12}
+Since Case 3-1 has larger $\max_{l}\kappa_{l}$ than Case 3-2, Adam (homo) is slower than Adam (hetero).
+To sum up, there is no conclusion on “whether Adam under homogeneity is faster or slower than Adam under heterogeneity ". Either case can happen.
+One possible source of confusion may come from the numerical examples (Case 3 and 4 in Section Section 4.1).
+If comparing two figures, Adam (homo) in Case 3 is slower than Adam (hetero) in Case 4.
+But as argued above, this is just one example, and it does not show Adam (homo) is always slower than Adam (hetero).
 Appendix C More Preliminaries
@@
 [82] show that SLQ converges faster than any other polynomial expansion method for spectrum estimation (e.g., Chebyshev methods used in [2]).
-See [83, Theorem 4.1] for a formal statement.
+See [82, Theorem 4.1] for a formal statement.
 We remark that there are at least four versions of the Lanczos algorithm in Step 3.
@@
 We first pre-train GPT2 on OpenwebText for 25B tokens and then fine-tune it on a subset of Alpaca Eval  [^11] https://huggingface.co/datasets/tatsu-lab/alpaca_eval.
-D.2 Implementation Details on Figure LABEL:fig_block_diagnal
-We use the code base of NanoGPT to train decoder-only Transformers on 4 consecutive tokens randomly selected from Openwebtext.
-We set the model configuration as: context window = 2, number of heads = 2 and the embedding dimension = 4.
-In MLP layers, all widths equal to 4.
-We choose the number of layers to be 2, 4, and 8.
-We remove all the LayerNorms in the model.
-The rest of the model configurations are set to their default values in the code base.
-We compute the Hessian on all the parameters blocks in attention and MLP layers.
-In Figure LABEL:fig_block_diagnal (a), the variables each block corresponds to the parameters in Query and Key; Value and Projection; and MLP, respectively.
-For better visualization, We take the absolute value of each entry to distinguish non-zero values (including negatives) from those near 0 and then report the average value in each block.
-Similarly for the rest of the figures.
-Due to the intensive overhead of computing and storing the whole Hessian, we have yet to check the block-diagonal structure on larger models.
-Rigorously speaking, so far we have not gotten sufficient evidence to claim this structure commonly holds in larger Transformers.
-It requires new numerical methods to efficiently check the block-diagonal Hessian structure without explicitly calculating them.
-We leave it as an interesting future direction.
-D.3 More Details on the MLP experiments in Figure LABEL:fig_mlp_gap
+D.2 Ablation Studies of SQL on a Small Tranformer
+On a small GPT model with 75k parameters, we compare (1) the true Hessian spectrum and (2) the estimated Hessian spectrum curve by SQL method.
+As shown in Figure LABEL:fig:sql_quality, we find that the SQL method can produce accurate estimation.
+D.3 Implementation Details on Figure LABEL:fig_block_diagnal
+We employ a synthetic dataset designed for binary classification, with 100 data points generated through the process outlined below.
+Our model is a 1-hidden-layer neural network, featuring an input size of 64 and a layer width of 8, utilizing the hyperbolic tangent (Tanh) as the activation function.
+We train this model over 1000 iterations using the Adam optimizer with a learning rate set to $1\times 10^{-4}$, achieving the classification accuracy of 100%.
+1def generate_data(n_samples_per_class, n_classes, input_dim):
+2 # Generate synthetic data for specified dimensions
+3 X = []
+4 y = []
+5 for i in range(n_classes):
+6 center = np.random.rand(input_dim) * 10 # Random class center
+7 class_samples = np.random.randn(n_samples_per_class, input_dim) * 0.5 + center # Add some noise
+8 X.append(class_samples)
+9 y.extend([i] * n_samples_per_class)
+10
+11 X = np.vstack(X) # Combine all class samples
+12 y = np.array(y) # Convert labels to a NumPy array
+13 return X, y
+D.4 Implementation Details on the MLP experiments in Figure LABEL:fig_mlp_gap
 We train a 4-layer MLP on MNIST.
@@
 $=$ $\max_{l\in[L]}\left(1-\frac{1}{\left(\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}}\right)\kappa_{l}}\right)\left(\mathcal{L}(w^{t})-\mathcal{L}^{*}\right).$
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The main results in the experimental sections match the claims in the abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The limitation is discussed in Section 5.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: The full set of assumptions are in Section 4.2 and the complete proof is in E.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: In Appendix D, we carefully describe all the needed information to reproduce the experimental results.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We provide code in the supplementary materials.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The experimental details are described in Appendix D.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: We did not provide the error bar since the experiments on large Transformers are too expensive to repeat for multiple runs.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The computation resource is described in Section D.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We conform with the NeurIPS code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Positive and negative social impacts are discussed in Section Broader Impacts.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: Our paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All assets are properly cited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: We do not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Our paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.